### v1

# Deep Reinforcement Learning-Based Active Flow Control of an Elliptical Cylinder: Transitioning from an Elliptical Cylinder to a Circular Cylinder and a Flat Plate

Abstract
This study investigates the effectiveness of Active Flow Control (AFC) technology supported by Deep Reinforcement Learning (DRL) in suppressing vortex shedding, reducing drag, and minimizing lift coefficients around elliptical cylinders. The aspect ratio ($Ar$) of the elliptical cylinder transitions from an ellipsoid ($Ar=2.0$) to a circular shape ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$). We utilize the Proximal Policy Optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on both the upper and lower surfaces of the cylinder. The focus is on optimizing fluid dynamics and examining the scalability of these control techniques across different geometric configurations. Our research findings indicate that, for elliptical cylinders with aspect ratios between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%. Notably, the DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities. Additionally, for aspect ratios from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%. This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding. This underscores the adaptability and potential of DRL-based AFC in managing complex fluid dynamics across diverse elliptical cylinder configurations.
preprint: AIP/123-QED
I Introduction
The development of active flow control technology has progressed from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineeringCollis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021). This technology aims to improve flow characteristics through active intervention in the flow field, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiencyCollis and Joslin (2004); Wang et al. (2018). Initially, research in active flow control primarily focused on theoretical analysis and simple experiments, exploring how to influence fluid flow with external inputs like jets, vibrations, or magnetic fieldsJahanmiri (2010); Wang et al. (2018); Awada, Younes, and Ilinca (2021). With the advancement of Computational Fluid Dynamics (CFD) and increased computational power, researchers have been able to gain a deeper understanding of complex flow responses through numerical simulations and develop more effective control strategiesAram et al. (2018); Jain, Yeo, and Chopra (2010); Li et al. (2022). In recent years, breakthroughs in machine learning and deep learning technologies have led to revolutionary progress in the field of active flow controlBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020). Researchers have begun using intelligent algorithms, such as Deep Reinforcement Learning (DRL), to automatically optimize control strategies, achieving precise control over highly nonlinear and variable flow fieldsBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021). These methods have not only improved the accuracy and real-time performance of control effects but also significantly expanded the application scope of active flow control technology, including but not limited to aerospace, automotive, energy, and environmental engineering sectorsVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020); Garnier et al. (2021); Razdan and Shah (2022).
DRL is an avant-garde methodology that amalgamates the principles of deep learning with reinforcement learning to optimize decision-making through interaction with an environmentArulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2022a). Deep learning excels in autonomously extracting intricate features from voluminous datasets, while reinforcement learning concentrates on how to make optimal decisions based on these features, guiding the learning process through rewardsJaniesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015). DRL leverages deep neural networks to represent the policies or value functions within reinforcement learning, thus empowering machines to tackle high-dimensional state spaces and complex decision-making challenges that were previously intractableArulkumaran et al. (2017); François-Lavet et al. (2018); Wang et al. (2022a). Notable for its robust non-linear modeling capabilities and adaptable learning mechanisms, DRL is capable of directly learning complex strategies from raw sensory data without the need for manually engineered featuresDargazany (2021). Moreover, DRL is inherently self-improving, continuously refining its strategies through ongoing interactions with the environmentWang et al. (2022a). This capacity for self-learning and adaptation bestows a significant advantage upon DRL, especially in resolving decision-making problems within dynamic settingsFrançois-Lavet et al. (2018); Wang et al. (2022a). The applications of DRL are extensive, encompassing domains such as gaming, autonomous vehicles, robotic control, natural language processing, and computer vision. In the gaming industry, breakthroughs like AlphaGo’s victory over world champions underscore DRL’s potentialGranter, Beck, and Papke (2017); in autonomous driving, DRL is employed for optimizing route planning and decision-makingSallab et al. (2017); in robotics, it enables robots to learn complex physical tasksZhao, Queralta, and Westerlund (2020). Additionally, DRL is being applied in scientific research areas like fluid mechanics and industrial optimization, showcasing its vast potential in managing complex systemsBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015).
In recent years, the application of DRL within the realm of fluid dynamics has seen a marked increase, with a primary focus on active flow control and shape optimizationBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al. (2021). Active flow control technology is a key challenge in the fields of fluid mechanics and engineering, including goals such as drag reduction, prevention of flow separation, control of vortex shedding, and enhancement of mixing efficiencyWang et al. (2018). Traditional control methodologies often rely on a priori knowledge and empirical rules, whereas DRL offers a novel pathway for autonomously learning optimal control strategies from dataCollis and Joslin (2004); Jahanmiri (2010). In the context of drag reduction applications, DRL is utilized to autonomously adjust the operation of control devices (e.g., jets, vibratory mechanisms) to minimize the fluidic drag around objects such as cylindrical bodies or car modelsGarnier et al. (2021); Razdan and Shah (2022). For vortex shedding control, DRL is capable of learning how to dynamically adjust control parameters to stabilize the flow field and reduce vibrations and noise caused by vortex sheddingVignon, Rabault, and Vinuesa (2023); Rabault et al. (2020). In the domain of shape optimization, DRL facilitates the automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal design strategiesBhola et al. (2023). With the continued expansion of computational resources and advancements in algorithms, the application of DRL in fluid dynamics is set to broaden further.
Fluid mechanics problems are characterized by a high degree of nonlinearity and multiscalar nature, involving high-dimensional state spaces and complex system dynamicsPozorski and Wacławczyk (2020); Reichstein et al. (2019). These challenges are often intractable for traditional mathematical models, demanding sophisticated control strategies, particularly when addressing phenomena such as unstable flows, turbulence, and multiphase flows, which require consideration of a wide range of spatiotemporal scalesSubramaniam (2020). The suitability of DRL methods for these kinds of fluid mechanics issues primarily rests on several key features: Firstly, DRL can handle the high-dimensional and complex state spaces intrinsic to fluid dynamics, effectively capturing the intricate dynamics presentXie et al. (2023). Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid mechanics systems with numerous variables and parameters, discerning patterns and regularities hidden within dataZhao et al. (2024). Secondly, DRL’s interactive and adaptive learning approach allows it to self-tune and optimize strategies in response to dynamic and uncertain fluid environments. Traditional control methods often struggle with such dynamicsBhansali et al. (2023); however, DRL models, through their interaction with the environment and trial-and-error processes, can independently learn and refine strategies to meet diverse fluidic conditions and task requirementsHusen, Chaudary, and Ahmad (2022). Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies. Control and optimization in fluid mechanics often necessitate foresight into long-term effectsVinuesa et al. (2022). DRL’s focus on long-term rewards enables models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanicsViquerat et al. (2022); Vinuesa (2024). In conclusion, the attributes of DRL render it highly suitable for addressing the control and optimization challenges in fluid mechanics.
Rabault et al.Rabault et al. (2019) pioneered the application of advanced DRL algorithms in active flow control. The author uses the DRL agent to control the mass flow rate of the synthetic jets on the upper and lower sides of the cylinder, effectively reducing the resistance around the cylinder by 8% and mitigating the detachment of the rear vortex. Rabault et al.; Tang et al.; Heess et al.; Ren, Rabault, and Tang; Jia and XuRabault et al. (2019); Tang et al. (2020); Heess et al. (2017); Ren, Rabault, and Tang (2021); Jia and Xu (2024a) applied the Proximal Policy Optimization (PPO) algorithm to active flow control (AFC) in the flow around a cylinder. The training results consistently showed a reduction in drag of 8% around the cylinder. Li and ZhangLi and Zhang (2022) investigated the flow stability and sensitivity in confined cylinder wakes under various blockage ratios. They used the analysis of flow stability and structure sensitivity to guide the manipulation of the flow field by DRL agents. Wang et al.Wang et al. (2022b) utilized an AFC method based on the PPO algorithm to reduce the drag on a cylindrical body by approximately 8%, and applied the Soft Actor-Critic (SAC) algorithm to decrease the drag around a square cylinder by about 13.6%. Jia and XuJia and Xu (2024b) employed the SAC algorithm within a DRL framework to achieve AFC around a square cylinder, effectively modulating the flow dynamics across varying Reynolds numbers. Wang et al.Wang et al. (2022c) used DRL methods for active control of flow through elliptical cylinders. Their training results prove that DRL can successfully learn active control strategies under four angles of attack and various long and short axis ratios. Fan et al.Fan et al. (2020) was the first to apply DRL to active flow control experiments, optimizing power efficiency by adjusting the rotational speed of two small cylinders behind the main cylinder. Ren, Wang, and TangRen, Wang, and Tang (2021) proposed a novel active flow control strategy to hide the hydrodynamic characteristics of bluff bodies, including strong shear forces and periodic shedding vortices, to avoid predators.
The application of AFC strategies based on DRL has been widely validated in various bluff body flow scenarios for their effectiveness in controlling vortex shedding and reducing drag. Among these, studies on flow around constrained cylinders under different blockage ratios represent the most thoroughly investigated areaRabault et al. (2019); Tang et al. (2020); Li and Zhang (2022); Wang et al. (2022b); Jia and Xu (2024a). By contrast, research related to flow around elliptic cylinders has seen relatively less exploration within this domain. Due to their unique shape, elliptic cylinders influence the development of separation points and boundary layers differently during flow, potentially leading to Hopf bifurcation and periodic vortex shedding at lower Reynolds numbers. These flow characteristics offer opportunities in aerospace, marine engineering, energy engineering, and modern architectural design for improving lift-to-drag ratios, reducing vortex-induced vibration, enhancing energy efficiency, and increasing structural stability. Therefore, an in-depth investigation of the flow characteristics around elliptic cylinders not only enhances our understanding of fluid dynamics but also promotes performance optimization and innovation in engineering design.
In summary,despite significant advancements in controlling flow around circular cylinders through DRL algorithms, the study of flow around elliptical cylinders presents a vast research landscape due to its more complex flow characteristics and control challenges. AFC based on DRL algorithms introduces a novel methodology that leverages learning from flow field information to intelligently adjust control strategies, aiming at reducing drag and suppressing vortex shedding oscillations. This research is dedicated to examining the impact of various $Ar$ of elliptical cylinders on flow stability and sensitivity to structural parameters. Building upon this, it explores the application of DRL-based AFC in managing flow around elliptical cylinders by utilizing the derived physical insights. By establishing a robust reinforcement learning environment and defining clear reward mechanisms, this study aims to identify and implement effective flow control strategies. The aim of this study is to enhance the flow stability around an elliptical cylinder using AFC technology based on DRL algorithm, serving as a reference for flow control around similarly complex-shaped objects.
The structure of this paper is organized as follows: Section II introduces the problem description and provides a comprehensive overview of Deep DRL and its framework for application in AFC. Section Section III presents the main results and analysis. Finally, Section IV concludes the paper by summarizing the key findings and discussing their implications.
II Problem description and methodology
II.1 Numerical simulation
Model configuration
The present study investigates the flow around two-dimensional elliptical cylinders with their major axis oriented perpendicular to the flow direction, as depicted in Fig. 1(a). The major axis of the constrained elliptical cylinder (denoted 2a) remains fixed, while the shape of the elliptical cylinder is changed by changing the value of 2b. The Reynolds number is calculated based on the length of the major axis, formulated as $Re=U(2a)/\nu$, where $U$ represents the velocity of the incoming uniform flow and $\nu$ denotes the kinematic viscosity. The aspect ratio ($Ar$) is defined as the ratio of the semi-minor axis ($b$) to the semi-major axis ($a$), expressed as $Ar=b/a$. For ease of comparison with circular cylinders, we define $2a$ as $D$, representing the diameter of an equivalent circular cylinder. Consequently, an $Ar$ of 0 designates a flat plate normal to the flow, while an $Ar$ of 1 denotes a circular cylinder. This study investigates a series of variations ranging from an elliptical cylinder $(Ar=2.0)$ to a circular cylinder $(Ar=1.0)$, and further to a flat plate $(Ar=0)$, as shown in Fig. 1(c).
Figure 1: (a) Definition of the problem, including relevant dimensions. (b) Establishment of boundary conditions. (c) This study examines an elliptical cylinder with varying aspect ratios: a normal flat plate corresponding to a cylinder with an aspect ratio of 0, a special case of a cylinder with an aspect ratio of 1, and an elliptical cylinder with an aspect ratio of 2.
Computational domain and coordinate system
An elliptical cylinder is placed within a computational domain bounded by a rectangular perimeter, where the diameter of the elliptical cylinder is defined as D, and the major axis length (2a) of the elliptical column is equivalent to D. The computational domain extends from -2D in front of the cylinder or elliptical column to 20D behind it in the flow direction, which is the direction of fluid flow. In the cross-flow direction, perpendicular to the flow direction, the domain ranges from -2D to 2.1D. Within this computational domain, a Cartesian coordinate system is established with the center of the cylinder or elliptical column as the origin O, where the direction of flow is defined along the x-axis, and the axis perpendicular to the flow direction is the y-axis. The configuration of the synthetic jets is depicted in Fig. 1(b), specifically positioned at two directions on the cylindrical or elliptical column, $\theta_{1}=90^{\circ}$ and $\theta_{2}=270^{\circ}$, representing symmetrical locations of the jet on either side of the object’s surface. The width of the jet opening is fixed at $\omega=10^{\circ}$.
Governing equations
Let $\Omega\subset\mathbb{R}^{nd}$ and $(0,T)$ be the spatial and temporal domains, respectively, where $nd$ is the number of space dimensions, and let $\Gamma$ denote the boundary of $\Omega$. The spatial and temporal coordinates are denoted by $x$ and $t$. This work investigates the dynamics of incompressible viscous fluid flow. The Navier–Stokes equations governing incompressible fluid flow are
$\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot(\nabla\mathbf{u})=-\nabla p+Re^{-1}\Delta\mathbf{u}\quad\text{on}\quad\Omega\times(0,T),$ (1a)
$\nabla\cdot\mathbf{u}=0\quad\text{on}\quad\Omega\times(0,T).$ (1b)
Here, $\bm{u}=(U,V)$ represents the velocity vector, where $U$ and $V$ are the velocity components along the $x$- and $y$-axes, respectively. $p$ is the thermodynamic pressure. The $Re$ is calculated based on the length of the major axis, formulated as $Re=U(2a)/\nu$, where $U$ represents the velocity of the incoming uniform flow and $\nu$ denotes the kinematic viscosity.
Boundary conditions
The upstream boundary of the computational domain, designated as $\Gamma_{\text{in}}$, is characterized by a parabolic velocity profile oriented in the $x$-axis, where the boundary conditions are specified such that the streamwise velocity component $U$ is equal to the mean velocity $U$ with a value of 1, and the transverse velocity component $V$ is set to zero. On the downstream boundary, denoted as $\Gamma_{\text{out}}$, a Neumann-type boundary condition is prescribed for the velocity field, ensuring that the stress vector is nullified. On the upper and lower boundaries $\Gamma_{\text{w}}$ , a slip-wall boundary condition is employed, which stipulates that both the velocity component perpendicular to these boundaries and the tangential stress vector component along them are assigned a value of zero. The synthetic jets boundaries are denoted by $\Gamma_{i}(i=1,2)$. The velocity within the jets adheres to a parabolic distribution, with the sign of the jet velocity indicative of fluid suction or injection. The design of the synthetic jets ensures mass flow balance, as expressed by the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, to maintain mass conservation in the system. Except for the jet, the non-slip solid wall boundary condition $\Gamma_{\text{w}}$ is adopted for the rest of the cylinder.
In compliance with the above requirements, the system is considered with the following boundary conditions
$-\rho\mathbf{n}\cdot\mathbf{p}+Re^{-1}(\mathbf{n}\cdot\nabla\mathbf{u})$ $=0\quad\text{on }\Gamma_{out},$ (2)
$\mathbf{u}$ $=0\quad\text{on }\Gamma_{w},$
$\mathbf{u}$ $=U\quad\text{on }\Gamma_{in},$
$\mathbf{u}$ $=f_{Q_{i}}\quad\text{on }\Gamma_{i},\quad i=1,2.$
Here, $U$ is the inflow velocity profile. $f_{Q_{i}}$ represents the radial velocity profiles simulating the suction or injection of fluid by the jets. The specific formulas of $U$ and $f_{Q_{i}}$ will be given below.
At the inlet $\Gamma_{\text{in}}$, the inflow velocity $U$ along the $x$-axis is prescribed by a parabolic velocity profile in the form,
$U=U_{m}\frac{(H-2y)(H+2y)}{H^{2}}.$ (3)
Here, $U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain.
The functions $f_{Q_{i}}$ represent radial velocity profiles that simulate the suction or injection mechanisms of the jets. These profiles are selected to ensure a smooth transition to the no-slip boundary conditions on the surfaces $\Gamma_{W}$ of the cylinder. Specifically, $f_{Q_{i}}$ is defined as $A(\theta;Q_{i})(x,y)$, with the modulation relying on the angular coordinate $\theta$ as illustrated in Fig. 2(b). For the jet characterized by a width of $\omega$ and centered at $\theta_{0}$, which is positioned on the cylinder with radius $R$, the modulation is established as follows:
$A(\theta;Q)=Q\frac{\pi}{2\omega R^{2}}\cos\left(\frac{\pi}{\omega}(\theta-\theta_{0})\right).$ (4)
Solver details
In this study, the incompressible flow solver is based on the open-source CFD package OpenFOAM, as described by Jasak et al.Jasak (2009); Jasak et al. (2007) OpenFOAM is a widely used and validated CFD software package that provides robust numerical algorithms for solving the Navier-Stokes equations. The solver in OpenFOAM utilizes the finite volume method to discretize the computational domain into a mesh made up of control volumes. To ensure numerical stability, the time step is chosen as $\Delta t$ = 0.0005.
Grid system
Taking the computational domain of an elliptical cylinder with an $Ar$ of 1 as an exemplar, the domain undergoes discretization utilizing a synergistic approach that incorporates both structured (quadrilateral) and unstructured (triangular) meshes. As delineated in Fig. 2, this process results in the domain being segmented into 18,484 mesh elements. Herein, the vicinity of the elliptical cylinder is discretized with multiple layers of quadrilateral meshes, whereas the remaining regions are discretized using triangular meshes. Fig. 2(a) reveals the global discretization scheme of the computational domain, highlighting the comprehensive approach to mesh division. Fig. 2(b) offers insight into the mesh partitioning adjacent to the elliptical cylinder, with the multi-layered quadrilateral mesh around the cylinder being significantly finer compared to the triangular mesh. Fig. 2(c) illustrates the multi-layered quadrilateral mesh division around the elliptical cylinder and details the integration strategy between structured and unstructured meshes. Lastly, Fig. 2(d) provides a closer inspection of the mesh division surrounding the elliptical cylinder, emphasizing the precision of the discretization near the structure. The mesh division method for ellipses with $Ar$ other than 1 follows the same approach as that for $Ar=1$.
Figure 2: Numerical simulation of water flowing through an elliptical cylinder, close-up view of the mesh at $Ar=1$. The computational domain is discretized into 18,484 mesh elements, incorporating both triangular and quadrilateral meshes. (a) Discretization of the computational domain is viewed from a global perspective. (b) A comprehensive display of the mesh around the elliptical cylinder. (c) Presentation of the multi-layered quadrilateral meshes surrounding the elliptical cylinder. (d) A magnified view of the quadrilateral meshes in the vicinity of the elliptical cylinder.
Quantities of interest
The lift coefficient ($C_{L}$) and drag coefficient ($C_{D}$) are defined as
$C_{L}=\frac{F_{L}}{0.5\rho\overline{U}^{2}D},\quad C_{D}=\frac{F_{D}}{0.5\rho\overline{U}^{2}D}.$ (5)
Here, $F_{L}$ and $F_{D}$ represent the lift and drag forces integrated on the surface of the elliptical cylinder, respectively, and $\rho$ is the fluid density. The Strouhal number $(St)$ is used to describe the characteristic frequency of oscillatory flow phenomena and is defined as follows:
$St=\frac{f\cdot D}{U}.$ (6)
Where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$. $D$ is the characteristic length of the elliptical cylinder, and $\overline{U}$ is the mean velocity of the upstream flow.
Grid independence
Taking the case with an $Ar$ of 1 as an example, this study explores mesh convergence and validates the numerical method utilized. Tab. 1 presents the simulation results using three meshes with differing resolutions, enabling a comparative analysis of the maximum drag coefficient $C_{D,\max}$, the mean of the maximum drag coefficient $C_{D,\text{mean}}$, the maximum lift coefficient $C_{L,\max}$, and the $St$. Furthermore, the results of the numerical simulations are compared with the findings of Rabault et al. Rabault et al. (2019) and Schäfer et al. Schäfer et al. (1996) for validation. The simulations conducted using a coarse mesh revealed significant deviations in the drag and lift coefficients when compared to the results by Rabault et al. Rabault et al. (2019), questioning the accuracy of such numerical simulations. On the other hand, the fine mesh, characterized by a high density of grid points, considerably slows down the training speed for reinforcement learning, presenting a challenge for efficient simulation execution. Our aim is to perform numerical simulations with the least number of mesh elements without compromising on computational accuracy. Consequently, the discretization strategy of the main mesh was adopted for subsequent numerical simulations and reinforcement learning training. This strategy not only satisfies the computational accuracy requirements but also keeps the mesh element count to a manageable level. Comparison of the computational results from the main mesh with those obtained by Rabault et al. Rabault et al. (2019). shows very little difference, indicating a high level of accuracy. Furthermore, the accuracy of the main mesh’s results lies within the upper and lower bounds provided by standard benchmarksSchäfer et al. (1996), confirming that the computational precision of the main mesh is adequate for further training purposes. The above analysis ensures the best balance between numerical accuracy and computational efficiency. The methodology for examining mesh dependency for ellipses with $Ar$ other than 1 is consistent with the procedure established for $Ar=1$.
Table 1: Grid Independence Test for an Elliptical Cylinder with $Ar=1$ at $Re=100$.
Configuration Mesh resolution Mesh $C_{D,max}$ $C_{D,mean}$ $C_{L,max}$ $St$
Rabault et al.Rabault et al. (2019) - 9262 3.205 -
Schäfer et al.Schäfer et al. (1996) - - 3.220–3.240 - 0.990–1.010 0.295–0.305
Coarse 10540 3.242 3.224 1.052 0.304
$Ar=1$ Medium 18484 3.225 3.205 0.990 0.300
Fine 25624 3.228 3.207 0.992 0.301
II.2 Deep Reinforcement learning
Deep learning
Deep learning, a subset of machine learning, employs multi-layer artificial neural networks to model and learn intricate nonlinear relationships. Fig. 3(a) demonstrates how deep learning captures the representation of input data across multiple layers of a neural network, including the input layer, several hidden layers, and the output layer. Each layer is composed of numerous neural units, with information transmitted between nodes via connection weights. The fundamental concept behind deep learning is its ability to autonomously learn data representations through successive layers. Utilizing the backpropagation algorithm, the network autonomously discerns the transformation process from raw data to sophisticated, high-level abstract features, eliminating the need for manual feature extraction.
(a) DL
(b) RL
(c) DRL
Figure 3: The basic architecture of deep learning, reinforcement learning, and deep reinforcement learning.
Reinforcement learning
In the domain of Reinforcement Learning, the interaction between an agent and its environment is the core mechanism by which learning is facilitated. As illustrated in Figure 3(b), this interaction proceeds in a cyclical pattern. At each timestep $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the relevant information necessary for making decisions. Based on this observation, the agent selects and executes an action $a_{t}$, according to its policy $\pi$. This policy, which is essentially a strategy that maps states to actions, may function deterministically or probabilistically. After the action is executed, the environment transitions to a new state $s_{t+1}$ and awards the agent with an immediate reward $r_{t+1}$. This reward, quantified as a numerical value, evaluates the action’s efficacy in progressing towards the agent’s objectives. Subsequently, the agent assimilates this feedback—comprised of both the reward and the new state—to refine its policy, employing RL algorithms designed to optimize the accumulation of long-term rewards.
Deep Reinforcement learning
Deep learning, a specialized subset of machine learning, excels at identifying patterns in extensive datasets through its multi-layer neural network architecture. This approach enables computers to derive complex representations directly from raw data, leading to revolutionary advances in fields like image recognition, natural language processing, and predictive analytics. DRL marries the potent perceptual capabilities of deep learning with the decision-making and optimization capabilities of reinforcement learning, effectively overcoming their individual limitations. Deep learning addresses the challenge of processing original, high-dimensional input data, which is problematic for traditional reinforcement learning, while reinforcement learning provides a structured decision-making framework and a goal-oriented learning mechanism that enhances the capabilities of deep learning. As depicted in Figure 3(c), a deep learning network transforms the intricate state of the environment into a compressed and efficient feature representation. This transformation yields a more manageable and interpretable state space for reinforcement learning algorithms. In DRL, reinforcement learning leverages these high-level features to learn and optimize decision-making policies. These policies are typically parameterized by deep neural networks, implying that the selection of actions is influenced not only by the current observation of the environmental state but also by the learned weights within the network.
II.3 PPO algorithm
PPO is an advanced DRL algorithm suitable for solving reinforcement learning problems in discrete and continuous action spaces. The core idea is to improve performance by maximizing approximate policy optimization while maintaining the stability of policy updates. In PPO, the agent interacts with the environment, collects empirical data, and uses this data to construct an objective function. The objective function takes into account dominance estimates from empirical data and is used to measure the direction of policy improvement. In the PPO algorithm, the agent interacts with the environment to generate a series of trajectories ($\tau$). A trajectory is a sequence of state-action pairs produced by the agent following the current policy interacting with the environment until reaching a terminal state or a predetermined maximum number of steps. Symbolically, a trajectory $\tau$ can be represented as:
$\tau=(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},s_{2},\ldots,s_{T-1},a_{T-1},r_{T},s_{T}),$ (7)
where, $s_{t}$ represents the state at time step $t$; $a_{t}$ represents the action at time step $t$; $r_{t+1}$ represents the immediate reward obtained after taking action $a_{t}$; $T$ represents the length of the trajectory; $s_{0}$ is the initial state, and $s_{T}$ is the terminal state. The cumulative reward $R(\tau)$ of a trajectory $\tau$ is obtained by summing the rewards $r_{t}$ obtained at each time step $t$:
$R(\tau)=\sum_{t=0}^{T}r_{t},$ (8)
The objective function $J(\pi_{\theta})$ in the PPO algorithm aims to maximize the expected cumulative reward obtained by following the policy $\pi_{\theta}$. Mathematically, this objective function is expressed as:
$J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)],$ (9)
where $J(\pi_{\theta})$ represents the expected return under policy $\pi_{\theta}$, with the expectation taken over trajectories sampled from the policy distribution $\pi_{\theta}$. The term $R(\tau)$ denotes the total reward accumulated along a trajectory $\tau$. The policy gradient $\nabla_{\theta}J(\pi_{\theta})$, indicating the direction of steepest ascent in the objective function $J(\pi_{\theta})$, is computed using the policy gradient theorem. The gradient of the expected return is expressed as the expectation of the gradient of the log-probability of trajectories weighted by the return:
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[\nabla_{\theta}\log P(\tau\mid\theta)R(\tau)],$ (10)
where $\nabla_{\theta}\log P(\tau\mid\theta)$ represents the gradient of the log-probability of trajectories with respect to the policy parameters $\theta$. This computation involves sampling trajectories from the policy distribution $\pi_{\theta}$ and weighting the log-probability gradients by their respective returns. The policy parameters $\theta$ are updated iteratively using gradient ascent, with the update rule:
$\theta_{k+1}=\theta_{k}+\alpha\nabla_{\theta}J(\pi_{\theta}),$ (11)
where $\theta_{k}$ represents the policy parameters at the $k$-th iteration, $\alpha$ is the learning rate controlling the step size, and $\nabla_{\theta}J(\pi_{\theta})$ denotes the policy gradient.
In the PPO algorithm, the trajectories generated by the agent are instrumental in shaping the policy update process. The objective function, $J(\pi_{\theta})$, is designed to enhance the agent’s performance by maximizing the expected cumulative reward under the policy $\pi_{\theta}$. This optimization involves iteratively updating the policy parameters $\theta$ using the policy gradient $\nabla_{\theta}J(\pi_{\theta})$, while maintaining stability and preventing substantial policy shifts. The policy gradient is estimated using a collection of trajectories, $\mathcal{D}=\{\tau_{i}\}_{i=1,\ldots,N}$, each generated by the agent interacting with the environment according to the policy. The policy gradient can be calculated as follows:
$\hat{g}=\frac{1}{|\mathcal{D}|}\sum_{\tau\in\mathcal{D}}\sum_{t=0}^{T}\nabla_{\theta}\log\left(\pi_{\theta}(a_{t}\mid s_{t})\right)R(\tau),$ (12)
where $|\mathcal{D}|$ represents the number of trajectories in $\mathcal{D}$, and $T$ is the length of each trajectory. This formulation offers a simplified approach to computing the policy gradient. Moreover, PPO commonly employs a clipped surrogate objective function to facilitate more stable policy updates. This objective is expressed as:
$L^{CLIP}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{\tau\in\mathcal{D}}\sum_{t=0}^{T}\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right),$ (13)
where $r_{t}(\theta)$ is the likelihood ratio between the old and new policies, $A_{t}$ denotes the advantage function, and $\epsilon$ is a hyperparameter that defines the clipping range. The inclusion of this clipped surrogate objective aids in moderating the extent of policy updates, thereby stabilizing the training process.
In summary, the PPO algorithm presents a robust framework for the training of deep reinforcement learning agents. It operates through an iterative process of updating policy parameters aimed at maximizing the expected cumulative reward. Crucially, PPO maintains stability and mitigates the risk of substantial policy deviations. Employing policy gradients along with clipped surrogate objectives, PPO ensures effective and stable training across a diverse array of reinforcement learning tasks. This methodology not only enhances the efficiency of learning but also contributes significantly to the consistency and reliability of the training process.
II.4 DRL-Enhanced Active Flow Control
This section outlines the design of a framework utilizing DRL algorithms to address active flow control problems. Firstly, it is necessary to define two fundamental elements in DRL: the agent and the environment. The agent can be designed as algorithms suitable for solving continuous action control problems, such as PPO agent or SAC agent. The CFD simulation of the flow around an elliptical cylinder is chosen as the environment for the agent to interact with. Rabault et al.Rabault et al. presented the main structure of the custom environment class, encouraging readers to delve into the detailed content. This framework design applies deep reinforcement learning methods to solve the active flow control problem, enabling the agent to continuously learn and optimize its behavior through interaction with the environment to achieve the control goal. Fig. 4 illustrates the fundamental framework of DRL, the computational process of a single episode during training, and the formulas utilized.
Figure 4: DRL leverages the concurrent execution of multiple environments within a parallel computational framework. Each iteration consists of two distinct phases: exploration and optimization. During the exploration phase, the agent engages with the environment via its current policy $\pi$, executing actions $a_{t}$, receiving states $s_{t}$ and rewards $r_{t}$, and noting terminal states. Upon the conclusion of the iteration, the parallel computation across $n$ environments yields a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, with each trajectory $\tau_{i}$ comprising the sequence of states, actions, and rewards as $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1},r_{1}),\ldots,(s_{T},a_{T},r_{T})\}$, where $(s_{t},a_{t},r_{t})$ represents the state, action , and reward at time-step $t$, respectively, for the $i$-th environment. In the optimization phase, the agent samples a minibatch of trajectories from the experience replay buffer to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$. The parameters $\theta$ of the policy network are updated using gradient ascent, i.e., $\theta\leftarrow\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $\alpha$ denotes the learning rate. This process iterates, with the agent deploying the updated policy $\pi_{\theta^{\prime}}$ to interact with the environment, computing the advantage function $A(s_{t},a_{t})$, and recalculating the policy gradient for subsequent updates of $\theta$. Through iterative refinement, the policy network incrementally enhances its action selection strategy within the environment. The formula delineates the update rule applied to the policy network during the optimization phase.
  • State $s_{t}$: Traditional state spaces are often too expansive for effective policy learning, requiring operations in partially observable environments. As a result, an agent’s performance heavily relies on the quality and relevance of the data it can observe. The impact of partial observability of states $s_{t}$ on control results was mentioned in Viquerat et al.Viquerat et al. (2022)’s research.Integrating instantaneous velocity or pressure information from CFD simulations to characterize the environmental state $s_{t}$, the DRL agent can effectively perceive and analyze the current flow conditions. To achieve the control objectives of reducing drag and lift around the elliptical cylinder and controlling the shedding of vortices in the wake field, multiple probes are strategically placed around the elliptical cylinder and within its wake field. This method is advantageous for gathering comprehensive data pertaining to the flow dynamics around the cylinder and subsequent wake characteristics. The design of probes placement in this study was based on the design principles articulated by Rabault et al.Rabault et al. (2019), which advocate strategic positioning of probes around the cylinder and within the wake region of the cylinder.In their studies, Jia and XuJia and Xu (2024a) and Wang et al.Wang et al. (2022b) all referred to the probes arrangement scheme proposed by Rabault et al.Rabault et al. (2019).
  • Action $a_{t}$ : The control action involves the blowing and suction of two synthetic jets on the elliptical cylinder. The flow rates of the two synthetic jets are constrained such that $Q_{\text{lower}}+Q_{\text{upper}}=0$. Therefore, the RL agent manipulates the velocity of the upper synthetic jets, while the lower synthetic jets has the same velocity but in the opposite direction.The velocity of the synthetic jets is set to be less than 2% of the inlet mean velocity. This restriction is imposed to prevent the agent from taking shortcuts and using excessively high jet velocities to achieve drag reduction, which contradicts the fundamental principle of active flow control: energy efficiency. Furthermore, each control action lasts for 50 time steps, termed as a viscous action. Allowing faster actions could harm overall performance as they need time to influence the environment. Ensuring smooth transitions between current and updated actions is crucial. Let $a_{t}$ and $a_{t+1}$ represent the action magnitudes at consecutive time steps. To ensure a gradual transition between these actions and mitigate abrupt changes in jet velocity, a smoothing function $S$ can be defined mathematically. The function $S$ interpolates between the previous action magnitude $V_{\Gamma_{1},T_{i-1}}$ and the target action magnitude $a$ at time step $i$, resulting in an updated value $V_{\Gamma_{1},T_{i}}$. This smoothing function can be expressed as:
    $S(V_{\Gamma_{1},T_{i}},a,V_{\Gamma_{1},T_{i-1}})=V_{\Gamma_{1},T_{i}}+\beta\cdot(a-V_{\Gamma_{1},T_{i-1}}),$ (14)
    where $V_{\Gamma_{1},T_{i}}$ represents the current value at time step $i$, $\beta$ is a coefficient determining the extent of adjustment towards the target action $a$, and $V_{\Gamma_{1},T_{i-1}}$ is the value at the previous time step $i-1$. This function ensures a smooth transition between consecutive actions, with the parameter $\beta$ controlling the smoothness of the transition.
  • Reward $r_{t}$ : This study incorporates the average drag coefficient (without control), the time-averaged value of the drag coefficient, and the time-averaged value of the lift coefficient (averaged over the duration of the action). The primary objective is to minimize the drag while maintaining a low lift coefficient; hence, the weight of the drag coefficient in the reward function is set to 1, while the lift coefficient is assigned a weight $w$, typically chosen to be 0.1 or 0.2. This hyperparameter $w$ is determined through tuning. Including a constant (the average drag coefficient without control) in the reward function helps guide it towards positive rewards, which positively affects the convergence speed of the learning curve. The reward function $f(\cdot)$ translates the drag and lift coefficients into a numerical reward assessing the agent’s behavior. The specific reward function takes the following form:
    $r_{T_{i}}=C_{D,0}-\left(C_{D}\right)_{T_{i}}-\omega\left|\left(C_{L}\right)_{T_{i}}\right|,$ (15)
    $C_{D,0}$ represents the baseline drag coefficient, serving as a reference point for drag reduction. $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective of minimizing this value relative to the baseline. $(C_{L})_{T_{i}}$ signifies the lift coefficient at time step $T_{i}$, whose absolute value is penalized to mitigate lift forces that may destabilize the flow around the bluff body. $\omega$ is a weighting factor that quantifies the trade-off between minimizing drag and controlling lift fluctuations.
  • Parameter. Time granularity in the DRL framework is crucially set based on the fluid dynamics problems studied and significantly affects agent performance and learning task complexity. Precise adjustments of this hyperparameter, informed by prior research and our ongoing strategy refinement, are vital for achieving optimized control outcomes. Each action is applied for a duration of 0.025 seconds (Corresponds to 25 numerical simulation time steps), and each episode consists of 100 control actions. The maximum total time for an episode is defined as $T_{\max}=T_{100}=2.5$ nondimensional time units. The shedding frequency of vortices around the elliptical cylinder in this study ranges from 0.3 to 0.4, corresponding to 6.25 to 8.33 vortex shedding cycles per episode. The length of episodes is intentionally chosen to encompass multiple shedding cycles, facilitating effective learning of the control algorithm by the agent through observation and adaptation of flow dynamics over several cycles.
II.5 High-Performance Computing in DRL and CFD
The DRL training and numerical simulation in this study were conducted on a high-performance computing(HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60GHz and comprising a total of 64 cores distributed evenly across two sockets, each containing 32 cores. Numerical simulations were performed using the eighth version of the open-source software platform, OpenFOAM® version 8.Jasak et al. (2007) The deployment of deep reinforcement learning algorithms was based on the Tensorforce library.Schaarschmidt, Kuhnle, and Fricke (2017) The ANN definitions and gradient descent algorithms were facilitated using the TensorFlow open-source library.Abadi et al. (2016) For DRL implementation, agent was loaded using the Tensorforce platform with the PPO algorithm, and the environment was constructed using the Gym interface.Brockman et al. (2016) Fig. 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process. This parallel strategy was inspired by the research of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al. (2022b), and also incorporated insights from the work of Wang et al.Wang et al. (2022b) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.
III Results
III.1 Baseline Flow
Vortex shedding phenomenon
Figure 5: Instantaneous Vorticity Contours for Flow Around an Elliptic Cylinder at $Re=100$, Varying from $Ar=2$ to $Ar=0$. Purple and orange indicate positive and negative vorticity, respectively. In all the depicted vorticity contours, purple represents the clockwise vortices shed from the upper side of the elliptical cylinder, while orange represents the counterclockwise vortices shed from the lower side.
Fig. 5 presents the vortex shedding patterns in the wake region formed around elliptical cylinders with varying $Ar$. The subfigures (a) to (j) depict the vortex structures that have reached a fully developed state, illustrating the progression from an aspect ratio of 0 to 2. At a Reynolds number of 100, around elliptical cylinders with $Ar$ varying from 0 to 2, persistent unstable vortices are observed at the computational domain’s outlet, indicating the continuation of fluid instability along the outflow direction. However, as the fluid progresses downstream, there is a diminishing trend in the peak vortex intensity. This attenuation of vorticity is predominantly attributed to viscous diffusion effects prevalent in low Reynolds number regimes. This is consistent with the study by Johnson, Thompson, and HouriganJohnson, Thompson, and Hourigan (2004). Between $Ar$ of 2 and 1.25, in addition to the recirculation bubbles downstream of the cylinder, two additional recirculation bubbles form upstream near the wall. As the $Ar$ decreases further, the confined flow exhibits only one recirculation bubble. Similar to observations by Sahin and OwensSahin and Owens (2004), under the influence of blockage ratio and Reynolds number, two additional recirculation bubbles appear on the solid sidewalls of the computational domain. At specific Reynolds numbers, this will lead to another Hopf bifurcation at $Re_{\text{c2}}$.
For an elliptical cylinder with an $Ar=2$, the wake field does not exhibit vortex shedding, and the zero vorticity contour lines show slight fluctuations. This suggests that despite the overall stability of the wake, there is still minor flow instability. As the $Ar$ decreases to 1.75 and 1.5, the wake field begins to exhibit more pronounced vortex shedding phenomena. The vortices become more compact, and there is increased interaction between positive and negative vortices through shearing, stretching, and detachment processes. As the $Ar$ decreases to 1, transforming the cylinder into a perfect circle, the vortex structure becomes more regular, forming a typical von Kármán vortex street. This "standard aligned BvK street" is characterized by a series of regularly spaced vortices that alternately form on both sides of the cylinder. As the $Ar$ decreases from $Ar=0.75$ to $Ar=0.25$, the recirculation area gradually diminishes, leading to a more compact vortex structure and an increase in shedding frequency. At $Ar=0.1$, for extremely low $Ar$, the recirculation area further decreases, resulting in an increase in both the frequency and intensity of vortex shedding. At $Ar=0$, when the cylinder degenerates into a flat plate, particularly in the zero-vortex region behind the ellipse, the acceleration of vortex shedding becomes more pronounced. This indicates an increase in flow instability and irregularity in vortex size and shape.
Fluctuation velocity
Strykowski and SreenivasanStrykowski and Sreenivasan (1990) demonstrated that the introduction of minor disturbances in a targeted region can alter vortex formation, effectively reducing or controlling vortex shedding. Similarly, our study employs small perturbations to disturb the flow field with the aim of decreasing drag and suppressing vortex shedding. Furthermore, the judicious incorporation of fluid physics knowledge within the control agents can streamline their architecture without compromising performance, as exemplified by Pastoor et al.Pastoor et al. (2008), where phase control in drag reduction of a bluff body is primarily founded upon the understanding of decoupled shear layer development and wake processes. In this study, the strategic placement of probes was thoughtfully engineered to align with the dynamic and controllable aspects of the flow field. Particular attention was paid to its inherent instability and pulsatility, ensuring that the probes could effectively capture the critical fluctuations necessary for a comprehensive analysis of the flow dynamics. The research conducted by Protas and WesfreidProtas and Wesfreid (2002) explores the dynamics of the total drag coefficient, $C_{D}$, which is defined by the equation $C_{D}=C_{D}^{b}+C_{D}^{0}$. Here, $C_{D}^{b}$ represents the base drag coefficient, a static value consistent across specific $Re$, encapsulating the inherent drag associated with the baseline flow. Conversely, the correctable component $C_{D}^{0}$ reflects the influence of oscillatory flow phenomena on the drag coefficient, providing an avenue for adjustment through strategic manipulation of Bénard–von Kármán vortices. In our study, the instantaneous velocity field is segmented into the mean velocity field and the fluctuating velocity field. We hypothesize that probes equipped to detect these pulsatile and oscillatory dynamics can leverage targeted minor disturbances to effectively neutralize these instabilities. Such interventions are anticipated to facilitate the suppression of vortex shedding, thereby enhancing the control of flow behaviors.
Figure 6: For the case of $Ar=1$, the distribution of probes is depicted across various velocity fields: (a) Instantaneous velocity, (b) Fluctuating velocity.
We delineates the instantaneous and fluctuating velocity fields behind an elliptical cylinder during full vortex development, as illustrated in Fig. 6. The strategic placement of probes is predicated on the objectives of reducing drag and lift around the elliptical cylinder while simultaneously suppressing vortex shedding. Probes are positioned around and downstream of the elliptical cylinder, specifically within the vortex shedding region. The probes around the elliptical cylinder aim to diminish the drag and lift coefficients integral to our control objectives. Downstream, probes are strategically deployed in the wake to facilitate the extraction of flow field information from the regions of maximum fluctuating velocity, where intense turbulent motion typically occurs due to vortex instability and enhanced turbulent energy. The probe placement in the cylinder’s wake encompasses the entire recirculation area, which is critical for capturing the most unstable and intensely turbulent information in the wake region. This configuration plays a significant role in reducing the fluctuating components of the drag coefficient, suppressing vortex shedding, and stabilizing the wake flow. In Fig. 6, an ellipse with an $Ar$ of 1 serves as the prototype. The methodology for designing probe placements for other $Ar$ adheres to the same principles established for $Ar=1$.
III.2 Effectiveness of AFC
This section describes the performance of AFC using the DRL algorithm, which manipulates the magnitude of synthetic jets to achieve targeted flow control. Within a flow configuration at $Re=100$, AFC is applied to elliptical cylinders with $Ar$ ranging from 2 to 0, as well as to cylindrical and flat plate configurations. The evaluation of the control’s effectiveness and performance is based on several criteria: comparisons of the velocity fields before and after control, reductions in drag coefficient, variations in lift coefficient, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
Figure 7: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline simulation. Left side of each row: result with active flow control. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$.
Fig. 7 and Fig. 8 provide flow field snapshots of elliptical cylinders under various $A$, encompassing both baseline (uncontrolled) and controlled scenarios. This section primarily describes the flow structures observed in the baseline cases. For elliptical cylinders with $Ar=2$, the baseline flow field inherently displays high stability, with no vortex shedding observed. Consequently, the AFC did not produce a noticeable effect. As the aspect ratio decreases to $Ar=1.75$, vortices begin to detach alternately from both sides of the elliptical cylinder at reduced frequencies and intensities. These detaching vortices organize into a relatively orderly and stable vortex street. By applying AFC based on DRL, the vortex shedding in the wake of the elliptical cylinder is completely suppressed, resulting in a stabilized wake field. At an aspect ratio of $Ar=1.5$, an alternating vortex structure forms behind the elliptical cylinder. These structures possess certain strengths and frequencies, yet the extent of vortex shedding remains comparatively mild. Further reduction in the aspect ratio to $Ar=1.25$ leads to a more pronounced and frequent shedding of vortical structures from the wake, indicating a transition towards more significant flow instability. The recirculation bubble downstream of the elliptical cylinder diminishes, oscillates, and becomes unstable. When the aspect ratio is $Ar=1.5$ and $Ar=1.25$, a stable, regular alternating vortex structure forms behind the elliptical cylinder. After the application of AFC, the shedding of vortices behind the elliptical cylinder is completely suppressed, transitioning the wake field into a stable flow structure. At $Ar=1$ and $Ar=0.75$, the area of the recirculation bubbles formed behind the elliptical cylinder decreases, leading to more prominent and frequent vortex shedding within the wake, indicating a transition towards significant flow instability. Following the implementation of AFC, the alternating vortex street behind the elliptical cylinder is transformed into a stable wake field.
Fig. 7 depicts the instantaneous velocity fields both baseline and post-implementation of control measures across a range of $Ar$ from 2 to 0.75 for elliptical cylinders. At $Ar=2$, the baseline wake field is notably stable. However, when the $Ar$ is reduced below 1.75, alterations in structural parameters provoke enhanced vortex shedding behind the ellipse, characterized by escalating frequency and intensity as the $Ar$ diminishes. The efficacy of AFC is underscored through the deployment of a control strategy, which effectively attenuates vortex shedding behind the elliptical cylinder for $Ar=1.75$, 1.5, 1.25, 1, and 0.75. This intervention demonstrates the significant capabilities of the DRL-based AFC technology in manipulating vortex shedding within the wake fields of elliptical cylinders across varying $Ar$, thereby enhancing the flow stability of the configurations under study.
Figure 8: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline flow. Left side of each row: result with active flow control. (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
As the $Ar$ is further reduced to 0.5, 0.25, and 0.1, eventually transitioning to a flat plate configuration, the vortex shedding phenomena behind the elliptical cylinder become more frequent and intense. It is evident that decreasing the $Ar$ leads to a significant increase in the circulation entering the wake, as the flow velocity increases when water moves around the top of the cylinder. This observation aligns with the findings of Thompson et al.Thompson et al. (2014). This underscores that as the $Ar$ of an elliptical cylinder gradually diminishes towards zero, it induces an enhancement in the flow instability of the wake field, consequently leading to the release of more energy in the form of vortices. This dynamic highlights the critical role of structural parameters in influencing the energy distribution and stability of the flow field as it interacts with varying geometrical configurations.
However, when utilizing the AFC technique to manipulate the flow field, the control outcomes differ from those observed at $Ar=0.75$ and higher $Ar$; vortex shedding behind the elliptical cylinder is not completely suppressed. At $Ar=0.5$, analysis of the instantaneous velocity field reveals that the recirculation area behind the elliptical cylinder is significantly elongated, and the shedding of vortices is notably subdued. Nevertheless, the flow field has not fully stabilized, as there remains a continual detachment of alternating vortex structures from behind the elliptical cylinder, resembling elongated vortex formations. At $Ar=0.25$, the recirculation bubble behind the elliptical cylinder is similarly elongated as seen at $Ar=0.5$, but with a notable difference: the bubble now exhibits significant oscillations. In the controlled wake field, the phenomenon of flow separation following fluid passage around the elliptical cylinder and subsequent reattachment is also observable. Despite these dynamics, the instability of the downstream velocity field has been relatively stabilized compared to the baseline state, indicating the effectiveness in delaying vortex shedding. At $Ar=0.1$ or 0, compared to the baseline condition, the control of the flow field has primarily delayed rather than completely suppressed vortex shedding behind the elliptical cylinder. The recirculation bubble behind the cylinder is elongated, yet it exhibits significant oscillations. This indicates that while the AFC has extended the duration before vortex detachment, it has not fully stabilized the vortex structures within the wake.
In summary, at higher $Ar$ ($Ar=0.75$, 1, 1.25, 1.5, 1.75), the DRL-based control strategy can fully suppress vortex formation, thereby stabilizing the wake field. However, at lower $Ar$ ($Ar=0.5$, 0.25, 0.1, 0), the effectiveness of the control strategy is limited to mitigating vortex shedding within the wake field to some extent. It delays the vortex detachment process and enlarges the recirculation area behind the elliptical cylinder, but oscillations in the recirculation bubble persist, and the downstream flow field continues to exhibit signs of instability.
Figure 9: Comparison of $C_{D}$ and $C_{L}$ before and after control.(a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
Fig. 9 presents a comparative analysis of the $C_{D}$ and $C_{L}$ before and after control intervention over $Ar$ ranging from 2 to 0.75. At an $Ar$ of 2, both $C_{D}$ and $C_{L}$ are stable, indicating that AFC does not have any effect. However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, upon activating synthetic jets control, there is an immediate, sharp response in $C_{D}$, which quickly descends to a value near its minimum. After half the total control duration, $C_{D}$ stabilizes and remains constant at this minimal value. Similarly, $C_{L}$ experiences significant fluctuations at the start of control but stabilizes near zero after half the total control duration. The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ are detailed in Tab. 2. These observations substantiate the effectiveness of the active flow control strategy underpinned by DRL algorithms. The activation of control swiftly alters the periodic behavior of both $C_{D}$ and $C_{L}$, facilitating the stabilization of $C_{D}$ and $C_{L}$ in the subsequent stages. This demonstrates the capability of the DRL-based active flow control approach to significantly reduce the drag around an elliptical cylinder to its lowest feasible value and diminish the lift to zero.
Fig. 10 illustrates the $C_{D}$ and $C_{L}$ for both baseline and controlled flow conditions around an elliptical cylinder at $Ar$ of 0.5, 0.25, 0.1, and 0. Under AFC, at $Ar=0.5$ and $Ar=0.25$, both $C_{D}$ and $C_{L}$ are reduced compared to the baseline flow, yet they do not achieve the stable state observed at $Ar=0.75$ and higher. Initially, both the $C_{D}$ and the $C_{L}$ experienced abrupt changes. However, they did not fully stabilize in the subsequent period. In the latter half of the control duration, $C_{D}$ and the $C_{L}$ exhibited fluctuations, but both the amplitude and frequency of these fluctuations were lower than those observed in the baseline flow. At $Ar=0.1$ and $Ar=0$, the controlled flow experiences a sharp decrease in the drag coefficient at the onset of control, followed by oscillations that persist at a lower frequency and amplitude than the baseline flow. Although complete stability is not achieved, the average value of the $C_{D}$ is significantly reduced. The $C_{L}$ undergoes an initial sharp decrease and approaches near zero, and after half the total duration of control, it oscillates at a lower frequency.
Figure 10: Snapshots of the velocity magnitude illustrating the effect of the active flow control. Left side of each row: baseline flow. Left side of each row: result with active flow control. (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
Figure 11: Action, which is the mass flow rate of the synthetic jets. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
Fig. 11 illustrates the time history curve of the action (the mass flow rate of synthetic jets), When performing AFC using DRL for flow around an elliptical cylinder with $Ar$ between 0 and 2. When $Ar=2$, the AFC is ineffective, and the value of the action, pertaining to the control input, remains close to zero. From $Ar=1.75$ to $Ar=0.75$, during the initial phases of control, the agent selects larger action values, corresponding to higher mass flow rates of synthetic jets. This indicates a deliberate strategy to assert flow control by introducing significant external energy inputs. As control progresses, the agent gradually diminishes the action values, thereby reducing the mass flow rate of the synthetic jets. Subsequently, in the mid to later stages, the agent’s action values converge towards zero, corresponding to a near-zero mass flow rate of synthetic jets. This behavior signifies the agent’s optimization towards an energy-saving state, maintaining flow characteristics with minimal control input. When $Ar=0.5$ and $Ar=0.25$, the value of the action initially surges but does not stabilize near zero in the later stages. For $Ar=0.1$ and $Ar=0$, the action value remains in a state of periodic fluctuation.
Overall, the results presented in Fig. 7 and Fig. 9 illustrate that the DRL-based control strategy exhibits energy efficiency and effectiveness within the range of elliptical $Ar=0.75$ to $Ar=1.75$. By progressively decreasing the mass flow rate of synthetic jets, the strategy achieves flow control, reduces drag and lift forces, and does so with reduced external energy input. However, at $Ar=0.5$ and $Ar=0.25$, the energy efficiency of the AFC strategy diminishes. At $Ar=0.1$ and $Ar=0$, the external energy utilized by the control strategy is relatively high compared to the inlet velocity, which will be detailed in Tab. 2.
Figure 12: The learning curve of PPO algorithm training when $Ar=2$ to 0. (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$. (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
In Fig. 12, we observe the learning curve during training of flow around an elliptical cylinder with different $Ar$. At $Ar=2$, the reward function value never reaches a positive figure, indicating that the DRL training did not achieve its objectives. For $Ar=1.75$ and $Ar=1.5$, the reward function initially increases from negative to positive values and then after 4000 episodes converges near the maximum reward value. At $Ar=1.25$, despite some initial fluctuations, the training reaches a relatively stable state after 3000 episodes and eventually settles into a stable plateau. For $Ar=1$ and $Ar=0.75$, there are significant initial fluctuations with the learning curve rising swiftly, indicating rapid progress in policy learning. After 2000 episodes, the training demonstrates a stable convergence trend, with minor fluctuations in the reward value between 2000 and 5000 episodes. This shows that the PPO algorithm has successfully learned an effective strategy, exhibiting commendable convergence and stability in performance. At $Ar=0.5$, the speed of training convergence begins to slow, trending towards convergence after 4000 episodes, with the reward values stabilizing between 4000 and 10000 episodes. At $Ar=0.25$, training approaches convergence after 8000 episodes, reaching a total of 20000 episodes, with a mid-to-late plateau marking stable convergence. At $Ar=0.1$ and $Ar=0$, the training convergence is extremely slow, with signs of convergence only appearing after nearly 12000 episodes.
Overall, the DRL algorithm exhibits rapid initial convergence and reaches a stable plateau during the mid-to-late stages for $Ar=1.25$, 1 and 0.75, showcasing excellent convergence and stability throughout the training process. Training also achieves stable convergence within 5000 episodes for $Ar=1.5$ and $Ar=1.75$. However, for $Ar=0.5$, 0.25, 0.1 and 0, as the flow phenomena become increasingly complex, the control challenge intensifies, leading to a noticeable slowdown in convergence speed. This indicates that for flow scenarios with higher degrees of instability, the convergence performance of DRL training is compromised, requiring a larger number of training iterations to reach stable convergence.
Tab. 2 presents the effectiveness of active flow control techniques based on DRL algorithms, targeting elliptical cylinders with $Ar$ ranging from 0.1 to 2. It meticulously details the values of $C_{D}$ and $C_{L}$ before and after the application of control, along with their respective rates of change. Additionally, it describes the magnitude of the mass flow rate of the synthetic jets utilized for control and its ratio relative to the inflow rate. Initially, at $Ar=2$, the wake field stability behind the elliptical cylinder is observed to be relatively high, resulting in no significant alteration in the drag experienced by the elliptical cylinder. When $Ar$ is reduced to 1.75, there is a noticeable decrease in drag around the elliptical cylinder by 0.9%. With further reduction in $Ar$, the rate of drag reduction progressively increases. In their study, Protas and Wesfreid (2002) elucidated that the drag coefficient consists of two components: (i) the invariant drag ($C_{D}^{b}$), which remains constant at a given Reynolds number; and (ii) the amendable drag correction ($C_{D}^{0}$), associated with oscillatory flow phenomena within the fluid, representing the additional resistance arising from fluid oscillations or fluctuations. This portion of drag can be adjusted through external interventions, indicating the influence of unstable flow regions and the potential for control. In this research, the flow field is controlled by utilizing "external excitation," such as synthetic jets, to mitigate the additional resistance caused by instability, thereby optimizing the overall fluid dynamic performance. As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, thereby enhancing the controllable component of the drag.
Table 2: The control effect of elliptical cylinder under DRL-based AFC.
$Ar$ Re $\overline{C}_{D,\text{Baseline}}$ $\overline{C}_{D,\text{Controlled}}$ Drag Reduction (%) $\overline{C}_{L,\text{Baseline}}$ $\overline{C}_{L,\text{Controlled}}$ Lift Reduction (%) $\overline{a}$ $\overline{a}$ ratio (%)
0 100 7.679 4.331 43.6 1.513 0.605 60.0 0.521 52.1
0.1 100 7.812 4.876 37.6 1.196 0.596 50.2 0.499 49.9
0.25 100 6.503 4.167 35.9 1.707 0.579 66.1 0.126 12.6
0.5 100 4.837 3.533 26.9 1.937 0.619 68.0 0.049 4.9
0.75 100 3.792 3.197 15.7 1.603 0.076 95.2 0.010 1.0
1 100 3.207 2.951 8.0 1.022 0.011 98.9 0.001 0.1
1.25 100 2.918 2.804 3.9 0.541 0.006 98.9 0.006 0.6
1.5 100 2.765 2.707 2.1 0.302 0.002 99.4 0.009 0.9
1.75 100 2.707 2.684 0.9 0.165 0.001 99.7 0.007 0.7
2 100 2.684 2.684 0.0 0.016 0.016 0 0 0
  • a. $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged drag coefficient of the elliptical cylinder once the baseline flow has fully developed.
  • b. $\overline{C}_{D,\text{Controlled}}$, $\overline{C}_{L,\text{Controlled}}$, and $\overline{a}$ denote the time-averaged values under stable conditions after AFC is applied, excluding the transient response phase.
  • c. $\overline{a}$ ratio (%) represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the inlet.
In Tab. 2, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the $C_{D}$ reduction rates are 0.9%, 2.1%, 3.9%, 8.0%, and 15.7% respectively, while the $C_{L}$ reduction rates exceed 95% in each case. These results highlight the efficacy of the DRL-based AFC technology in effectively suppressing both $C_{D}$ and $C_{L}$ within the $Ar$ range of 1.75 to 0.75, demonstrating the efficiency of the DRL-based intelligent control strategy. However, as $Ar$ decreases to 0.5, 0.25, 0.1, and 0, even though the $C_{D}$ reduction rates significantly increase to 26.9%, 35.9%, 37.6%, and 43.6%, the $C_{L}$ reduction rates are 68.0%, 66.1%, 50.2%, and 60.0% respectively. This indicates that while the DRL-based control strategy achieves significant reductions in $C_{D}$, its effect on suppressing $C_{L}$ is limited and does not completely eliminate $C_{L}$ around the elliptical cylinder. As the $Ar$ further decreases, the AFC technique based on the DRL algorithm does not fully suppress vortex shedding behind the elliptical cylinder, but merely delays the shedding process. The reduction in $C_{L}$ coefficients aligns with the flow snapshots presented in Fig. 8, showing that vortex shedding still occurs, albeit with controlled frequency and intensity. These phenomena suggest that at very low $Ar$ values, the control strategy faces challenges in managing more complex and unstable flow conditions.
[…]
The control strategies derived from DRL-based AFC technology exhibit greater complexity compared to conventional harmonic forcing implemented in previous works (Bergmann, Cordier, and BrancherBergmann, Cordier, and Brancher (2005)), highlighting the value of employing ANN as controllers. Typically, approximately 93% of the drag induced by the shedding of vortices is suppressed by the discovered control laws, with minimal jet intensities required for drag reduction. For elliptical cylinders with aspect ratios $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.
III.3 Analysis of the Controlled Flow
Fig. 13 illustrates the suppression of vortex shedding in the wake of elliptical cylinders under the influence of active flow control based on DRL. The elliptical cylinders examined have $Ar$ of 1.75, 1.5, 1.25, 1.0, and 0.75, scenarios in which vortex shedding can be completely suppressed. At the initial moment of control (T1), vortices in the wake of the elliptical cylinders are fully developed. During this phase, vortices are observed detaching from the cylinder at a consistent frequency and intensity, forming a Kármán vortex street in the downstream flow. As the control process advances (T2), external excitation methods, such as synthetic jets, are introduced to suppress the vortices in the wake of the elliptical cylinders. The objective of the control is to alter the characteristics of the flow field to weaken or eliminate the formation and persistence of vortices. With the progression of the control measures, a snapshot of the flow field is captured at the conclusion of the control (T3). At this juncture, vortices within the wake of the elliptical cylinders are observed to be suppressed, demonstrating the effectiveness of the DRL-based active flow control strategy in stabilizing complex fluid dynamics.
(a) $Ar=2$
(b) $Ar=1.75$
(c) $Ar=1.5$
(d) $Ar=1.25$
(e) $Ar=1$
(f) $Ar=0.75$
Figure 13: When $Ar=2$, 1.75, 1.5, 1.25, 1, 0.75, the vortex shedding of the elliptical cylinder can be completely suppressed. The streamwise snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as T1, T2, and T3. These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the conclusion of control, respectively.
In the study of elliptical cylinders, the phenomenon of vortex shedding and its control are crucial for understanding the structure of flow fields. For elliptical cylinders with different $Ar$, the behavior of the flow fields shows significant differences. At $Ar=2$, the instantaneous vorticity maps at times T1, T2, and T3 demonstrate a high degree of stability in the flow field, with no evident vortex shedding observed. For elliptical cylinders with $Ar$ values of 1.75, 1.5, and 1.25, at T1, alternating vortices form behind the cylinder, influenced by the elliptical shape effect, which results in lower intensity and frequency of vortex shedding. By T2, with the initiation of control, synthetic jets quickly intervene in the regular vortex shedding process, leading to a delay in vortex detachment and the formation of "elongated vortex structures." The original vortices are significantly elongated and narrowed, and the instantaneous recirculation bubbles are also elongated longitudinally and tend towards stability. By T3, the original vortex shedding has been completely suppressed, with no new vortices forming, and the instantaneous recirculation bubbles reach their maximum extent and remain stable, demonstrating the effectiveness of the control strategy.
At $Ar=1$ and $0.75$, during the T1 phase, a standard Bénard–von Kármán vortex street forms behind the elliptical cylinder, characterized by a series of regularly spaced vortices. These vortices alternately shed from both sides of the cylinder, creating a regular pattern of vortex arrangement. Moving into the second phase (T2), similar to cases with higher aspect ratios ($Ar=1.25$, 1.5, 1.75), the recirculation bubbles behind the cylinder are elongated longitudinally, and the originally shed vortices are also elongated and narrowed, with shearing interactions occurring between them. By the third time point (T3), the vortices behind the elliptical cylinder are completely suppressed, and the regions of positive and negative vorticity in the flow field nearly achieve symmetry, demonstrating high stability.
(a) $Ar=0.5$
(b) $Ar=0.25$
(c) $Ar=0.1$
(d) $Ar=0$
Figure 14: When $Ar=0.5$, 0.25,0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed. The streamwise snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as T1, T2, and T3. These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.
Fig. 14 delineates scenarios where the suppression of vortex shedding from elliptical cylinders is not entirely effective for aspect ratios $Ar$ of 0.5, 0.25, 0.1, and 0. At T1, the vortex shedding behind the cylinder is fully developed, T2 represents the control phase, and T3 provides a snapshot of the flow field after control cessation. For $Ar=0.5$ and $Ar=0.25$, During the T1 phase, vortices are observed detaching from the cylinder with certain frequency and intensity, forming a Kármán vortex street in the downstream wake. In the T2 phase, as control commences, the recirculation bubble is noted to elongate longitudinally to some extent, with the recirculation area appearing relatively symmetric and stable. The intervention of synthetic jets delays the detachment of vortices, which are stretched and deformed due to shear with the side walls. By T3, the initially generated vortices are significantly elongated and deformed, exiting the computational domain, yet the recirculation bubble behind the cylinder remains unstable and continues to oscillate, leading to new vortex shedding. The emergence of new irregular vortex shedding highlights persistent instability behind the elliptical cylinder. For elliptical cylinders with $Ar=0.5$ and $Ar=0.25$, the DRL-based control strategy does not completely suppress vortex shedding. The instability of the wake field is mitigated, demonstrating partial efficacy of the control strategy in manipulating flow dynamics, although full stability is not achieved.
For $Ar=0.1$ and $Ar=0$, during T1, the vortex shedding frequency and intensity from the cylinder are more severe compared to $Ar=0.5$, with enhanced flow instability. Multi-level vortex shedding occurs, with primary and secondary vortices differing in shape, frequency, and intensity. At T2, vortices behind the cylinder are elongated and delayed, and the recirculation area significantly enlarges. The instability of the recirculation area manifests as asymmetry, swinging, and oscillation phenomena. By T3, although the original shedding vortices are elongated and expelled from the computational domain, new, unstable, and irregular shedding occurs. Shed vortices swinging to the side walls are sheared, stretched, and undergo deformation, resulting in irregularly shaped vortices. This highlights the persistent challenges in managing flow dynamics and instability at lower aspect ratios.
IV Conclusions
In this study, we investigate the use of DRL-based AFC to suppress vortex shedding, drag, and lift coefficients of an elliptical cylinder with a Reynolds number of 100. The $Ar$ of the elliptical cylinder ranges from an ellipse ($Ar=2.0$) to a circle ($Ar=1.0$), and finally to a flat plate ($Ar=0$). We employ the PPO algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of the elliptical cylinder. The primary objective of this work is to gain further insights into the control capabilities of deep reinforcement learning algorithms for complex flow systems. The main findings of this study can be summarized as follows:
  • We evaluated the impact of $Ar$ on flow control around elliptical cylinders. When the $Ar$ is 2, the baseline flow is highly stable, and AFC shows negligible effect. However, for aspect ratios of 1.75, 1.5, 1.25, 1, and 0.75, the reduction rates in the drag coefficient around the cylinder are respectively 0.9%, 2.1%, 3.9%, 8.0%, and 15.7%, while the reduction rates in lift coefficient are 99.7%, 99.4%, 98.9%, 98.9%, and 95.2%. These data indicate that the AFC strategy based on DRL is capable of significantly minimizing the cyclically fluctuating drag coefficient to minimal values and stabilizing it, while also maintaining the cyclically oscillating lift coefficient near zero. This demonstrates the efficacy of DRL in AFC applications, effectively reducing the lift and drag around elliptical cylinder with $Ar$ ranging from 0.75 to 1.75. Velocity cloud diagrams reveal that the original periodically shedding vortices transition to a stable flow field with no vortex shedding after flow control. Instantaneous vorticity maps during the control process show that with the activation of synthetic jets, the vortices shedding downstream of the elliptical cylinder are progressively elongated and squeezed out of the computational domain. The originally unstable recirculation area behind the cylinder also transforms into a stable, symmetrical recirculation bubble, ceasing the production of new vortices. These results underscore the robust adaptability of DRL-based AFC across different elliptical cylinder shapes ($Ar$ between 0.75 and 1.75), capable of learning optimal control strategies directly through interaction with the environment to achieve a reduction in drag and lift coefficients around the cylinder. Moreover, for elliptical cylinders with aspect ratios of 1.75, 1.5, 1.25, 1, and 0.75, the mass flow rate ratios of synthetic jets relative to the inlet flow rate are 0.7%, 0.9%, 0.6%, 0.1%, and 1.0% respectively. This underscores that the flow control strategy efficiently utilizes less than 1% of external excitation energy, exemplifying its energy-saving capability and fully satisfying the low-energy consumption requirements essential for effective flow control.
  • When the $Ar$ of the elliptical cylinder decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the drag coefficient and lift coefficient are 26.9%, 35.9%, 37.6%, and 43.6%, respectively, while the corresponding reduction rates for the lift coefficient are 68.0%, 66.1%, 50.2%, and 60.1%. Despite the effective reduction in both drag and lift coefficients achieved by the DRL-based AFC strategy, they have not yet stabilized compared to cases with larger $Ar$ values. The initial phase of control exhibits distinct fluctuations, followed by low-frequency and low-amplitude oscillations in the drag and lift coefficients, which are lower than the baseline. Particularly, for $Ar$ values of 0.1 and 1, the drag and lift coefficients experience a sharp decrease initially, followed by low-amplitude, low-frequency periodic oscillations. This indicates that for elliptical cylinders with $Ar$ ranging from 0.5 to 0, the AFC control strategy can partially suppress lift and reduce drag. Correspondingly, for elliptical cylinders with $Ar$ values of 0.5, 0.25, 0.1, and 0, the ratios of the mass flow rates of the synthetic jets relative to the incoming flow rate are 4.9%, 12.6%, 49.9%, and 52.1%, respectively. This suggests that despite the use of higher external energy, the lift coefficient cannot be completely controlled to prevent oscillations and fluctuations. From the vortex plots of the control process, for $Ar$ values of 0.5 and 0.25, the recirculation region behind the elliptical cylinder is elongated during the initial phase of control, with symmetric recirculation bubbles. As the control progresses, the recirculation region becomes unstable, exhibiting lateral oscillations and the shedding of long, narrow vortices. Similarly, for $Ar$ values of 0.5 and 0.25, the recirculation region undergoes elongation during the control phase, with high-frequency and large-amplitude oscillations, accompanied by the shedding of new vortices. After control cessation, the shedding of vortices is somewhat alleviated, with a decrease in both the frequency and intensity of vortex shedding.
In conclusion, for elliptical cylinders with $Ar$ ranging from 1.75 to 0.75, intelligent agents trained using the PPO algorithm have adeptly learned to adopt an energy-efficient flow control strategy. The control policies based on the PPO algorithm significantly reduce lift around the elliptical cylinder, greatly diminish drag, and suppress vortex shedding in the wake. Regarding elliptical cylinder with $Ar$ values ranging from 0.5 to 0, the control strategies obtained from agents trained with the PPO algorithm can moderately alleviate vortex shedding and reduce the amplitude and frequency of fluctuations in lift and drag coefficients. These control outcomes demonstrate several unique capabilities of AFC strategies based on DRL, making them particularly suitable for managing complex fluid dynamics systems. These capabilities include adaptability, real-time optimization, robustness to uncertainties, the ability to handle complex problems, minimal human intervention, and high control efficiency.
Acknowledgements.
The authors would like to express their gratitude to Dr. Jean Rabault (University of Oslo, Oslo, Norway), Dr. Jichao Li (National University of Singapore, Singapore), and Mr. Qiulei Wang (The University of Hong Kong, Hong Kong SAR, China) for making their open-source codes for deep reinforcement learning and numerical simulation available online. These resources have significantly contributed to the research presented in this paper. The relevant code repositories can be accessed at https://github.com/jerabaul29/Cylinder2DFlowControlDRLParallel Rabault and Kuhnle (2019b); Rabault et al. (2019), https://github.com/npuljc/RL_control_Nek5000 Li (2021); Li and Zhang (2022), and https://github.com/venturi123/DRLinFluids Qiulei Wang (2022); Wang et al. (2022b).
Declaration of interests
The authors report no conflict of interest.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Author ORCIDs

### v2

@@
 Abstract
-This study investigates the effectiveness of Active Flow Control (AFC) technology supported by Deep Reinforcement Learning (DRL) in suppressing vortex shedding, reducing drag, and minimizing lift coefficients around elliptical cylinders.
-The aspect ratio ($Ar$) of the elliptical cylinder transitions from an ellipsoid ($Ar=2.0$) to a circular shape ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$).
+This study investigates the effectiveness of active flow control (AFC) technology supported by deep reinforcement learning (DRL) applied to flows around elliptical cylinders at $Re=100$.
+We vary the aspect ratio ($Ar$) of the elliptical cylinder from an ellipsoid ($Ar=2.0$) to a circular shape ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$).
 We utilize the proximal policy optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on both the upper and lower surfaces of the cylinder.
-The focus is on optimizing fluid dynamics and examining the scalability of these control techniques across different geometric configurations.
-Our research findings indicate that, for elliptical cylinders with aspect ratios between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%.
+The control objective focuses on reducing drag, minimizing lift and suppressing vortex shedding.
+We examine the robustness and adaptability of DRL-based control techniques across different geometric configurations.
+Our research findings indicate that, for elliptical cylinders with $Ar$ between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%.
 Notably, the DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities.
-Additionally, for aspect ratios from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%.
+Additionally, for $Ar$ from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%.
 This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding.
-This underscores the adaptability and potential of DRL-based AFC in managing complex fluid dynamics across diverse elliptical cylinder configurations.
+Overall, our findings underscore the adaptability and potential of DRL-based AFC in controlling complex fluid dynamics across diverse geometric configurations.
 preprint: AIP/123-QED
 I INTRODUCTION
-The development of active flow control technology has progressed from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineeringCollis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021).
-This technology aims to improve flow characteristics through active intervention in the flow field, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiencyCollis and Joslin (2004); Wang et al.
-(2018).
-Initially, research in active flow control primarily focused on theoretical analysis and simple experiments, exploring how to influence fluid flow with external inputs like jets, vibrations, or magnetic fieldsJahanmiri (2010); Wang et al.
-(2018); Awada, Younes, and Ilinca (2021).
-With the advancement of Computational Fluid Dynamics (CFD) and increased computational power, researchers have been able to gain a deeper understanding of complex flow responses through numerical simulations and develop more effective control strategiesAram et al.
+The development of active flow control (AFC) technology has progressed from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineering.Collis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021) This technology actively intervenes the flow field to improve flow characteristics, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiency.Collis and Joslin (2004); Wang et al.
+(2018) In its early stages, AFC research primarily focused on theoretical analysis and simple experiments to explore the influence of external inputs, such as jets, vibrations, or magnetic fields, on fluid flow dynamics.Jahanmiri (2010); Wang et al.
+(2018); Awada, Younes, and Ilinca (2021) With the advancement of computational fluid dynamics (CFD) and increased computational power, researchers have been able to delve deeper into complex flow responses through numerical simulations and develop more effective control strategies.Aram et al.
 (2018); Jain, Yeo, and Chopra (2010); Li et al.
-(2022).
-In recent years, breakthroughs in machine learning and deep learning technologies have led to revolutionary progress in the field of active flow controlBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020).
-Researchers have begun using intelligent algorithms, such as Deep Reinforcement Learning (DRL), to automatically optimize control strategies, achieving precise control over highly nonlinear and variable flow fieldsBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al.
-(2021).
-These methods have not only improved the accuracy and real-time performance of control effects but also significantly expanded the application scope of active flow control technology, including but not limited to aerospace, automotive, energy, and environmental engineering sectorsVignon, Rabault, and Vinuesa (2023); Rabault et al.
-(2020); Garnier et al.
-(2021); Razdan and Shah (2022).
-DRL is an avant-garde methodology that amalgamates the principles of deep learning with reinforcement learning to optimize decision-making through interaction with an environmentArulkumaran et al.
+(2022) However, as fluid problems are typically characterized by a high degree of nonlinearity and multiscale nature, involving high-dimensional state spaces and complex system dynamicsPozorski and Wacławczyk (2020); Reichstein et al.
+(2019), many engineering problems are often intractable using existing models.
+Sophisticated control strategies that require consideration of a wide range of spatiotemporal scales are indispensable when addressing phenomena such as unstable flows, turbulence, and multiphase flows.Subramaniam (2020)
+Recently, the field of AFC has witnessed revolutionary progress, thanks to breakthroughs in machine learning and deep learning technologies.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020) Deep reinforcement learning (DRL) is an avant-garde technique that combines the principles of deep learning with reinforcement learning to optimize decision-making through interactions with an environment.Arulkumaran et al.
 (2017); François-Lavet et al.
 (2018); Wang et al.
-(2022a).
-Deep learning excels in autonomously extracting intricate features from voluminous datasets, while reinforcement learning concentrates on how to make optimal decisions based on these features, guiding the learning process through rewardsJaniesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015).
-DRL leverages deep neural networks to represent the policies or value functions within reinforcement learning, thus empowering machines to tackle high-dimensional state spaces and complex decision-making challenges that were previously intractableArulkumaran et al.
+(2020) Deep learning excels at autonomously extracting intricate features from voluminous datasets, while reinforcement learning concentrates on how to make optimal decisions based on these features by guiding the learning process through rewards.Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015) By utilizing deep neural networks to represent policies or value functions within reinforcement learning, DRL enables machines to tackle complex decision-making challenges and high-dimensional state spaces that were previously considered intractable.Arulkumaran et al.
 (2017); François-Lavet et al.
 (2018); Wang et al.
-(2022a).
-Notable for its robust non-linear modeling capabilities and adaptable learning mechanisms, DRL is capable of directly learning complex strategies from raw sensory data without the need for manually engineered featuresDargazany (2021).
-Moreover, DRL is inherently self-improving, continuously refining its strategies through ongoing interactions with the environmentWang et al.
-(2022a).
-This capacity for self-learning and adaptation bestows a significant advantage upon DRL, especially in resolving decision-making problems within dynamic settingsFrançois-Lavet et al.
+(2020) Renowned for its robust non-linear modeling capabilities and adaptive learning mechanisms, DRL has the capability to directly learn complex strategies from raw sensory data without relying on manually engineered features.Dargazany (2021) Furthermore, DRL is inherently self-improving, continuously refining its strategies through ongoing interactions with the environment.Wang et al.
+(2020) The self-learning and adaptive nature of DRL grants a significant advantage, particularly in resolving decision-making problems in dynamic settings.François-Lavet et al.
 (2018); Wang et al.
-(2022a).
-The applications of DRL are extensive, encompassing domains such as gaming, autonomous vehicles, robotic control, natural language processing, and computer vision.
-In the gaming industry, breakthroughs like AlphaGo’s victory over world champions underscore DRL’s potentialGranter, Beck, and Papke (2017); in autonomous driving, DRL is employed for optimizing route planning and decision-makingSallab et al.
-(2017); in robotics, it enables robots to learn complex physical tasksZhao, Queralta, and Westerlund (2020).
-Additionally, DRL is being applied in scientific research areas like fluid mechanics and industrial optimization, showcasing its vast potential in managing complex systemsBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015).
-In recent years, the application of DRL within the realm of fluid dynamics has seen a marked increase, with a primary focus on active flow control and shape optimizationBrunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al.
-(2021).
-Active flow control technology is a key challenge in the fields of fluid mechanics and engineering, including goals such as drag reduction, prevention of flow separation, control of vortex shedding, and enhancement of mixing efficiencyWang et al.
-(2018).
-Traditional control methodologies often rely on a priori knowledge and empirical rules, whereas DRL offers a novel pathway for autonomously learning optimal control strategies from dataCollis and Joslin (2004); Jahanmiri (2010).
-In the context of drag reduction applications, DRL is utilized to autonomously adjust the operation of control devices (e.g., jets, vibratory mechanisms) to minimize the fluidic drag around objects such as cylindrical bodies or car modelsGarnier et al.
-(2021); Razdan and Shah (2022).
-For vortex shedding control, DRL is capable of learning how to dynamically adjust control parameters to stabilize the flow field and reduce vibrations and noise caused by vortex sheddingVignon, Rabault, and Vinuesa (2023); Rabault et al.
-(2020).
-In the domain of shape optimization, DRL facilitates the automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal design strategiesBhola et al.
-(2023).
-With the continued expansion of computational resources and advancements in algorithms, the application of DRL in fluid dynamics is set to broaden further.
-Fluid mechanics problems are characterized by a high degree of nonlinearity and multiscalar nature, involving high-dimensional state spaces and complex system dynamicsPozorski and Wacławczyk (2020); Reichstein et al.
-(2019).
-These challenges are often intractable for traditional mathematical models, demanding sophisticated control strategies, particularly when addressing phenomena such as unstable flows, turbulence, and multiphase flows, which require consideration of a wide range of spatiotemporal scalesSubramaniam (2020).
-The suitability of DRL methods for these kinds of fluid mechanics issues primarily rests on several key features: Firstly, DRL can handle the high-dimensional and complex state spaces intrinsic to fluid dynamics, effectively capturing the intricate dynamics presentXie et al.
-(2023).
-Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid mechanics systems with numerous variables and parameters, discerning patterns and regularities hidden within dataZhao et al.
-(2024).
-Secondly, DRL’s interactive and adaptive learning approach allows it to self-tune and optimize strategies in response to dynamic and uncertain fluid environments.
-Traditional control methods often struggle with such dynamicsBhansali et al.
-(2023); however, DRL models, through their interaction with the environment and trial-and-error processes, can independently learn and refine strategies to meet diverse fluidic conditions and task requirementsHusen, Chaudary, and Ahmad (2022).
-Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies.
-Control and optimization in fluid mechanics often necessitate foresight into long-term effectsVinuesa et al.
-(2022).
-DRL’s focus on long-term rewards enables models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanicsViquerat et al.
-(2022); Vinuesa (2024).
-In conclusion, the attributes of DRL render it highly suitable for addressing the control and optimization challenges in fluid mechanics.
-Rabault et al.Rabault et al.
-(2019) pioneered the application of advanced DRL algorithms in active flow control.
-The author uses the DRL agent to control the mass flow rate of the synthetic jets on the upper and lower sides of the cylinder, effectively reducing the resistance around the cylinder by 8% and mitigating the detachment of the rear vortex.
-Rabault et al.; Tang et al.; Heess et al.; Ren, Rabault, and Tang; Jia and XuRabault et al.
-(2019); Tang et al.
-(2020); Heess et al.
-(2017); Ren, Rabault, and Tang (2021); Jia and Xu (2024a) applied the Proximal Policy Optimization (PPO) algorithm to active flow control (AFC) in the flow around a cylinder.
-The training results consistently showed a reduction in drag of 8% around the cylinder.
-Li and ZhangLi and Zhang (2022) investigated the flow stability and sensitivity in confined cylinder wakes under various blockage ratios.
-They used the analysis of flow stability and structure sensitivity to guide the manipulation of the flow field by DRL agents.
-Wang et al.Wang et al.
-(2022b) utilized an AFC method based on the PPO algorithm to reduce the drag on a cylindrical body by approximately 8%, and applied the Soft Actor-Critic (SAC) algorithm to decrease the drag around a square cylinder by about 13.6%.
-Jia and XuJia and Xu (2024b) employed the SAC algorithm within a DRL framework to achieve AFC around a square cylinder, effectively modulating the flow dynamics across varying Reynolds numbers.
-Wang et al.Wang et al.
-(2022c) used DRL methods for active control of flow through elliptical cylinders.
-Their training results prove that DRL can successfully learn active control strategies under four angles of attack and various long and short axis ratios.
-Fan et al.Fan et al.
-(2020) was the first to apply DRL to active flow control experiments, optimizing power efficiency by adjusting the rotational speed of two small cylinders behind the main cylinder.
-Ren, Wang, and TangRen, Wang, and Tang (2021) proposed a novel active flow control strategy to hide the hydrodynamic characteristics of bluff bodies, including strong shear forces and periodic shedding vortices, to avoid predators.
-The application of AFC strategies based on DRL has been widely validated in various bluff body flow scenarios for their effectiveness in controlling vortex shedding and reducing drag.
-Among these, studies on flow around constrained cylinders under different blockage ratios represent the most thoroughly investigated areaRabault et al.
+(2020) The applications of DRL are vast and encompass various domains, including gaming, autonomous vehicles, robotic control, natural language processing, and computer vision.
+In the gaming industry, distinguished achievements like AlphaGo’s victory over world champions highlight the potential of DRL.Granter, Beck, and Papke (2017) In the realm of autonomous driving, DRL is employed to optimize route planning and decision-making.Sallab et al.
+(2017) In the field of robotics, it enables robots to learn complex physical tasks.Zhao, Queralta, and Westerlund (2020) Overall, DRL offers the potential to develop adaptive and intelligent control strategies that can adapt to changing environments and optimize performance in a wide range of applications.
+With progressive advancement in the DRL technology, there has been a significant increase in the application of DRL in the field of fluid dynamics as well.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al.
+(2021); Razdan and Shah (2022) DRL methods are highly suited to handle the high-dimensional and complex state spaces intrinsic to fluid dynamics.Xie et al.
+(2023) Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid systems and discerning flow patterns and structures.Zhao et al.
+(2024) DRL models, through their interaction with the environment and trial-and-error processes, can independently learn and refine strategies to meet diverse fluid conditions.Husen, Chaudary, and Ahmad (2022); Collis and Joslin (2004); Jahanmiri (2010) For instance, DRL has been applied to shape optimization of fluid systems and enabled automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal strategies.Bhola et al.
+(2023) Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies.
+This unique feature has allowed DRL to shine in many AFC problems where DRL’s focus on long-term rewards enables control models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanics.Vinuesa et al.
+(2022); Vignon, Rabault, and Vinuesa (2023); Viquerat et al.
+(2022); Vinuesa (2024)
+In the academic literature, flow control around blunt bodies using DRL has attracted considerable interest due to its pivotal role in reducing drag, a key aspect of fluid dynamics research.
+This pursuit carries several important implications, including enhanced energy efficiency, improved performance, noise and vibration reduction, and industrial process optimization.
+As flow around blunt body tends to involve strong non-linearities such as flow separation and recirculation, traditional control methods often struggle in achieving control objectives whereas DRL models have shown promising results.
+Pioneered by Rabault et al.
+Rabault et al.
+(2019), various researchers such as Tang et al.
+Tang et al.
+(2020), Heess et al.
+Heess et al.
+(2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024) applied the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder.
+Consistently, they achieved an approximate 8% reduction in drag in their validation experiments.
+Furthermore, Li and ZhangLi and Zhang (2022) ,Wang et al.Wang et al.
+(2022a), Jia and XuJia and Xu (2024), and Wang et al.Wang et al.
+(2022b) respectively conducted flow control studies on confined circular cylinders and square cylinders using DRL.
+In addition, Fan et al.Fan et al.
+(2020) initiated the application of DRL to AFC experiments, optimizing power efficiency by adjusting the rotational speed of auxiliary circular cylinders positioned behind another circular cylinder.
+Ren, Wang, and TangRen, Wang, and Tang (2021) proposed an AFC strategy to conceal the hydrodynamic characteristics of circular cylinders, such as strong shear forces and periodic shedding vortices.
+Studies on flow around constrained circular cylinders under different blockage ratios have also received extensive attention in the field Rabault et al.
 (2019); Tang et al.
 (2020); Li and Zhang (2022); Wang et al.
-(2022b); Jia and Xu (2024a).
-By contrast, research related to flow around elliptic cylinders has seen relatively less exploration within this domain.
-Due to their unique shape, elliptic cylinders influence the development of separation points and boundary layers differently during flow, potentially leading to Hopf bifurcation and periodic vortex shedding at lower Reynolds numbers.
-These flow characteristics offer opportunities in aerospace, marine engineering, energy engineering, and modern architectural design for improving lift-to-drag ratios, reducing vortex-induced vibration, enhancing energy efficiency, and increasing structural stability.
-Therefore, an in-depth investigation of the flow characteristics around elliptic cylinders not only enhances our understanding of fluid dynamics but also promotes performance optimization and innovation in engineering design.
-In summary,despite significant advancements in controlling flow around circular cylinders through DRL algorithms, the study of flow around elliptical cylinders presents a vast research landscape due to its more complex flow characteristics and control challenges.
-AFC based on DRL algorithms introduces a novel methodology that leverages learning from flow field information to intelligently adjust control strategies, aiming at reducing drag and suppressing vortex shedding oscillations.
-This research is dedicated to examining the impact of various $Ar$ of elliptical cylinders on flow stability and sensitivity to structural parameters.
-Building upon this, it explores the application of DRL-based AFC in managing flow around elliptical cylinders by utilizing the derived physical insights.
-By establishing a robust reinforcement learning environment and defining clear reward mechanisms, this study aims to identify and implement effective flow control strategies.
-The aim of this study is to enhance the flow stability around an elliptical cylinder using AFC technology based on DRL algorithm, serving as a reference for flow control around similarly complex-shaped objects.
-The structure of this paper is organized as follows: Section II introduces the problem description and provides a comprehensive overview of Deep DRL and its framework for application in AFC.
-Section Section III presents the main results and analysis.
-Finally, Section IV concludes the paper by summarizing the key findings and discussing their implications.
+(2022a); Wang and Xu (2024).
+The existing studies have successfully demonstrated the effectiveness of using DRL for flow control, particularly in actively manipulating the flow around circular cylinders and square cylinders.
+Although active flow control strategies for both circular and square cylinders are pivotal in various engineering applications for managing aerodynamic and hydrodynamic forcesVignon, Rabault, and Vinuesa (2023); Rabault et al.
+(2020); Garnier et al.
+(2021); Razdan and Shah (2022), their extension to more intricate, non-ideal geometries poses significant challenges.
+The effectiveness of DRL strategies in such contexts remains largely unexplored, marked by uncertainties regarding their capacity to generate physically reasonable and adaptable control measures.
+Current literature indicates a distinct lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies.
+This research proposes to bridge these gaps by focusing on the flow dynamics around parameterized ellipsoids, examining how variations in the aspect ratio influence flow fields and their corresponding DRL-driven control strategies.
+By extending the investigation from elongated ellipsoids to circular cylinders and subsequently normal flat plates, this study aims to develop a deeper understanding of the physical validity of DRL approaches and their adaptability across varying body geometries.
+This work extends the studies conducted by Wang and XuWang and Xu (2024); Jia and Xu (2024) by applying the DRL-based AFC approach to elliptical cylinders, which present more complex flow characteristics and additional control challenges.
+The objective is to develop effective control strategies to reduce drag, improve flow stability, and enhance overall performance.
+The structure of this paper is organized as follows: Sec.
+II introduces the problem description and provides a comprehensive overview of DRL and its framework for application in AFC.
+Sec.
+III presents the main results and analysis.
+Finally, Sec.
+IV concludes the paper by summarizing the key findings and discussing their implications.
 II PROBLEM DESCRIPTION AND METHODOLOGY
@@
 Model configuration
-The present study investigates the flow around two-dimensional elliptical cylinders with their major axis oriented perpendicular to the flow direction, as depicted in Fig.
-1(a).
-The major axis of the constrained elliptical cylinder (denoted 2a) remains fixed, while the shape of the elliptical cylinder is changed by changing the value of 2b.
-The Reynolds number is calculated based on the length of the major axis, formulated as $Re=U(2a)/\nu$, where $U$ represents the velocity of the incoming uniform flow and $\nu$ denotes the kinematic viscosity.
-The aspect ratio ($Ar$) is defined as the ratio of the semi-minor axis ($b$) to the semi-major axis ($a$), expressed as $Ar=b/a$.
-For ease of comparison with circular cylinders, we define $2a$ as $D$, representing the diameter of an equivalent circular cylinder.
-Consequently, an $Ar$ of 0 designates a flat plate normal to the flow, while an $Ar$ of 1 denotes a circular cylinder.
-This study investigates a series of variations ranging from an elliptical cylinder $(Ar=2.0)$ to a circular cylinder $(Ar=1.0)$, and further to a flat plate $(Ar=0)$, as shown in Fig.
-1(c).
-Figure 1: (a) Definition of the problem, including relevant dimensions.
-(b) Establishment of boundary conditions.
-(c) This study examines an elliptical cylinder with varying aspect ratios: a normal flat plate corresponding to a cylinder with an aspect ratio of 0, a special case of a cylinder with an aspect ratio of 1, and an elliptical cylinder with an aspect ratio of 2.
-Computational domain and coordinate system
-An elliptical cylinder is placed within a computational domain bounded by a rectangular perimeter, where the diameter of the elliptical cylinder is defined as D, and the major axis length (2a) of the elliptical column is equivalent to D.
-The computational domain extends from -2D in front of the cylinder or elliptical column to 20D behind it in the flow direction, which is the direction of fluid flow.
-In the cross-flow direction, perpendicular to the flow direction, the domain ranges from -2D to 2.1D.
-Within this computational domain, a Cartesian coordinate system is established with the center of the cylinder or elliptical column as the origin O, where the direction of flow is defined along the x-axis, and the axis perpendicular to the flow direction is the y-axis.
-The configuration of the synthetic jets is depicted in Fig.
-1(b), specifically positioned at two directions on the cylindrical or elliptical column, $\theta_{1}=90^{\circ}$ and $\theta_{2}=270^{\circ}$, representing symmetrical locations of the jet on either side of the object’s surface.
-The width of the jet opening is fixed at $\omega=10^{\circ}$.
+[…]
+Figure 1: (a) Definition of the problem, including the dimensions of the computational domain and the establishment of a Cartesian coordinate system.
+(b) Establishment of boundary conditions; note that the dimensions of the elliptical cylinder and the computational domain in the figure are not to scale, focusing solely on the delineation of boundary conditions.
+(c) This study investigates elliptical cylinders with different aspect ratio: a normal flat plate corresponding to $Ar=0$, a circular cylinder with $Ar=1$, and an elliptical cylinder with $Ar=2$.
+The dimension $2a$ is fixed as $D$, and the $Ar$ is varied by changing the length of 2b.
 Governing equations
 Let $\Omega\subset\mathbb{R}^{nd}$ and $(0,T)$ be the spatial and temporal domains, respectively, where $nd$ is the number of space dimensions, and let $\Gamma$ denote the boundary of $\Omega$.
-The spatial and temporal coordinates are denoted by $x$ and $t$.
-This work investigates the dynamics of incompressible viscous fluid flow.
-The Navier–Stokes equations governing incompressible fluid flow are
+The Navier-Stokes equations that describe the dynamics of this incompressible viscous fluid flow are given by:
 $\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot(\nabla\mathbf{u})=-\nabla p+Re^{-1}\Delta\mathbf{u}\quad\text{on}\quad\Omega\times(0,T),$ (1a)
-$\nabla\cdot\mathbf{u}=0\quad\text{on}\quad\Omega\times(0,T).$ (1b)
-Here, $\bm{u}=(U,V)$ represents the velocity vector, where $U$ and $V$ are the velocity components along the $x$- and $y$-axes, respectively.
-$p$ is the thermodynamic pressure.
-The $Re$ is calculated based on the length of the major axis, formulated as $Re=U(2a)/\nu$, where $U$ represents the velocity of the incoming uniform flow and $\nu$ denotes the kinematic viscosity.
+$\nabla\cdot\mathbf{u}=0\quad\text{on}\quad\Omega\times(0,T),$ (1b)
+where $\mathbf{u}$, $t$ and $p$ are the velocity vector, time, and pressure, respectively. In this study, the $Re$ is set to 100.
 Boundary conditions
-The upstream boundary of the computational domain, designated as $\Gamma_{\text{in}}$, is characterized by a parabolic velocity profile oriented in the $x$-axis, where the boundary conditions are specified such that the streamwise velocity component $U$ is equal to the mean velocity $U$ with a value of 1, and the transverse velocity component $V$ is set to zero.
+The upstream boundary of the computational domain, designated as $\Gamma_{\text{in}}$, is characterized by a parabolic velocity profile oriented in the $x$-axis, where the boundary conditions are specified such that the transverse velocity component is set to zero, and the velocity along the stream direction shows a parabolic profile distribution.
 On the downstream boundary, denoted as $\Gamma_{\text{out}}$, a Neumann-type boundary condition is prescribed for the velocity field, ensuring that the stress vector is nullified.
@@
 The design of the synthetic jets ensures mass flow balance, as expressed by the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, to maintain mass conservation in the system.
-Except for the jet, the non-slip solid wall boundary condition $\Gamma_{\text{w}}$ is adopted for the rest of the cylinder.
+Except for the synthetic jets, the non-slip solid wall boundary condition $\Gamma_{\text{w}}$ is adopted for the rest of the cylinder.
 In compliance with the above requirements, the system is considered with the following boundary conditions
@@
 Here, $U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain.
+The mean velocity magnitude is $\overline{U}=2U_{m}/3=1$ and is used to define the Reynolds number $Re=\overline{U}D/\nu$.
 The functions $f_{Q_{i}}$ represent radial velocity profiles that simulate the suction or injection mechanisms of the jets.
@@
 Specifically, $f_{Q_{i}}$ is defined as $A(\theta;Q_{i})(x,y)$, with the modulation relying on the angular coordinate $\theta$ as illustrated in Fig.
-2(b).
+2.
 For the jet characterized by a width of $\omega$ and centered at $\theta_{0}$, which is positioned on the cylinder with radius $R$, the modulation is established as follows:
@@
 Grid system
-Taking the computational domain of an elliptical cylinder with an $Ar$ of 1 as an exemplar, the domain undergoes discretization utilizing a synergistic approach that incorporates both structured (quadrilateral) and unstructured (triangular) meshes.
+Considering an elliptical cylinder computational domain with $Ar=1$, the domain is discretized employing a hybrid meshing approach that integrates both structured (quadrilateral) and unstructured (triangular) meshes.
 As delineated in Fig.
-2, this process results in the domain being segmented into 18,484 mesh elements.
-Herein, the vicinity of the elliptical cylinder is discretized with multiple layers of quadrilateral meshes, whereas the remaining regions are discretized using triangular meshes.
-Fig.
-2(a) reveals the global discretization scheme of the computational domain, highlighting the comprehensive approach to mesh division.
-Fig.
-2(b) offers insight into the mesh partitioning adjacent to the elliptical cylinder, with the multi-layered quadrilateral mesh around the cylinder being significantly finer compared to the triangular mesh.
-Fig.
-2(c) illustrates the multi-layered quadrilateral mesh division around the elliptical cylinder and details the integration strategy between structured and unstructured meshes.
-Lastly, Fig.
-2(d) provides a closer inspection of the mesh division surrounding the elliptical cylinder, emphasizing the precision of the discretization near the structure.
-The mesh division method for ellipses with $Ar$ other than 1 follows the same approach as that for $Ar=1$.
-Figure 2: Numerical simulation of water flowing through an elliptical cylinder, close-up view of the mesh at $Ar=1$.
-The computational domain is discretized into 18,484 mesh elements, incorporating both triangular and quadrilateral meshes.
+2, the computational domain is divided into 18,484 mesh cells.
+Specifically, the area surrounding the elliptical cylinder is discretized using a multi-layer quadrilateral mesh, while the remaining regions are discretized with triangular meshes.
+Fig.2(a) displays the global discretization scheme for the computational domain, highlighting the comprehensive approach to mesh division.
+Fig.2(b) provides a detailed view of the mesh partitioning near the elliptical cylinder, where the multi-layered quadrilateral meshes surrounding the cylinder are notably finer compared to the triangular meshes.
+Fig.2(c) illustrates the division of multi-layered quadrilateral meshes around the elliptical cylinder, emphasizing the precision of discretization near the structure.
+Finally, fig.2(d) showcases the mesh division around the elliptical cylinder and the integration strategy between quadrilateral and triangular meshes.
+The method for mesh division for ellipses with $Ar$ values other than 1 follows the same approach as that for $Ar=1$, and is not separately demonstrated.
+Figure 2: For $Ar=1$, the computational domain is discretized into 18,484 mesh cells, including both triangular and quadrilateral meshes.
 (a) Discretization of the computational domain is viewed from a global perspective.
@@
 The Strouhal number $(St)$ is used to describe the characteristic frequency of oscillatory flow phenomena and is defined as follows:
-$St=\frac{f\cdot D}{U}.$ (6)
-Where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$.
-$D$ is the characteristic length of the elliptical cylinder, and $\overline{U}$ is the mean velocity of the upstream flow.
+$St=\frac{f_{s}\cdot D}{\overline{U}}$ (6)
+Where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$. $D$ is the characteristic length of the elliptical cylinder, and $\overline{U}=1$.
 Grid independence
@@
 On the other hand, the fine mesh, characterized by a high density of grid points, considerably slows down the training speed for reinforcement learning, presenting a challenge for efficient simulation execution.
-Our aim is to perform numerical simulations with the least number of mesh elements without compromising on computational accuracy.
+Our aim is to perform numerical simulations with the least number of mesh cells without compromising on computational accuracy.
 Consequently, the discretization strategy of the main mesh was adopted for subsequent numerical simulations and reinforcement learning training.
@@
 Configuration Mesh resolution Mesh $C_{D,max}$ $C_{D,mean}$ $C_{L,max}$ $St$
+Schäfer et al.Schäfer et al. (1996) - - 3.220–3.240 - 0.990–1.010 0.295–0.305
 Rabault et al.Rabault et al. (2019) - 9262 3.205 -
-Schäfer et al.Schäfer et al. (1996) - - 3.220–3.240 - 0.990–1.010 0.295–0.305
-Coarse 10540 3.242 3.224 1.052 0.304
-$Ar=1$ Medium 18484 3.225 3.205 0.990 0.300
-Fine 25624 3.228 3.207 0.992 0.301
+Coarse 10,540 3.242 3.224 1.052 0.304
+$Ar=1$ Medium 18,484 3.225 3.205 0.990 0.300
+Fine 25,624 3.228 3.207 0.992 0.301
 II.2 Deep Reinforcement learning
@@
 Fig.
-3(a) demonstrates how deep learning captures the representation of input data across multiple layers of a neural network, including the input layer, several hidden layers, and the output layer.
+3 demonstrates how deep learning captures the representation of input data across multiple layers of a neural network, including the input layer, several hidden layers, and the output layer.
 Each layer is composed of numerous neural units, with information transmitted between nodes via connection weights.
@@
 Utilizing the backpropagation algorithm, the network autonomously discerns the transformation process from raw data to sophisticated, high-level abstract features, eliminating the need for manual feature extraction.
-(a) DL
-(b) RL
-(c) DRL
-Figure 3: The basic architecture of deep learning, reinforcement learning, and deep reinforcement learning.
+Figure 3: DRL combines deep learning’s ability to interpret complex data with reinforcement learning’s strategy to make informed decisions.
 Reinforcement learning
 In the domain of Reinforcement Learning, the interaction between an agent and its environment is the core mechanism by which learning is facilitated.
-As illustrated in Figure 3(b), this interaction proceeds in a cyclical pattern.
+As illustrated in Figure 3, this interaction proceeds in a cyclical pattern.
 At each timestep $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the relevant information necessary for making decisions.
@@
 Deep learning addresses the challenge of processing original, high-dimensional input data, which is problematic for traditional reinforcement learning, while reinforcement learning provides a structured decision-making framework and a goal-oriented learning mechanism that enhances the capabilities of deep learning.
-As depicted in Figure 3(c), a deep learning network transforms the intricate state of the environment into a compressed and efficient feature representation.
+As depicted in Figure 3, a deep learning network transforms the intricate state of the environment into a compressed and efficient feature representation.
 This transformation yields a more manageable and interpretable state space for reinforcement learning algorithms.
@@
 These policies are typically parameterized by deep neural networks, implying that the selection of actions is influenced not only by the current observation of the environmental state but also by the learned weights within the network.
-II.3 PPO algorithm
-PPO is an advanced DRL algorithm suitable for solving reinforcement learning problems in discrete and continuous action spaces.
-The core idea is to improve performance by maximizing approximate policy optimization while maintaining the stability of policy updates.
-In PPO, the agent interacts with the environment, collects empirical data, and uses this data to construct an objective function.
-The objective function takes into account dominance estimates from empirical data and is used to measure the direction of policy improvement.
-In the PPO algorithm, the agent interacts with the environment to generate a series of trajectories ($\tau$).
-A trajectory is a sequence of state-action pairs produced by the agent following the current policy interacting with the environment until reaching a terminal state or a predetermined maximum number of steps.
-Symbolically, a trajectory $\tau$ can be represented as:
+II.3 Proximal Policy Optimization algorithm
+PPO is an advanced DRL algorithm suitable for addressing reinforcement learning challenges within both discrete and continuous action spaces.
+The fundamental concept of PPO revolves around enhancing performance by maximizing a surrogate objective function while ensuring the stability of policy updates.
+In PPO, an agent interacts with an environment to gather experience data, which is then utilized to construct the objective function.
+This function incorporates dominant estimates from the experience data to gauge the direction of policy improvement.
+During these interactions, the agent generates a sequence of trajectories ($\tau$).
+A trajectory consists of a series of state-action pairs produced as the agent operates under the current policy, continuing until a terminal state is reached or a predetermined maximum number of steps is exceeded.
+Symbolically, a trajectory $\tau$ can be represented as follows:
 $\tau=(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},s_{2},\ldots,s_{T-1},a_{T-1},r_{T},s_{T}),$ (7)
@@
 The inclusion of this clipped surrogate objective aids in moderating the extent of policy updates, thereby stabilizing the training process.
-In summary, the PPO algorithm presents a robust framework for the training of deep reinforcement learning agents.
+In summary, the PPO algorithm presents a robust framework for the training of DRL agents.
 It operates through an iterative process of updating policy parameters aimed at maximizing the expected cumulative reward.
@@
 II.4 DRL-Enhanced Active Flow Control
-This section outlines the design of a framework utilizing DRL algorithms to address active flow control problems.
+This section outlines the design of a framework utilizing DRL algorithms to address AFC problems.
 Firstly, it is necessary to define two fundamental elements in DRL: the agent and the environment.
@@
 Rabault et al.Rabault et al.
-presented the main structure of the custom environment class, encouraging readers to delve into the detailed content.
-This framework design applies deep reinforcement learning methods to solve the active flow control problem, enabling the agent to continuously learn and optimize its behavior through interaction with the environment to achieve the control goal.
+(2020) presented the main structure of the custom environment class, encouraging readers to delve into the detailed content.
+This framework design applies DRL methods to solve the AFC problem, enabling the agent to continuously learn and optimize its behavior through interaction with the environment to achieve the control goal.
 Fig.
@@
 During the exploration phase, the agent engages with the environment via its current policy $\pi$, executing actions $a_{t}$, receiving states $s_{t}$ and rewards $r_{t}$, and noting terminal states.
-Upon the conclusion of the iteration, the parallel computation across $n$ environments yields a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, with each trajectory $\tau_{i}$ comprising the sequence of states, actions, and rewards as $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1},r_{1}),\ldots,(s_{T},a_{T},r_{T})\}$, where $(s_{t},a_{t},r_{t})$ represents the state, action , and reward at time-step $t$, respectively, for the $i$-th environment.
+Upon the result of the iteration, the parallel computation across $n$ environments yields a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, with each trajectory $\tau_{i}$ comprising the sequence of states, actions, and rewards as $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1},r_{1}),\ldots,(s_{T},a_{T},r_{T})\}$, where $(s_{t},a_{t},r_{t})$ represents the state, action , and reward at time-step $t$, respectively, for the $i$-th environment.
 In the optimization phase, the agent samples a minibatch of trajectories from the experience replay buffer to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$.
@@
 Through iterative refinement, the policy network incrementally enhances its action selection strategy within the environment.
-The formula delineates the update rule applied to the policy network during the optimization phase.
   • State $s_{t}$: Traditional state spaces are often too expansive for effective policy learning, requiring operations in partially observable environments.
@@
 The design of probes placement in this study was based on the design principles articulated by Rabault et al.Rabault et al.
-(2019), which advocate strategic positioning of probes around the cylinder and within the wake region of the cylinder.In their studies, Jia and XuJia and Xu (2024a) and Wang et al.Wang et al.
-(2022b) all referred to the probes arrangement scheme proposed by Rabault et al.Rabault et al.
+(2019), which advocate strategic positioning of probes around the cylinder and within the wake region of the cylinder.In their studies, Wang and XuWang and Xu (2024) and Wang et al.Wang et al.
+(2022a) all referred to the probes arrangement scheme proposed by Rabault et al.Rabault et al.
 (2019).
@@
 The flow rates of the two synthetic jets are constrained such that $Q_{\text{lower}}+Q_{\text{upper}}=0$.
-Therefore, the RL agent manipulates the velocity of the upper synthetic jets, while the lower synthetic jets has the same velocity but in the opposite direction.The velocity of the synthetic jets is set to be less than 2% of the inlet mean velocity.
-This restriction is imposed to prevent the agent from taking shortcuts and using excessively high jet velocities to achieve drag reduction, which contradicts the fundamental principle of active flow control: energy efficiency.
+Therefore, the RL agent manipulates the velocity of the upper synthetic jets, while the lower synthetic jets has the same velocity but in the opposite direction.
+The velocity of the synthetic jets is set to be less than 2% of the inlet mean velocity.
+This restriction is imposed to prevent the agent from taking shortcuts and using excessively high jet velocities to achieve drag reduction, which contradicts the fundamental principle of AFC: energy efficiency.
 Furthermore, each control action lasts for 50 time steps, termed as a viscous action.
@@
 Numerical simulations were performed using the eighth version of the open-source software platform, OpenFOAM® version 8.Jasak et al.
-(2007) The deployment of deep reinforcement learning algorithms was based on the Tensorforce library.Schaarschmidt, Kuhnle, and Fricke (2017) The ANN definitions and gradient descent algorithms were facilitated using the TensorFlow open-source library.Abadi et al.
+(2007) The deployment of DRL algorithms was based on the Tensorforce library.Schaarschmidt et al.
+(2018) The ANN definitions and gradient descent algorithms were facilitated using the TensorFlow open-source library.Abadi et al.
 (2016) For DRL implementation, agent was loaded using the Tensorforce platform with the PPO algorithm, and the environment was constructed using the Gym interface.Brockman et al.
@@
 This parallel strategy was inspired by the research of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al.
-(2022b), and also incorporated insights from the work of Wang et al.Wang et al.
-(2022b) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.
+(2022a), and also incorporated insights from the work of Wang et al.Wang et al.
+(2022a) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.
 III RESULTS
@@
 Vortex shedding phenomenon
-Figure 5: Instantaneous Vorticity Contours for Flow Around an Elliptic Cylinder at $Re=100$, Varying from $Ar=2$ to $Ar=0$.
+Fig.
+5 depicts the vortex shedding patterns within the wake field of elliptical cylinders with $Ar$ from 0 to 2.
+Starting from Fig.
+5(a) depicting an elliptical cylinder with $Ar=2$, the wake field does not exhibit vortex shedding.
+This is expected as the elongated elliptical cylinder results in weak adverse pressure gradients that allow the boundary layers to stay attached for a prolonged distance.
+The zero vorticity contour lines in the wake extend all the way to the outlet with slight fluctuations.
+This suggests that the wake is largely stable, with minor flow instability due to shear in the flow.
+It is worth highlighting that this test case of $Ar=2$ serves as a special verification study for subsequent DRL control experiments.
+For a stable flow field with no vortex shedding, we expect the DRL-based flow control algorithm to conform to the underlying flow physics and to not introduce any actuation since the current baseline is essentially the optimal flow configuration.
+To the best of our knowledge, this sanity test that verifies minimal DRL-based control is often overlooked in literature.
+We include this to showcase the robustness of DRL-based control across a spectrum of flow physics.
+We delay the discussion of DRL control to Sec.
+III.2.
+Figure 5: Contours of instantaneous vorticity for flow around an elliptic cylinder at $Re=100$, varying from $Ar=2$ to $Ar=0$.
 Purple and orange indicate positive and negative vorticity, respectively.
 In all the depicted vorticity contours, purple represents the clockwise vortices shed from the upper side of the elliptical cylinder, while orange represents the counterclockwise vortices shed from the lower side.
-Fig.
-5 presents the vortex shedding patterns in the wake region formed around elliptical cylinders with varying $Ar$.
-The subfigures (a) to (j) depict the vortex structures that have reached a fully developed state, illustrating the progression from an aspect ratio of 0 to 2.
-At a Reynolds number of 100, around elliptical cylinders with $Ar$ varying from 0 to 2, persistent unstable vortices are observed at the computational domain’s outlet, indicating the continuation of fluid instability along the outflow direction.
-However, as the fluid progresses downstream, there is a diminishing trend in the peak vortex intensity.
-This attenuation of vorticity is predominantly attributed to viscous diffusion effects prevalent in low Reynolds number regimes.
-This is consistent with the study by Johnson, Thompson, and HouriganJohnson, Thompson, and Hourigan (2004).
-Between $Ar$ of 2 and 1.25, in addition to the recirculation bubbles downstream of the cylinder, two additional recirculation bubbles form upstream near the wall.
-As the $Ar$ decreases further, the confined flow exhibits only one recirculation bubble.
-Similar to observations by Sahin and OwensSahin and Owens (2004), under the influence of blockage ratio and Reynolds number, two additional recirculation bubbles appear on the solid sidewalls of the computational domain.
-At specific Reynolds numbers, this will lead to another Hopf bifurcation at $Re_{\text{c2}}$.
-For an elliptical cylinder with an $Ar=2$, the wake field does not exhibit vortex shedding, and the zero vorticity contour lines show slight fluctuations.
-This suggests that despite the overall stability of the wake, there is still minor flow instability.
-As the $Ar$ decreases to 1.75 and 1.5, the wake field begins to exhibit more pronounced vortex shedding phenomena.
-The vortices become more compact, and there is increased interaction between positive and negative vortices through shearing, stretching, and detachment processes.
-As the $Ar$ decreases to 1, transforming the cylinder into a perfect circle, the vortex structure becomes more regular, forming a typical von Kármán vortex street.
+As the $Ar$ decreases to 1.75, the recirculation region behind the cylinder starts to become unstable and begins to exhibit vortex shedding phenomena.
+As $Ar$ decreases further to 1.5 and 1.25, the vortices become more compact with increasing frequency, and there is increasing interaction between pairs of vortices through shearing, stretching, and detachment processes.
+It is also evident that the onset of vortex shedding shifts upstream with decreasing $Ar$.
+As the $Ar$ decreases to 1, which transforms the elliptical cylinder into a perfect circle, we observe the typical von Kármán vortex street reported in many literature.
 This "standard aligned BvK street" is characterized by a series of regularly spaced vortices that alternately form on both sides of the cylinder.
@@
 At $Ar=0.1$, for extremely low $Ar$, the recirculation area further decreases, resulting in an increase in both the frequency and intensity of vortex shedding.
-At $Ar=0$, when the cylinder degenerates into a flat plate, particularly in the zero-vortex region behind the ellipse, the acceleration of vortex shedding becomes more pronounced.
+At $Ar=0$, when the cylinder degenerates into a vertical flat plate, flow separates right at the two sides of the flat plate and vortex shedding becomes more pronounced.
 This indicates an increase in flow instability and irregularity in vortex size and shape.
-Fluctuation velocity
-Strykowski and SreenivasanStrykowski and Sreenivasan (1990) demonstrated that the introduction of minor disturbances in a targeted region can alter vortex formation, effectively reducing or controlling vortex shedding.
-Similarly, our study employs small perturbations to disturb the flow field with the aim of decreasing drag and suppressing vortex shedding.
-Furthermore, the judicious incorporation of fluid physics knowledge within the control agents can streamline their architecture without compromising performance, as exemplified by Pastoor et al.Pastoor et al.
+Across all test cases, persistent instabilities in the form of vortices are observed at the exit of the computational domain, indicating that the fluid flow remains unstable in the downstream direction.
+As the fluid advances downstream, there is a tendency for the peak vortex intensity to decrease.
+This attenuation of vorticity is primarily attributed to the viscous diffusion effects commonly present in low Reynolds number regions.
+This finding is consistent with the studies by Johnson, Thompson, and Hourigan Johnson, Thompson, and Hourigan (2004).
+Between $Ar$ of 2 and 1.25, in addition to the recirculation bubble downstream of the cylinder, two additional recirculation bubbles form near the upstream wall.
+As $Ar$ is further reduced, the confined flow exhibits only one recirculation bubble.
+This observation is similar to that of Sahin and Owens Sahin and Owens (2004), where two additional recirculation bubbles appear on the solid sidewalls of the computational domain under the influence of blockage ratio and Reynolds number.
+These phenomena in the wake field of elliptical cylinders represent some intriguing physical behaviors.
+This serves as a foundation for subsequent flow control studies.
+Velocity fluctuation and design of probe layout
+In addition to the baseline flow field, we analyze the fluctuating component of velocity, which serves as a good indicator to inform the location of probes during control.
+The judicious incorporation of fluid physics knowledge within the control agents can streamline their architecture without compromising performance, as exemplified by Pastoor et al.Pastoor et al.
 (2008), where phase control in drag reduction of a bluff body is primarily founded upon the understanding of decoupled shear layer development and wake processes.
-In this study, the strategic placement of probes was thoughtfully engineered to align with the dynamic and controllable aspects of the flow field.
-Particular attention was paid to its inherent instability and pulsatility, ensuring that the probes could effectively capture the critical fluctuations necessary for a comprehensive analysis of the flow dynamics.
-The research conducted by Protas and WesfreidProtas and Wesfreid (2002) explores the dynamics of the total drag coefficient, $C_{D}$, which is defined by the equation $C_{D}=C_{D}^{b}+C_{D}^{0}$.
-Here, $C_{D}^{b}$ represents the base drag coefficient, a static value consistent across specific $Re$, encapsulating the inherent drag associated with the baseline flow.
-Conversely, the correctable component $C_{D}^{0}$ reflects the influence of oscillatory flow phenomena on the drag coefficient, providing an avenue for adjustment through strategic manipulation of Bénard–von Kármán vortices.
-In our study, the instantaneous velocity field is segmented into the mean velocity field and the fluctuating velocity field.
-We hypothesize that probes equipped to detect these pulsatile and oscillatory dynamics can leverage targeted minor disturbances to effectively neutralize these instabilities.
-Such interventions are anticipated to facilitate the suppression of vortex shedding, thereby enhancing the control of flow behaviors.
-Figure 6: For the case of $Ar=1$, the distribution of probes is depicted across various velocity fields: (a) Instantaneous velocity, (b) Fluctuating velocity.
-We delineates the instantaneous and fluctuating velocity fields behind an elliptical cylinder during full vortex development, as illustrated in Fig.
-6.
-The strategic placement of probes is predicated on the objectives of reducing drag and lift around the elliptical cylinder while simultaneously suppressing vortex shedding.
-Probes are positioned around and downstream of the elliptical cylinder, specifically within the vortex shedding region.
-The probes around the elliptical cylinder aim to diminish the drag and lift coefficients integral to our control objectives.
-Downstream, probes are strategically deployed in the wake to facilitate the extraction of flow field information from the regions of maximum fluctuating velocity, where intense turbulent motion typically occurs due to vortex instability and enhanced turbulent energy.
-The probe placement in the cylinder’s wake encompasses the entire recirculation area, which is critical for capturing the most unstable and intensely turbulent information in the wake region.
-This configuration plays a significant role in reducing the fluctuating components of the drag coefficient, suppressing vortex shedding, and stabilizing the wake flow.
+We also utilize the physical information from the wake field to guide the flow control strategies of the agents.
+This aspect is particularly reflected in the practical design of the probe positions.
+The study by Protas and Wesfreid Protas and Wesfreid (2002) elaborates on the adjustable component of the drag coefficient, $C_{D}^{0}$, which reflects the influence of oscillatory flow phenomena on the drag coefficient.
+This provides a pathway for strategic manipulation of the von Kármán vortex street to adjust the drag through targeted interventions.
+Building on this foundation, we tailor the physical information observed by the agents to specifically face the factors contributing to instability in the wake field.
+Figure 6: The distribution of probes within the fluctuating velocity magnitude fields for different $Ar$ : (a) $Ar=0.25$; (b) $Ar=0.1$; (c) $Ar=0$.
+As depicted in Fig.
+6, the fluctuating magnitude velocity fields around elliptical cylinders with $Ar$ of 0, 0.1, and 0.25 are illustrated.
+The positions of the black dots on the diagram represent the locations of the probes.
+These probes are strategically placed around the elliptical cylinder and in the downstream wake region.
+On one hand, the arrangement of probes around the elliptical cylinder is designed to capture physical information pertinent to the lift and drag forces, which are critical components of our control objectives.
+On the other hand, probes within the wake of the elliptical cylinder are positioned to encompass areas of maximum velocity fluctuations, facilitating the extraction of flow field data from regions where velocity fluctuations and vortex instabilities are most intense.
+The extraction of physical information from these locations enables the agents to acutely observe changes in the flow field.
 In Fig.
-6, an ellipse with an $Ar$ of 1 serves as the prototype.
-The methodology for designing probe placements for other $Ar$ adheres to the same principles established for $Ar=1$.
+6, three tests with different $Ar$ are shown as examples.
+The configurations for other $Ar$ are fundamentally similar and are not elaborated further.
 III.2 Effectiveness of AFC
 This section describes the performance of AFC using the DRL algorithm, which manipulates the magnitude of synthetic jets to achieve targeted flow control.
-Within a flow configuration at $Re=100$, AFC is applied to elliptical cylinders with $Ar$ ranging from 2 to 0, as well as to cylindrical and flat plate configurations.
-The evaluation of the control’s effectiveness and performance is based on several criteria: comparisons of the velocity fields before and after control, reductions in drag coefficient, variations in lift coefficient, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
-Figure 7: Snapshots of the velocity magnitude illustrating the effect of the active flow control.
-Left side of each row: baseline simulation.
-Left side of each row: result with active flow control.
-(a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$.
+At $Re=100$, AFC is applied to elliptical cylinders with $Ar$ ranging from 2 to 0, from elongated body to circular cylinder and then to a normal flat plate.
+The evaluation of the control’s effectiveness and performance is based on several criteria: intensity of vortex shedding after control, reductions in drag coefficient, variations in lift coefficient, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
 Fig.
 7 and Fig.
-8 provide flow field snapshots of elliptical cylinders under various $A$, encompassing both baseline (uncontrolled) and controlled scenarios.
-This section primarily describes the flow structures observed in the baseline cases.
-For elliptical cylinders with $Ar=2$, the baseline flow field inherently displays high stability, with no vortex shedding observed.
-Consequently, the AFC did not produce a noticeable effect.
-As the aspect ratio decreases to $Ar=1.75$, vortices begin to detach alternately from both sides of the elliptical cylinder at reduced frequencies and intensities.
-These detaching vortices organize into a relatively orderly and stable vortex street.
-By applying AFC based on DRL, the vortex shedding in the wake of the elliptical cylinder is completely suppressed, resulting in a stabilized wake field.
-At an aspect ratio of $Ar=1.5$, an alternating vortex structure forms behind the elliptical cylinder.
-These structures possess certain strengths and frequencies, yet the extent of vortex shedding remains comparatively mild.
-Further reduction in the aspect ratio to $Ar=1.25$ leads to a more pronounced and frequent shedding of vortical structures from the wake, indicating a transition towards more significant flow instability.
-The recirculation bubble downstream of the elliptical cylinder diminishes, oscillates, and becomes unstable.
-When the aspect ratio is $Ar=1.5$ and $Ar=1.25$, a stable, regular alternating vortex structure forms behind the elliptical cylinder.
-After the application of AFC, the shedding of vortices behind the elliptical cylinder is completely suppressed, transitioning the wake field into a stable flow structure.
-At $Ar=1$ and $Ar=0.75$, the area of the recirculation bubbles formed behind the elliptical cylinder decreases, leading to more prominent and frequent vortex shedding within the wake, indicating a transition towards significant flow instability.
-Following the implementation of AFC, the alternating vortex street behind the elliptical cylinder is transformed into a stable wake field.
-Fig.
-7 depicts the instantaneous velocity fields both baseline and post-implementation of control measures across a range of $Ar$ from 2 to 0.75 for elliptical cylinders.
-At $Ar=2$, the baseline wake field is notably stable.
-However, when the $Ar$ is reduced below 1.75, alterations in structural parameters provoke enhanced vortex shedding behind the ellipse, characterized by escalating frequency and intensity as the $Ar$ diminishes.
-The efficacy of AFC is underscored through the deployment of a control strategy, which effectively attenuates vortex shedding behind the elliptical cylinder for $Ar=1.75$, 1.5, 1.25, 1, and 0.75.
-This intervention demonstrates the significant capabilities of the DRL-based AFC technology in manipulating vortex shedding within the wake fields of elliptical cylinders across varying $Ar$, thereby enhancing the flow stability of the configurations under study.
-Figure 8: Snapshots of the velocity magnitude illustrating the effect of the active flow control.
-Left side of each row: baseline flow.
-Left side of each row: result with active flow control.
+8 display the velocity magnitude snapshots of elliptical cylinders with $Ar$ ranging from 2 to 0 under baseline and controlled conditions.
+More specifically, Fig.
+7 shows those for $Ar$ from 2 to 0.75, the vortices in the wake field of the controlled elliptical cylinders can be completely suppressed.
+In contrast, Fig.
+8 illustrates those for $Ar$ from 0.5 to 0, the vortices in the wake field of the controlled elliptical cylinders are mitigated by not fully suppressed.
+Figure 7: Comparison of the velocity magnitude snapshots for an elliptical cylinder before and after control.
+Each row displays: on the left, the baseline simulation; on the right, the AFC results.
+(a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$.
+Starting from Fig.
+7(a) for the elliptical cylinder with $Ar=2$, the baseline flow field (left) exhibits high stability with no observable vortex shedding.
+We applied a DRL-based flow control to this stable, non-shedding wake field, and the controlled flow field (right) behind the elliptical cylinder remained stable and nearly identical to the baseline.
+This sanity test indicates that the DRL-based model is consistent with flow physics and does not generate any unreasonable perturbation during the control process.
+As the $Ar$ decreases to $Ar=1.75$, the application of DRL-based AFC technology successfully suppresses the vortices that originally separated alternately from both sides of the elliptical cylinder.
+Further reducing the $Ar$ results in a more pronounced and frequent shedding of vortex structures in the uncontrolled wake, while the recirculation bubble downstream of the elliptical cylinder gradually decreases and becomes oscillatory.
+Although the flow instability gradually increases, under the conditions of $Ar=1.5$, 1.25, 1, and 0.75, the alternating vortex streets behind the elliptical cylinders are transformed into a stable wake field after the implementation of AFC in all cases.
+The DRL-based AFC technology demonstrates significant vortex shedding suppression capabilities under these conditions, highlighting the effectiveness of the AFC control strategy in enhancing the flow stability of the wake field.
+Figure 8: Comparison of the velocity magnitude snapshots for an elliptical cylinder before and after control.
+Each row displays: on the left, the baseline simulation; on the right, the AFC results.
 (a) $Ar=0.5$; (b) $Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
-As the $Ar$ is further reduced to 0.5, 0.25, and 0.1, eventually transitioning to a flat plate configuration, the vortex shedding phenomena behind the elliptical cylinder become more frequent and intense.
-It is evident that decreasing the $Ar$ leads to a significant increase in the circulation entering the wake, as the flow velocity increases when water moves around the top of the cylinder.
-This observation aligns with the findings of Thompson et al.Thompson et al.
-(2014).
-This underscores that as the $Ar$ of an elliptical cylinder gradually diminishes towards zero, it induces an enhancement in the flow instability of the wake field, consequently leading to the release of more energy in the form of vortices.
-This dynamic highlights the critical role of structural parameters in influencing the energy distribution and stability of the flow field as it interacts with varying geometrical configurations.
-However, when utilizing the AFC technique to manipulate the flow field, the control outcomes differ from those observed at $Ar=0.75$ and higher $Ar$; vortex shedding behind the elliptical cylinder is not completely suppressed.
-At $Ar=0.5$, analysis of the instantaneous velocity field reveals that the recirculation area behind the elliptical cylinder is significantly elongated, and the shedding of vortices is notably subdued.
-Nevertheless, the flow field has not fully stabilized, as there remains a continual detachment of alternating vortex structures from behind the elliptical cylinder, resembling elongated vortex formations.
-At $Ar=0.25$, the recirculation bubble behind the elliptical cylinder is similarly elongated as seen at $Ar=0.5$, but with a notable difference: the bubble now exhibits significant oscillations.
-In the controlled wake field, the phenomenon of flow separation following fluid passage around the elliptical cylinder and subsequent reattachment is also observable.
-Despite these dynamics, the instability of the downstream velocity field has been relatively stabilized compared to the baseline state, indicating the effectiveness in delaying vortex shedding.
-At $Ar=0.1$ or 0, compared to the baseline condition, the control of the flow field has primarily delayed rather than completely suppressed vortex shedding behind the elliptical cylinder.
-The recirculation bubble behind the cylinder is elongated, yet it exhibits significant oscillations.
-This indicates that while the AFC has extended the duration before vortex detachment, it has not fully stabilized the vortex structures within the wake.
-In summary, at higher $Ar$ ($Ar=0.75$, 1, 1.25, 1.5, 1.75), the DRL-based control strategy can fully suppress vortex formation, thereby stabilizing the wake field.
-However, at lower $Ar$ ($Ar=0.5$, 0.25, 0.1, 0), the effectiveness of the control strategy is limited to mitigating vortex shedding within the wake field to some extent.
-It delays the vortex detachment process and enlarges the recirculation area behind the elliptical cylinder, but oscillations in the recirculation bubble persist, and the downstream flow field continues to exhibit signs of instability.
-Figure 9: Comparison of $C_{D}$ and $C_{L}$ before and after control.(a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
+In Fig.
+8, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the baseline flow becomes increasingly unstable, and the area of the recirculation region gradually decreases while the strength and frequency of vortex shedding increase.
+When $Ar=0.5$, the frequent alternating vortex shedding is effectively suppressed after flow control is applied.
+The recirculation region behind the elliptical cylinder is significantly elongated, and the shedding of vortices is controlled, although the wake flow still exhibits alternating oscillations.
+For the controlled flow at $Ar=0.25$, the recirculation bubble behind the elliptical cylinder is elongated, resembling the case at $Ar=0.5$.
+However, there is a notable difference: the recirculation bubble now demonstrates prominent oscillations.
+The vortex shedding behind the cylinder continues alternately, akin to the elongated vortex structures being stretched longitudinally.
+Despite these instabilities, vortex shedding in the wake fields at $Ar=0.5$ and $Ar=0.25$ is reasonably mitigated compared to the baseline state.
+At $Ar=0.1$ or 0, compared to baseline conditions, the recirculation bubbles behind the cylinder are elongated, and the initiation point of vortex shedding is extended further downstream.
+We do not observe complete suppression of vortex shedding due to the highly chaotic nature of these wake behind vertical thin bodies.
+Nevertheless, the vortex shedding is reduced to a lower frequency and intensity.
+These observations reflect the effectiveness of the control strategies in suppressing vortex shedding and stabilizing the wake field.
+Figure 9: Comparison of $C_{D}$ and $C_{L}$ before and after control.
+The background color represents the unstable stage of control, and the absence of background color means that the corresponding drag or lift has been in a stable stage during this period.
+(a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
 Fig.
@@
 2.
-These observations substantiate the effectiveness of the active flow control strategy underpinned by DRL algorithms.
+These observations substantiate the effectiveness of the AFC strategy underpinned by DRL algorithms.
 The activation of control swiftly alters the periodic behavior of both $C_{D}$ and $C_{L}$, facilitating the stabilization of $C_{D}$ and $C_{L}$ in the subsequent stages.
-This demonstrates the capability of the DRL-based active flow control approach to significantly reduce the drag around an elliptical cylinder to its lowest feasible value and diminish the lift to zero.
+This demonstrates the capability of the DRL-based AFC approach to significantly reduce the drag around an elliptical cylinder to its lowest feasible value and diminish the lift to zero.
 Fig.
@@
 Although complete stability is not achieved, the average value of the $C_{D}$ is significantly reduced.
-The $C_{L}$ undergoes an initial sharp decrease and approaches near zero, and after half the total duration of control, it oscillates at a lower frequency.
-Figure 10: Snapshots of the velocity magnitude illustrating the effect of the active flow control.
-Left side of each row: baseline flow.
-Left side of each row: result with active flow control.
+Similarly, the $C_{L}$ undergoes an initial sharp decrease and approaches near zero, then later starts to oscillates, but at a lower frequency than the baseline $C_{L}$ oscillation.
+Figure 10: Comparison of $C_{D}$ and $C_{L}$ before and after control.
+The background color indicates that the drag or lift coefficient is still in an unstable stage after flow control.
 (a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
 Figure 11: Action, which is the mass flow rate of the synthetic jets.
+The background color represents that the value of the action is in an unstable stage, and the absence of a background color means that the action has been in a stable stage during this period.
 (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$.
@@
 11 illustrates the time history curve of the action (the mass flow rate of synthetic jets), When performing AFC using DRL for flow around an elliptical cylinder with $Ar$ between 0 and 2.
-When $Ar=2$, the AFC is ineffective, and the value of the action, pertaining to the control input, remains close to zero.
+When $Ar=2$, the AFC is negligible, and the value of the action, pertaining to the control input, remains close to zero.
+This is a strong indication that the DRL-based model is not generating nonphysical perturbations for a stable flow field.
 From $Ar=1.75$ to $Ar=0.75$, during the initial phases of control, the agent selects larger action values, corresponding to higher mass flow rates of synthetic jets.
@@
 7 and Fig.
-9 illustrate that the DRL-based control strategy exhibits energy efficiency and effectiveness within the range of elliptical $Ar=0.75$ to $Ar=1.75$.
+9 illustrate that the DRL-based control strategy exhibits energy efficiency and effectiveness within the range of elliptical $Ar=2$ to $Ar=0.75$.
 By progressively decreasing the mass flow rate of synthetic jets, the strategy achieves flow control, reduces drag and lift forces, and does so with reduced external energy input.
@@
 2.
-Figure 12: The learning curve of PPO algorithm training when $Ar=2$ to 0.
+Figure 12: When $Ar$ of the elliptical cylinder is in the range of 2-0, the learning curve of PPO algorithm training.
 (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$.
@@
 In Fig.
-12, we observe the learning curve during training of flow around an elliptical cylinder with different $Ar$.
-At $Ar=2$, the reward function value never reaches a positive figure, indicating that the DRL training did not achieve its objectives.
-For $Ar=1.75$ and $Ar=1.5$, the reward function initially increases from negative to positive values and then after 4000 episodes converges near the maximum reward value.
-At $Ar=1.25$, despite some initial fluctuations, the training reaches a relatively stable state after 3000 episodes and eventually settles into a stable plateau.
+12, we observe the learning curve during training of flow around elliptical cylinders with different $Ar$.
+At $Ar=2$, the learning curve has entered the plateau phase after the initial improvement phase, indicating that the DRL training has converged.
+For $Ar=1.75$ and $Ar=1.5$, the reward function initially increases from negative to positive values and then after 4,000 episodes converges near the maximum reward value.
+At $Ar=1.25$, despite some initial fluctuations, the training reaches a relatively stable state after 3,000 episodes and eventually settles into a stable plateau.
 For $Ar=1$ and $Ar=0.75$, there are significant initial fluctuations with the learning curve rising swiftly, indicating rapid progress in policy learning.
-After 2000 episodes, the training demonstrates a stable convergence trend, with minor fluctuations in the reward value between 2000 and 5000 episodes.
+After 2,000 episodes, the training demonstrates a stable convergence trend, with minor fluctuations in the reward value between 2,000 and 5,000 episodes.
 This shows that the PPO algorithm has successfully learned an effective strategy, exhibiting commendable convergence and stability in performance.
-At $Ar=0.5$, the speed of training convergence begins to slow, trending towards convergence after 4000 episodes, with the reward values stabilizing between 4000 and 10000 episodes.
-At $Ar=0.25$, training approaches convergence after 8000 episodes, reaching a total of 20000 episodes, with a mid-to-late plateau marking stable convergence.
-At $Ar=0.1$ and $Ar=0$, the training convergence is extremely slow, with signs of convergence only appearing after nearly 12000 episodes.
-Overall, the DRL algorithm exhibits rapid initial convergence and reaches a stable plateau during the mid-to-late stages for $Ar=1.25$, 1 and 0.75, showcasing excellent convergence and stability throughout the training process.
-Training also achieves stable convergence within 5000 episodes for $Ar=1.5$ and $Ar=1.75$.
+At $Ar=0.5$, the speed of training convergence begins to slow, trending towards convergence after 4,000 episodes, with the reward values stabilizing between 4,000 and 10,000 episodes.
+At $Ar=0.25$, training approaches convergence after 8,000 episodes, reaching a total of 20,000 episodes, with a mid-to-late plateau marking stable convergence.
+At $Ar=0.1$ and $Ar=0$, the training convergence is extremely slow, with signs of convergence only appearing after nearly 12,000 episodes.
+The DRL algorithm exhibits rapid initial convergence and reaches a stable plateau during the mid-to-late stages for $Ar=1.25$, 1 and 0.75, showcasing excellent convergence and stability throughout the training process.
+Training also achieves stable convergence within 5,000 episodes for $Ar=1.5$ and $Ar=1.75$.
 However, for $Ar=0.5$, 0.25, 0.1 and 0, as the flow phenomena become increasingly complex, the control challenge intensifies, leading to a noticeable slowdown in convergence speed.
 This indicates that for flow scenarios with higher degrees of instability, the convergence performance of DRL training is compromised, requiring a larger number of training iterations to reach stable convergence.
-Tab.
-2 presents the effectiveness of active flow control techniques based on DRL algorithms, targeting elliptical cylinders with $Ar$ ranging from 0.1 to 2.
-It meticulously details the values of $C_{D}$ and $C_{L}$ before and after the application of control, along with their respective rates of change.
-Additionally, it describes the magnitude of the mass flow rate of the synthetic jets utilized for control and its ratio relative to the inflow rate.
-Initially, at $Ar=2$, the wake field stability behind the elliptical cylinder is observed to be relatively high, resulting in no significant alteration in the drag experienced by the elliptical cylinder.
-When $Ar$ is reduced to 1.75, there is a noticeable decrease in drag around the elliptical cylinder by 0.9%.
-With further reduction in $Ar$, the rate of drag reduction progressively increases.
-In their study, Protas and Wesfreid (2002) elucidated that the drag coefficient consists of two components: (i) the invariant drag ($C_{D}^{b}$), which remains constant at a given Reynolds number; and (ii) the amendable drag correction ($C_{D}^{0}$), associated with oscillatory flow phenomena within the fluid, representing the additional resistance arising from fluid oscillations or fluctuations.
-This portion of drag can be adjusted through external interventions, indicating the influence of unstable flow regions and the potential for control.
-In this research, the flow field is controlled by utilizing "external excitation," such as synthetic jets, to mitigate the additional resistance caused by instability, thereby optimizing the overall fluid dynamic performance.
-As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, thereby enhancing the controllable component of the drag.
 Table 2: The control effect of elliptical cylinder under DRL-based AFC.
 $Ar$ Re $\overline{C}_{D,\text{Baseline}}$ $\overline{C}_{D,\text{Controlled}}$ Drag Reduction (%) $\overline{C}_{L,\text{Baseline}}$ $\overline{C}_{L,\text{Controlled}}$ Lift Reduction (%) $\overline{a}$ $\overline{a}$ ratio (%)
+2 100 2.684 2.684 0.0 0.016 0.016 0 0 0
+1.75 100 2.707 2.684 0.9 0.165 0.001 99.7 0.007 0.7
+1.5 100 2.765 2.707 2.1 0.302 0.002 99.4 0.009 0.9
+1.25 100 2.918 2.804 3.9 0.541 0.006 98.9 0.006 0.6
+1 100 3.207 2.951 8.0 1.022 0.011 98.9 0.001 0.1
+0.75 100 3.792 3.197 15.7 1.603 0.076 95.2 0.010 1.0
+0.5 100 4.837 3.533 26.9 1.937 0.619 68.0 0.049 4.9
+0.25 100 6.503 4.167 35.9 1.707 0.579 66.1 0.126 12.6
+0.1 100 7.812 4.876 37.6 1.196 0.596 50.2 0.499 49.9
 0 100 7.679 4.331 43.6 1.513 0.605 60.0 0.521 52.1
-0.1 100 7.812 4.876 37.6 1.196 0.596 50.2 0.499 49.9
-0.25 100 6.503 4.167 35.9 1.707 0.579 66.1 0.126 12.6
-0.5 100 4.837 3.533 26.9 1.937 0.619 68.0 0.049 4.9
-0.75 100 3.792 3.197 15.7 1.603 0.076 95.2 0.010 1.0
-1 100 3.207 2.951 8.0 1.022 0.011 98.9 0.001 0.1
-1.25 100 2.918 2.804 3.9 0.541 0.006 98.9 0.006 0.6
-1.5 100 2.765 2.707 2.1 0.302 0.002 99.4 0.009 0.9
-1.75 100 2.707 2.684 0.9 0.165 0.001 99.7 0.007 0.7
-2 100 2.684 2.684 0.0 0.016 0.016 0 0 0
   • a. $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged drag coefficient of the elliptical cylinder once the baseline flow has fully developed.
@@
   • c. $\overline{a}$ ratio (%) represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the inlet.
-In Tab.
-2, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the $C_{D}$ reduction rates are 0.9%, 2.1%, 3.9%, 8.0%, and 15.7% respectively, while the $C_{L}$ reduction rates exceed 95% in each case.
+Tab.
+2 presents the effectiveness of AFC techniques based on DRL algorithms, targeting elliptical cylinders with $Ar$ ranging from 0 to 2.
+It meticulously details the values of $C_{D}$ and $C_{L}$ before and after the application of control, along with their respective rates of change.
+Additionally, it describes the magnitude of the mass flow rate of the synthetic jets utilized for control and its ratio relative to the inflow rate.
+When the $Ar$ is set to 2, the baseline flow has reached a steady state.
+During the application of flow control, the agent does not use external energy to disrupt this stable flow field, which demonstrates the intelligence of the DRL-based AFC.
+This reflects the adaptability and intelligent decision-making of AFC control in maintaining flow stability.
+For elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the $C_{D}$ reduction rates are 0.9%, 2.1%, 3.9%, 8.0%, and 15.7% respectively, while the $C_{L}$ reduction rates exceed 95% in each case.
 These results highlight the efficacy of the DRL-based AFC technology in effectively suppressing both $C_{D}$ and $C_{L}$ within the $Ar$ range of 1.75 to 0.75, demonstrating the efficiency of the DRL-based intelligent control strategy.
@@
 […]
+In their research, Protas and WesfreidProtas and Wesfreid (2002) elucidated that the drag coefficient consists of two components: (i) the invariant drag ($C_{D}^{b}$), which remains constant at a given Reynolds number; and (ii) the amendable drag correction ($C_{D}^{0}$), associated with oscillatory flow phenomena within the fluid, representing the additional resistance arising from fluid oscillations or fluctuations.
+This portion of drag can be adjusted through external interventions, indicating the influence of unstable flow regions and the potential for control.
+In our research, the flow field is controlled by utilizing external excitation, such as synthetic jets, to mitigate the additional resistance caused by instability, thereby optimizing the overall fluid dynamic performance.
+As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, thereby enhancing the controllable component of the drag.
+This is the reason why the drag reduction rate gradually increases as $Ar$ decreases in Tab.
+2.
 The control strategies derived from DRL-based AFC technology exhibit greater complexity compared to conventional harmonic forcing implemented in previous works (Bergmann, Cordier, and BrancherBergmann, Cordier, and Brancher (2005)), highlighting the value of employing ANN as controllers.
 Typically, approximately 93% of the drag induced by the shedding of vortices is suppressed by the discovered control laws, with minimal jet intensities required for drag reduction.
-For elliptical cylinders with aspect ratios $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.
+In our research, the control strategy derived from the DRL-based AFC system dynamically responds in real time to variations in the flow state.
+This emphasizes its capability to adaptively adjust to changing conditions, demonstrating the robustness and responsiveness of the AFC framework.
+Furthermore, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.
 III.3 Analysis of the Controlled Flow
 Fig.
-13 illustrates the suppression of vortex shedding in the wake of elliptical cylinders under the influence of active flow control based on DRL.
-The elliptical cylinders examined have $Ar$ of 1.75, 1.5, 1.25, 1.0, and 0.75, scenarios in which vortex shedding can be completely suppressed.
-At the initial moment of control (T1), vortices in the wake of the elliptical cylinders are fully developed.
+13 illustrates the suppression of vortex shedding in the wake of elliptical cylinders under the influence of AFC based on DRL.
+The elliptical cylinders examined have $Ar$ of 2, 1.75, 1.5, 1.25, 1.0, and 0.75, scenarios in which vortex shedding can be completely suppressed.
+At the initial moment of control ($T_{1}$), vortices in the wake of the elliptical cylinders are fully developed.
 During this phase, vortices are observed detaching from the cylinder at a consistent frequency and intensity, forming a Kármán vortex street in the downstream flow.
-As the control process advances (T2), external excitation methods, such as synthetic jets, are introduced to suppress the vortices in the wake of the elliptical cylinders.
+As the control process advances ($T_{2}$), external excitation methods, such as synthetic jets, are introduced to suppress the vortices in the wake of the elliptical cylinders.
 The objective of the control is to alter the characteristics of the flow field to weaken or eliminate the formation and persistence of vortices.
-With the progression of the control measures, a snapshot of the flow field is captured at the conclusion of the control (T3).
-At this juncture, vortices within the wake of the elliptical cylinders are observed to be suppressed, demonstrating the effectiveness of the DRL-based active flow control strategy in stabilizing complex fluid dynamics.
-(a) $Ar=2$
-(b) $Ar=1.75$
-(c) $Ar=1.5$
-(d) $Ar=1.25$
-(e) $Ar=1$
-(f) $Ar=0.75$
+With the progression of the control measures, a snapshot of the flow field is captured at the result of the control ($T_{3}$).
+At this juncture, vortices within the wake of the elliptical cylinders are observed to be suppressed, demonstrating the effectiveness of the DRL-based AFC strategy in stabilizing complex fluid dynamics.
 Figure 13: When $Ar=2$, 1.75, 1.5, 1.25, 1, 0.75, the vortex shedding of the elliptical cylinder can be completely suppressed.
-The streamwise snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as T1, T2, and T3.
-These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the conclusion of control, respectively.
+The streamwise snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$.
+These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the result of control, respectively.
 In the study of elliptical cylinders, the phenomenon of vortex shedding and its control are crucial for understanding the structure of flow fields.
 For elliptical cylinders with different $Ar$, the behavior of the flow fields shows significant differences.
-At $Ar=2$, the instantaneous vorticity maps at times T1, T2, and T3 demonstrate a high degree of stability in the flow field, with no evident vortex shedding observed.
-For elliptical cylinders with $Ar$ values of 1.75, 1.5, and 1.25, at T1, alternating vortices form behind the cylinder, influenced by the elliptical shape effect, which results in lower intensity and frequency of vortex shedding.
-By T2, with the initiation of control, synthetic jets quickly intervene in the regular vortex shedding process, leading to a delay in vortex detachment and the formation of "elongated vortex structures." The original vortices are significantly elongated and narrowed, and the instantaneous recirculation bubbles are also elongated longitudinally and tend towards stability.
-By T3, the original vortex shedding has been completely suppressed, with no new vortices forming, and the instantaneous recirculation bubbles reach their maximum extent and remain stable, demonstrating the effectiveness of the control strategy.
-At $Ar=1$ and $0.75$, during the T1 phase, a standard Bénard–von Kármán vortex street forms behind the elliptical cylinder, characterized by a series of regularly spaced vortices.
+At $Ar=2$, both before and after control, the instantaneous vorticity maps at moments $T_{1}$, $T_{2}$, and $T_{3}$ demonstrate a high degree of stability in the flow field, with no evident vortex shedding observed.
+For elliptical cylinders with $Ar$ values of 1.75, 1.5, and 1.25, at $T_{1}$, alternating vortices form behind the cylinder, influenced by the elliptical shape effect, which results in lower intensity and frequency of vortex shedding.
+By $T_{2}$, with the initiation of control, synthetic jets quickly intervene in the regular vortex shedding process, leading to a delay in vortex detachment and the formation of "elongated vortex structures." The original vortices are significantly elongated and narrowed, and the instantaneous recirculation bubbles are also elongated longitudinally and tend towards stability.
+By $T_{3}$, the original vortex shedding has been completely suppressed, with no new vortices forming, and the instantaneous recirculation bubbles reach their maximum extent and remain stable, demonstrating the effectiveness of the control strategy.
+At $Ar=1$ and $0.75$, during the $T_{1}$ phase, a standard Bénard–von Kármán vortex street forms behind the elliptical cylinder, characterized by a series of regularly spaced vortices.
 These vortices alternately shed from both sides of the cylinder, creating a regular pattern of vortex arrangement.
-Moving into the second phase (T2), similar to cases with higher aspect ratios ($Ar=1.25$, 1.5, 1.75), the recirculation bubbles behind the cylinder are elongated longitudinally, and the originally shed vortices are also elongated and narrowed, with shearing interactions occurring between them.
-By the third time point (T3), the vortices behind the elliptical cylinder are completely suppressed, and the regions of positive and negative vorticity in the flow field nearly achieve symmetry, demonstrating high stability.
-(a) $Ar=0.5$
-(b) $Ar=0.25$
-(c) $Ar=0.1$
-(d) $Ar=0$
+Moving into the second phase ($T_{2}$), similar to cases with higher aspect ratios ($Ar=1.75$, 1.5, 1.25), the recirculation bubbles behind the cylinder are elongated longitudinally, and the originally shed vortices are also elongated and narrowed, with shearing interactions occurring between them.
+By the third time point ($T_{3}$), the vortices behind the elliptical cylinder are completely suppressed, and the regions of positive and negative vorticity in the flow field nearly achieve symmetry, demonstrating high stability.
 Figure 14: When $Ar=0.5$, 0.25, 0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed.
-The streamwise snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as T1, T2, and T3.
+The streamwise snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$.
 These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.
 Fig.
-14 delineates scenarios where the suppression of vortex shedding from elliptical cylinders is not entirely effective for aspect ratios $Ar$ of 0.5, 0.25, 0.1, and 0.
-At T1, the vortex shedding behind the cylinder is fully developed, T2 represents the control phase, and T3 provides a snapshot of the flow field after control cessation.
-For $Ar=0.5$ and $Ar=0.25$, During the T1 phase, vortices are observed detaching from the cylinder with certain frequency and intensity, forming a Kármán vortex street in the downstream wake.
-In the T2 phase, as control commences, the recirculation bubble is noted to elongate longitudinally to some extent, with the recirculation area appearing relatively symmetric and stable.
+14 delineates scenarios where the suppression of vortex shedding from elliptical cylinders is not entirely effective for $Ar$ of 0.5, 0.25, 0.1, and 0.
+At $T_{1}$, the vortex shedding behind the cylinder is fully developed, $T_{2}$ represents the control phase, and $T_{3}$ provides a snapshot of the flow field after control cessation.
+For $Ar=0.5$ and $Ar=0.25$, During the $T_{1}$ phase, vortices are observed detaching from the cylinder with certain frequency and intensity, forming a Kármán vortex street in the downstream wake.
+In the $T_{2}$ phase, as control commences, the recirculation bubble is noted to elongate longitudinally to some extent, with the recirculation area appearing relatively symmetric and stable.
 The intervention of synthetic jets delays the detachment of vortices, which are stretched and deformed due to shear with the side walls.
-By T3, the initially generated vortices are significantly elongated and deformed, exiting the computational domain, yet the recirculation bubble behind the cylinder remains unstable and continues to oscillate, leading to new vortex shedding.
+By $T_{3}$, the initially generated vortices are significantly elongated and deformed, exiting the computational domain, yet the recirculation bubble behind the cylinder remains unstable and continues to oscillate, leading to new vortex shedding.
 The emergence of new irregular vortex shedding highlights persistent instability behind the elliptical cylinder.
@@
 The instability of the wake field is mitigated, demonstrating partial efficacy of the control strategy in manipulating flow dynamics, although full stability is not achieved.
-For $Ar=0.1$ and $Ar=0$, during T1, the vortex shedding frequency and intensity from the cylinder are more severe compared to $Ar=0.5$, with enhanced flow instability.
+For $Ar=0.1$ and $Ar=0$, during $T_{1}$, the vortex shedding frequency and intensity from the cylinder are more severe compared to $Ar=0.5$, with enhanced flow instability.
 Multi-level vortex shedding occurs, with primary and secondary vortices differing in shape, frequency, and intensity.
@@
 The instability of the recirculation area manifests as asymmetry, swinging, and oscillation phenomena.
-By T3, although the original shedding vortices are elongated and expelled from the computational domain, new, unstable, and irregular shedding occurs.
+By $T_{3}$, although the original shedding vortices are elongated and expelled from the computational domain, new, unstable, and irregular shedding occurs.
 Shed vortices swinging to the side walls are sheared, stretched, and undergo deformation, resulting in irregularly shaped vortices.
-This highlights the persistent challenges in managing flow dynamics and instability at lower aspect ratios.
+This highlights the persistent challenges in managing flow dynamics and instability at lower $Ar$.
 IV CONCLUSIONS
-In this study, we investigate the use of DRL-based AFC to suppress vortex shedding, drag, and lift coefficients of an elliptical cylinder with a Reynolds number of 100.
+In this study, we investigate the use of DRL-based AFC to suppress vortex shedding, reduce drag, and mitigate lift fluctuations of elliptical cylinders at a Reynolds number of 100.
 The $Ar$ of the elliptical cylinder ranges from an ellipse ($Ar=2.0$) to a circle ($Ar=1.0$), and finally to a flat plate ($Ar=0$).
 We employ the PPO algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of the elliptical cylinder.
-The primary objective of this work is to gain further insights into the control capabilities of deep reinforcement learning algorithms for complex flow systems.
+The primary objective of this work is to gain further insights into the control capabilities of DRL algorithms for complex flow systems.
 The main findings of this study can be summarized as follows:
-  • We evaluated the impact of $Ar$ on flow control around elliptical cylinders.
-When the $Ar$ is 2, the baseline flow is highly stable, and AFC shows negligible effect.
-However, for aspect ratios of 1.75, 1.5, 1.25, 1, and 0.75, the reduction rates in the drag coefficient around the cylinder are respectively 0.9%, 2.1%, 3.9%, 8.0%, and 15.7%, while the reduction rates in lift coefficient are 99.7%, 99.4%, 98.9%, 98.9%, and 95.2%.
-These data indicate that the AFC strategy based on DRL is capable of significantly minimizing the cyclically fluctuating drag coefficient to minimal values and stabilizing it, while also maintaining the cyclically oscillating lift coefficient near zero.
+  • At $Ar=2$, the baseline flow is already in a stable state.
+When AFC is applied, the agent does not perform any actions that would disturb the flow field.
+This indicates that the DRL agent is capable of recognizing a stable flow condition and making intelligent decisions accordingly.
+However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the reduction rates in the drag coefficient around the cylinder are respectively 0.9%, 2.1%, 3.9%, 8.0%, and 15.7%, while the reduction rates in lift coefficient are 99.7%, 99.4%, 98.9%, 98.9%, and 95.2%.
+These data indicate that the AFC strategy based on DRL is capable of significantly minimizing the periodically fluctuating drag coefficient to minimal values and stabilizing it, while also maintaining the oscillating lift coefficient near zero.
 This demonstrates the efficacy of DRL in AFC applications, effectively reducing the lift and drag around elliptical cylinder with $Ar$ ranging from 0.75 to 1.75.
-Velocity cloud diagrams reveal that the original periodically shedding vortices transition to a stable flow field with no vortex shedding after flow control.
-Instantaneous vorticity maps during the control process show that with the activation of synthetic jets, the vortices shedding downstream of the elliptical cylinder are progressively elongated and squeezed out of the computational domain.
-The originally unstable recirculation area behind the cylinder also transforms into a stable, symmetrical recirculation bubble, ceasing the production of new vortices.
-These results underscore the robust adaptability of DRL-based AFC across different elliptical cylinder shapes ($Ar$ between 0.75 and 1.75), capable of learning optimal control strategies directly through interaction with the environment to achieve a reduction in drag and lift coefficients around the cylinder.
-Moreover, for elliptical cylinders with aspect ratios of 1.75, 1.5, 1.25, 1, and 0.75, the mass flow rate ratios of synthetic jets relative to the inlet flow rate are 0.7%, 0.9%, 0.6%, 0.1%, and 1.0% respectively.
-This underscores that the flow control strategy efficiently utilizes less than 1% of external excitation energy, exemplifying its energy-saving capability and fully satisfying the low-energy consumption requirements essential for effective flow control.
+Visualizations of velocity fields reveal that the original periodically shedding vortices transition to stable flow fields with no vortex shedding after flow control.
+Instantaneous vorticity maps during the control process show that with the activation of synthetic jets, the vortices downstream of the elliptical cylinder are progressively elongated.
+The originally unstable recirculation area behind the cylinder transforms into a stable, symmetrical recirculation bubble, ceasing the production of new vortices.
+These results underscore the robust adaptability of DRL-based AFC across different elliptical cylinder shapes ($Ar$ between 0.75 and 1.75), and its capability of learning optimal control strategies directly through interaction with the environment to achieve a reduction in drag and lift coefficients around the cylinder.
+Moreover, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the mass flow rate ratios of synthetic jets relative to the inlet flow rate are 0.7%, 0.9%, 0.6%, 0.1%, and 1.0% respectively.
+This underscores that the flow control strategy efficiently utilizes less than 1% of external excitation energy, exemplifying its energy-saving capability for effective flow control.
   • When the $Ar$ of the elliptical cylinder decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the drag coefficient and lift coefficient are 26.9%, 35.9%, 37.6%, and 43.6%, respectively, while the corresponding reduction rates for the lift coefficient are 68.0%, 66.1%, 50.2%, and 60.1%.
-Despite the effective reduction in both drag and lift coefficients achieved by the DRL-based AFC strategy, they have not yet stabilized compared to cases with larger $Ar$ values.
+Despite the effective reduction in both drag and lift coefficients achieved by the DRL-based AFC strategy, vortex shedding phenomena have not yet stabilized compared to cases with larger $Ar$ values.
 The initial phase of control exhibits distinct fluctuations, followed by low-frequency and low-amplitude oscillations in the drag and lift coefficients, which are lower than the baseline.
-Particularly, for $Ar$ values of 0.1 and 1, the drag and lift coefficients experience a sharp decrease initially, followed by low-amplitude, low-frequency periodic oscillations.
+Particularly, for $Ar$ values of 0.1 and 0, the drag and lift coefficients experience a sharp decrease initially, followed by moderate-amplitude, low-frequency periodic oscillations.
 This indicates that for elliptical cylinders with $Ar$ ranging from 0.5 to 0, the AFC control strategy can partially suppress lift and reduce drag.
+From the vorticity plots of the control process, for $Ar$ values of 0.5 and 0.25, the recirculation region behind the elliptical cylinder is elongated during the initial phase of control, with symmetric recirculation bubbles.
+As the control progresses, the recirculation region becomes unstable, exhibiting lateral oscillations and the shedding of long, narrow vortices.
+Similarly, for $Ar$ values of 0.1 and 0, the recirculation region undergoes elongation during the control phase, with high-frequency and large-amplitude oscillations, accompanied by the shedding of new vortices.
+After a significant period of control, the shedding of vortices is somewhat alleviated, with a decrease in both the frequency and intensity of vortex shedding.
 Correspondingly, for elliptical cylinders with $Ar$ values of 0.5, 0.25, 0.1, and 0, the ratios of the mass flow rates of the synthetic jets relative to the incoming flow rate are 4.9%, 12.6%, 49.9%, and 52.1%, respectively.
 This suggests that despite the use of higher external energy, the lift coefficient cannot be completely controlled to prevent oscillations and fluctuations.
-From the vortex plots of the control process, for $Ar$ values of 0.5 and 0.25, the recirculation region behind the elliptical cylinder is elongated during the initial phase of control, with symmetric recirculation bubbles.
-As the control progresses, the recirculation region becomes unstable, exhibiting lateral oscillations and the shedding of long, narrow vortices.
-Similarly, for $Ar$ values of 0.5 and 0.25, the recirculation region undergoes elongation during the control phase, with high-frequency and large-amplitude oscillations, accompanied by the shedding of new vortices.
-After control cessation, the shedding of vortices is somewhat alleviated, with a decrease in both the frequency and intensity of vortex shedding.
 In conclusion, for elliptical cylinders with $Ar$ ranging from 1.75 to 0.75, intelligent agents trained using the PPO algorithm have adeptly learned to adopt an energy-efficient flow control strategy.
-The control policies based on the PPO algorithm significantly reduce lift around the elliptical cylinder, greatly diminish drag, and suppress vortex shedding in the wake.
+The control policies based on the PPO algorithm significantly reduce lift around the elliptical cylinder, greatly diminish drag, and completely suppress vortex shedding in the wake.
 Regarding elliptical cylinder with $Ar$ values ranging from 0.5 to 0, the control strategies obtained from agents trained with the PPO algorithm can moderately alleviate vortex shedding and reduce the amplitude and frequency of fluctuations in lift and drag coefficients.
-These control outcomes demonstrate several unique capabilities of AFC strategies based on DRL, making them particularly suitable for managing complex fluid dynamics systems.
-These capabilities include adaptability, real-time optimization, robustness to uncertainties, the ability to handle complex problems, minimal human intervention, and high control efficiency.
-Acknowledgements.
+These control results demonstrate the adaptability of DRL-based AFC strategies to various geometric shapes, their robustness against uncertainties, and their ability to handle complex problems.
+This makes DRL-based AFC particularly suitable for controlling complex fluid dynamic systems.
+ACKNOWLEDGMENTS
 The authors would like to express their gratitude to Dr.
@@
 Jichao Li (National University of Singapore, Singapore), and Mr.
-Qiulei Wang (The University of Hong Kong, Hong Kong SAR, China) for making their open-source codes for deep reinforcement learning and numerical simulation available online.
+Qiulei Wang (The University of Hong Kong, Hong Kong SAR, China) for making their open-source codes for DRL and numerical simulation available online.
 These resources have significantly contributed to the research presented in this paper.
@@
 (2019), https://github.com/npuljc/RL_control_Nek5000 Li (2021); Li and Zhang (2022), and https://github.com/venturi123/DRLinFluids Qiulei Wang (2022); Wang et al.
-(2022b).
-Declaration of interests
+(2022a).
+AUTHOR DECLARATIONS
+IV.1 Conflict of Interest
 The authors report no conflict of interest.
+IV.2 Author Contributions
+Wang Jia: Conceptualization (equal); Data curation (lead); Formal analysis (lead); Methodology(lead); Validation (lead); Visualization (lead); Writing – original draft (lead).
+Hang Xu: Conceptualization (equal); Investigation (equal); Supervision (lead); Validation (equal); Writing –review & editing (lead).
 DATA AVAILABILITY

### v3

@@
 Abstract
-This study investigates the effectiveness of active flow control (AFC) technology supported by deep reinforcement learning (DRL) applied to flows around elliptical cylinders at $Re=100$.
-We vary the aspect ratio ($Ar$) of the elliptical cylinder from an ellipsoid ($Ar=2.0$) to a circular shape ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$).
-We utilize the proximal policy optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on both the upper and lower surfaces of the cylinder.
-The control objective focuses on reducing drag, minimizing lift and suppressing vortex shedding.
-We examine the robustness and adaptability of DRL-based control techniques across different geometric configurations.
+This study investigates the effectiveness of active flow control (AFC) technology supported by deep reinforcement learning (DRL) applied to flows around an elliptical cylinder at $Re=100$.
+We vary the aspect ratio ($Ar$) of the elliptical cylinder from an elliptical cylinder ($Ar=2.0$) to a circular cylinder ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$).
+We utilize the Proximal Policy Optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of a cylinder to achieve reduction in drag, minimization of lift, and suppression of vortex shedding.
+The robustness and adaptability of DRL-based control techniques are examined across different geometric configurations.
 Our research findings indicate that, for elliptical cylinders with $Ar$ between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%.
-Notably, the DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities.
+The DRL-based control strategy not only significantly reduces lift and drag, but also completely suppresses vortex shedding while using less than 1% of external excitation energy, demonstrating its efficiency and energy-saving capabilities.
 Additionally, for $Ar$ from 0.5 to 0, the reduction in drag coefficient ranges from 26.9% to 43.6%, and the reduction in lift coefficient from 50.2% to 68.0%.
 This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding.
-Overall, our findings underscore the adaptability and potential of DRL-based AFC in controlling complex fluid dynamics across diverse geometric configurations.
+It is expected underscore the adaptability and potential of DRL-based AFC in controlling complex fluid dynamics across diverse geometric configurations.
 preprint: AIP/123-QED
@@
 Sophisticated control strategies that require consideration of a wide range of spatiotemporal scales are indispensable when addressing phenomena such as unstable flows, turbulence, and multiphase flows.Subramaniam (2020)
-Recently, the field of AFC has witnessed revolutionary progress, thanks to breakthroughs in machine learning and deep learning technologies.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020) Deep reinforcement learning (DRL) is an avant-garde technique that combines the principles of deep learning with reinforcement learning to optimize decision-making through interactions with an environment.Arulkumaran et al.
+Recent advances in machine learning and deep learning technologies have revolutionized the field of artificial fluid control Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020).
+Deep Reinforcement Learning (DRL) combines deep learning and reinforcement learning to optimize decision-making through interactions with the environment Arulkumaran et al.
 (2017); François-Lavet et al.
 (2018); Wang et al.
-(2020) Deep learning excels at autonomously extracting intricate features from voluminous datasets, while reinforcement learning concentrates on how to make optimal decisions based on these features by guiding the learning process through rewards.Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015) By utilizing deep neural networks to represent policies or value functions within reinforcement learning, DRL enables machines to tackle complex decision-making challenges and high-dimensional state spaces that were previously considered intractable.Arulkumaran et al.
+(2020).
+DRL excels at extracting complex features from large datasets and making optimal decisions based on these features Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015).
+By using deep neural networks, DRL is capable of handling high-dimensional state spaces and enables machines to tackle previously intractable decision-making challenges Arulkumaran et al.
 (2017); François-Lavet et al.
 (2018); Wang et al.
-(2020) Renowned for its robust non-linear modeling capabilities and adaptive learning mechanisms, DRL has the capability to directly learn complex strategies from raw sensory data without relying on manually engineered features.Dargazany (2021) Furthermore, DRL is inherently self-improving, continuously refining its strategies through ongoing interactions with the environment.Wang et al.
-(2020) The self-learning and adaptive nature of DRL grants a significant advantage, particularly in resolving decision-making problems in dynamic settings.François-Lavet et al.
+(2020).
+DRL’s nonlinear modeling capabilities and adaptive learning mechanisms allow it to learn complex strategies directly from raw sensory data without manual feature engineering Dargazany (2021).
+It continuously improves its strategies through interactions with the environment Wang et al.
+(2020), making it advantageous for dynamic decision-making problems François-Lavet et al.
 (2018); Wang et al.
-(2020) The applications of DRL are vast and encompass various domains, including gaming, autonomous vehicles, robotic control, natural language processing, and computer vision.
-In the gaming industry, distinguished achievements like AlphaGo’s victory over world champions highlight the potential of DRL.Granter, Beck, and Papke (2017) In the realm of autonomous driving, DRL is employed to optimize route planning and decision-making.Sallab et al.
-(2017) In the field of robotics, it enables robots to learn complex physical tasks.Zhao, Queralta, and Westerlund (2020) Overall, DRL offers the potential to develop adaptive and intelligent control strategies that can adapt to changing environments and optimize performance in a wide range of applications.
+(2020).
+DRL has found applications in various fields including gaming, autonomous vehicles, robotics, natural language processing, and computer vision.
+In particular, it has achieved distinguished success, such as AlphaGo’s victory over world champions in gaming Granter, Beck, and Papke (2017), route planning and decision-making optimization in autonomous driving Sallab et al.
+(2017), and complex physical tasks learning in robotics Zhao, Queralta, and Westerlund (2020).
 With progressive advancement in the DRL technology, there has been a significant increase in the application of DRL in the field of fluid dynamics as well.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al.
@@
 (2022); Vinuesa (2024)
-In the academic literature, flow control around blunt bodies using DRL has attracted considerable interest due to its pivotal role in reducing drag, a key aspect of fluid dynamics research.
-This pursuit carries several important implications, including enhanced energy efficiency, improved performance, noise and vibration reduction, and industrial process optimization.
-As flow around blunt body tends to involve strong non-linearities such as flow separation and recirculation, traditional control methods often struggle in achieving control objectives whereas DRL models have shown promising results.
+In literature, DRL has attracted considerable interest for flow control around blunt bodies, as flow around blunt body tends to involve strong non-linearities such as flow separation and recirculation, traditional control methods often struggle in achieving control objectives whereas DRL models have shown promising results.
 Pioneered by Rabault et al.
@@
 Heess et al.
-(2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024) applied the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder.
+(2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024a) applied the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder.
 Consistently, they achieved an approximate 8% reduction in drag in their validation experiments.
 Furthermore, Li and ZhangLi and Zhang (2022) ,Wang et al.Wang et al.
-(2022a), Jia and XuJia and Xu (2024), and Wang et al.Wang et al.
-(2022b) respectively conducted flow control studies on confined circular cylinders and square cylinders using DRL.
+(2022a), Wang and XuWang and Xu (2024b) and Wang et al.Wang et al.
+(2022b)He et al.
+(2023) respectively conducted flow control studies on confined circular cylinder, square cylinder and airfoil using DRL.
 In addition, Fan et al.Fan et al.
@@
 (2020); Li and Zhang (2022); Wang et al.
-(2022a); Wang and Xu (2024).
-The existing studies have successfully demonstrated the effectiveness of using DRL for flow control, particularly in actively manipulating the flow around circular cylinders and square cylinders.
+(2022a); Wang and Xu (2024a).
 Although active flow control strategies for both circular and square cylinders are pivotal in various engineering applications for managing aerodynamic and hydrodynamic forcesVignon, Rabault, and Vinuesa (2023); Rabault et al.
@@
 Current literature indicates a distinct lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies.
-This research proposes to bridge these gaps by focusing on the flow dynamics around parameterized ellipsoids, examining how variations in the aspect ratio influence flow fields and their corresponding DRL-driven control strategies.
-By extending the investigation from elongated ellipsoids to circular cylinders and subsequently normal flat plates, this study aims to develop a deeper understanding of the physical validity of DRL approaches and their adaptability across varying body geometries.
-This work extends the studies conducted by Wang and XuWang and Xu (2024); Jia and Xu (2024) by applying the DRL-based AFC approach to elliptical cylinders, which present more complex flow characteristics and additional control challenges.
-The objective is to develop effective control strategies to reduce drag, improve flow stability, and enhance overall performance.
+Based on the studies conducted by Jia and Xu Wang and Xu (2024a, b), we focus on the development of the DRL technqiue for active flow control around elliptical cylinders, with the objective of developing effective control strategies for drag reduction, flow stability enhancement, and overall performance improvement, addressing the research gap in applying DRL to different geometric shapes of elliptical cylinders.
 The structure of this paper is organized as follows: Sec.
-II introduces the problem description and provides a comprehensive overview of DRL and its framework for application in AFC.
+II introduces the problem description and provides a comprehensive procedure for implementing DRL.
 Sec.
@@
 IV concludes the paper by summarizing the key findings and discussing their implications.
-II PROBLEM DESCRIPTION AND METHODOLOGY
-II.1 Numerical simulation
+II PROBLEM AND METHODOLOGY
+II.1 Problem statement
 Model configuration
@@
 $\nabla\cdot\mathbf{u}=0\quad\text{on}\quad\Omega\times(0,T),$ (1b)
-where $\mathbf{u}$, $t$ and $p$ are the velocity vector, time, and pressure, respectively. In this study, the $Re$ is set to 100.
+where $\mathbf{u}$, $t$ and $p$ are the velocity vector, time, and pressure, respectively.
+The Reynolds number, denoted by $Re$, is defined as $Re=\frac{\overline{U}D}{\nu}$, where $\overline{U}=2U_{m}/3=1$ represents the mean velocity magnitude, $U_{m}$ is the maximum velocity, and $\nu$ is the kinematic viscosity.
+In this study, $Re=100$ is utilized.
 Boundary conditions
-The upstream boundary of the computational domain, designated as $\Gamma_{\text{in}}$, is characterized by a parabolic velocity profile oriented in the $x$-axis, where the boundary conditions are specified such that the transverse velocity component is set to zero, and the velocity along the stream direction shows a parabolic profile distribution.
-On the downstream boundary, denoted as $\Gamma_{\text{out}}$, a Neumann-type boundary condition is prescribed for the velocity field, ensuring that the stress vector is nullified.
-On the upper and lower boundaries $\Gamma_{\text{w}}$ , a slip-wall boundary condition is employed, which stipulates that both the velocity component perpendicular to these boundaries and the tangential stress vector component along them are assigned a value of zero.
-The synthetic jets boundaries are denoted by $\Gamma_{i}(i=1,2)$.
-The velocity within the jets adheres to a parabolic distribution, with the sign of the jet velocity indicative of fluid suction or injection.
-The design of the synthetic jets ensures mass flow balance, as expressed by the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, to maintain mass conservation in the system.
-Except for the synthetic jets, the non-slip solid wall boundary condition $\Gamma_{\text{w}}$ is adopted for the rest of the cylinder.
-In compliance with the above requirements, the system is considered with the following boundary conditions
-$-\rho\mathbf{n}\cdot\mathbf{p}+Re^{-1}(\mathbf{n}\cdot\nabla\mathbf{u})$ $=0\quad\text{on }\Gamma_{out},$ (2)
-$\mathbf{u}$ $=0\quad\text{on }\Gamma_{w},$
-$\mathbf{u}$ $=U\quad\text{on }\Gamma_{in},$
-$\mathbf{u}$ $=f_{Q_{i}}\quad\text{on }\Gamma_{i},\quad i=1,2.$
-Here, $U$ is the inflow velocity profile.
-$f_{Q_{i}}$ represents the radial velocity profiles simulating the suction or injection of fluid by the jets.
-The specific formulas of $U$ and $f_{Q_{i}}$ will be given below.
-At the inlet $\Gamma_{\text{in}}$, the inflow velocity $U$ along the $x$-axis is prescribed by a parabolic velocity profile in the form,
-$U=U_{m}\frac{(H-2y)(H+2y)}{H^{2}}.$ (3)
-Here, $U_{m}$ is the maximum velocity magnitude of the parabolic profile, and $H=4.1D$ represents the total height of the rectangular domain.
-The mean velocity magnitude is $\overline{U}=2U_{m}/3=1$ and is used to define the Reynolds number $Re=\overline{U}D/\nu$.
-The functions $f_{Q_{i}}$ represent radial velocity profiles that simulate the suction or injection mechanisms of the jets.
-These profiles are selected to ensure a smooth transition to the no-slip boundary conditions on the surfaces $\Gamma_{W}$ of the cylinder.
-Specifically, $f_{Q_{i}}$ is defined as $A(\theta;Q_{i})(x,y)$, with the modulation relying on the angular coordinate $\theta$ as illustrated in Fig.
+The boundary conditions for the computational domain are as follows: The upstream boundary ($\Gamma_{\text{in}}$) is characterized by a parabolic velocity profile, the downstream boundary ($\Gamma_{\text{out}}$) is subjected to a Neumann-type boundary condition to nullify the stress vector, the upper and lower boundaries ($\Gamma_{\text{w}}$) employ a slip-wall boundary condition with zero velocity and tangential stress, and the boundaries of the synthetic jets ($\Gamma_{i}$, where $i=1,2$) are assigned a prescribed parabolic velocity profile representing fluid suction or injection.
+The synthetic jets feature a parabolic velocity distribution, where the sign of the jet velocity indicates fluid suction or injection.
+The design ensures mass flow balance through the condition $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, preserving mass conservation in the system.
+The non-slip solid wall boundary condition is applied to the rest of the elliptical cylinder except for the synthetic jets.
+Mathematically, the boundary conditions are written as:
+$\begin{array}[]{cl}-\rho\mathbf{n}\cdot\mathbf{p}+Re^{-1}(\mathbf{n}\cdot\nabla\mathbf{u})=0&\text{on\quad}\;\Gamma_{out},\\ \mathbf{u}=0&\text{on \quad}\Gamma_{w},\\ \;\mathbf{u}=U&\text{on \quad}\Gamma_{in},\\ \;\;\;\mathbf{u}=f_{Q_{i}}&\text{on \quad}\Gamma_{i},\quad i=1,2.\end{array}$ (6)
+Here, $\mathbf{n}$ denotes the normal vector, $U$ is the inflow velocity, and $f_{Q_{i}}$ represents the radial velocity profiles simulating the suction or injection of fluid by the jets.
+At the inlet $\Gamma_{\text{in}}$, the inflow velocity $U$ along the $x$-axis is prescribed by a parabolic velocity profile given by:
+$U=U_{m}\frac{(H-2y)(H+2y)}{H^{2}}$ (7)
+where $H=4.1D$ represents the total height of the rectangular domain.
+In particular, we choose $f_{Q_{i}}=A(\theta;Q_{i})(x,y)$, where the modulation relies on the angular coordinate $\theta$ as illustrated in Fig.
+1(b).
+For the jet, which has a width of $\omega$ and is centered at $\theta_{0}$ on the cylinder with radius $R$, the modulation is defined as follows:
+$A(\theta;Q)=Q\frac{\pi}{2\omega R^{2}}\cos\left(\frac{\pi}{\omega}(\theta-\theta_{0})\right).$ (8)
+This modulation function ensures a smooth transition to the no-slip boundary conditions on the cylinder surfaces, enabling a comprehensive analysis of the flow characteristics and control challenges associated with the elliptical cylinder geometry.
+Quantities of interest
+The lift coefficient ($C_{L}$) and drag coefficient ($C_{D}$) are physically important quantities in active flow control, and they are defined as follows:
+$C_{L}=\frac{F_{L}}{0.5\rho\overline{U}^{2}D},\quad C_{D}=\frac{F_{D}}{0.5\rho\overline{U}^{2}D}.$ (9)
+Here, $F_{L}$ and $F_{D}$ represent the lift and drag forces integrated over the surface of the elliptical cylinder, respectively, and $\rho$ is the fluid density.
+The Strouhal number ($St$) is used to characterize the characteristic frequency of oscillatory flow phenomena, and it is defined as follows:
+$St=\frac{f_{s}\cdot D}{\overline{U}}$ (10)
+where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$.
+II.2 Numerical simulation
+Solver details
+We utilize the open-source CFD package OpenFOAM as the incompressible flow solver, following the methodology described by Jasak et al.Jasak (2009); Jasak et al.
+(2007).
+This approach is widely employed due to its ability to provide reliable numerical algorithms for solving the Navier-Stokes equations.
+The solver implemented in OpenFOAM uses the finite volume method to discretize the computational domain into a mesh consisting of control volumes.
+To ensure numerical stability, a time step size of $\Delta t=0.0005$ is selected in the simulation.
+Grid scheme
+As an example, we examine a computational domain comprising an elliptical cylinder with an aspect ratio ($Ar=1$) of 1.
+To discretize this domain accurately, a hybrid meshing approach is adopted, integrating both structured (quadrilateral) and unstructured (triangular) meshes, as shown in Fig.
 2.
-For the jet characterized by a width of $\omega$ and centered at $\theta_{0}$, which is positioned on the cylinder with radius $R$, the modulation is established as follows:
-$A(\theta;Q)=Q\frac{\pi}{2\omega R^{2}}\cos\left(\frac{\pi}{\omega}(\theta-\theta_{0})\right).$ (4)
-Solver details
-In this study, the incompressible flow solver is based on the open-source CFD package OpenFOAM, as described by Jasak et al.Jasak (2009); Jasak et al.
-(2007) OpenFOAM is a widely used and validated CFD software package that provides robust numerical algorithms for solving the Navier-Stokes equations.
-The solver in OpenFOAM utilizes the finite volume method to discretize the computational domain into a mesh made up of control volumes.
-To ensure numerical stability, the time step is chosen as $\Delta t$ = 0.0005.
-Grid system
-Considering an elliptical cylinder computational domain with $Ar=1$, the domain is discretized employing a hybrid meshing approach that integrates both structured (quadrilateral) and unstructured (triangular) meshes.
-As delineated in Fig.
-2, the computational domain is divided into 18,484 mesh cells.
-Specifically, the area surrounding the elliptical cylinder is discretized using a multi-layer quadrilateral mesh, while the remaining regions are discretized with triangular meshes.
-Fig.2(a) displays the global discretization scheme for the computational domain, highlighting the comprehensive approach to mesh division.
-Fig.2(b) provides a detailed view of the mesh partitioning near the elliptical cylinder, where the multi-layered quadrilateral meshes surrounding the cylinder are notably finer compared to the triangular meshes.
-Fig.2(c) illustrates the division of multi-layered quadrilateral meshes around the elliptical cylinder, emphasizing the precision of discretization near the structure.
-Finally, fig.2(d) showcases the mesh division around the elliptical cylinder and the integration strategy between quadrilateral and triangular meshes.
+The computational domain is discretized into 18,484 mesh cells, with a multi-layer quadrilateral mesh used for the area surrounding the elliptical cylinder and triangular meshes used for the remaining regions.
+In detail, Fig.
+2(a) provides an overview of the global discretization scheme, demonstrating the comprehensive approach to mesh division employed in the computational domain.
+Fig.
+2(b) offers a closer look at the mesh partitioning near the elliptical cylinder, highlighting the finer multi-layered quadrilateral meshes surrounding the cylinder compared to the triangular meshes.
+Fig.
+2(c) illustrates the division of the multi-layered quadrilateral meshes around the elliptical cylinder, emphasizing the precision in discretization near the structure.
+Fig.
+2(d) illustrates the mesh division strategy around the elliptical cylinder, demonstrating the integration between the quadrilateral and triangular meshes.
 The method for mesh division for ellipses with $Ar$ values other than 1 follows the same approach as that for $Ar=1$, and is not separately demonstrated.
@@
 (d) A magnified view of the quadrilateral meshes in the vicinity of the elliptical cylinder.
-Quantities of interest
-The lift coefficient ($C_{L}$) and drag coefficient ($C_{D}$) are defined as
-$C_{L}=\frac{F_{L}}{0.5\rho\overline{U}^{2}D},\quad C_{D}=\frac{F_{D}}{0.5\rho\overline{U}^{2}D}.$ (5)
-Here, $F_{L}$ and $F_{D}$ represent the lift and drag forces integrated on the surface of the elliptical cylinder, respectively, and $\rho$ is the fluid density.
-The Strouhal number $(St)$ is used to describe the characteristic frequency of oscillatory flow phenomena and is defined as follows:
-$St=\frac{f_{s}\cdot D}{\overline{U}}$ (6)
-Where $f_{s}$ is the shedding frequency calculated based on the periodic evolution of the $C_{L}$. $D$ is the characteristic length of the elliptical cylinder, and $\overline{U}=1$.
 Grid independence
-Taking the case with an $Ar$ of 1 as an example, this study explores mesh convergence and validates the numerical method utilized.
+Similarly, the case where the $Ar$ value is 1 is also used as an example to explore grid convergence and validate the numerical method utilized.
 Tab.
-1 presents the simulation results using three meshes with differing resolutions, enabling a comparative analysis of the maximum drag coefficient $C_{D,\max}$, the mean of the maximum drag coefficient $C_{D,\text{mean}}$, the maximum lift coefficient $C_{L,\max}$, and the $St$.
+1 presents simulation results using three different mesh resolutions, allowing for a comparative analysis of the maximum drag coefficient $C_{D,\max}$, the mean of the maximum drag coefficient $C_{D,\text{mean}}$, the maximum lift coefficient $C_{L,\max}$, and the $St$.
 Furthermore, the results of the numerical simulations are compared with the findings of Rabault et al.
@@
 (1996) for validation.
-The simulations conducted using a coarse mesh revealed significant deviations in the drag and lift coefficients when compared to the results by Rabault et al.
+The simulations conducted using a coarse mesh reveal significant deviations in the drag and lift coefficients when compared to the results obtained by Rabault et al.
 Rabault et al.
-(2019), questioning the accuracy of such numerical simulations.
-On the other hand, the fine mesh, characterized by a high density of grid points, considerably slows down the training speed for reinforcement learning, presenting a challenge for efficient simulation execution.
-Our aim is to perform numerical simulations with the least number of mesh cells without compromising on computational accuracy.
-Consequently, the discretization strategy of the main mesh was adopted for subsequent numerical simulations and reinforcement learning training.
-This strategy not only satisfies the computational accuracy requirements but also keeps the mesh element count to a manageable level.
-Comparison of the computational results from the main mesh with those obtained by Rabault et al.
+(2019), which raises doubts about the accuracy of such numerical simulations.
+On the other hand, using a fine mesh with a high density of grid points significantly slows down the training speed for reinforcement learning, posing challenges for efficient simulation execution.
+It is always desirable to use the minimum number of grid elements for numerical simulation without affecting computational accuracy.
+As a result, the discretization strategy of the main mesh is adopted for subsequent numerical simulations and reinforcement learning training.
+The comparison of the computational results obtained from the main mesh with those obtained by Rabault et al.
 Rabault et al.
-(2019).
-shows very little difference, indicating a high level of accuracy.
-Furthermore, the accuracy of the main mesh’s results lies within the upper and lower bounds provided by standard benchmarksSchäfer et al.
-(1996), confirming that the computational precision of the main mesh is adequate for further training purposes.
-The above analysis ensures the best balance between numerical accuracy and computational efficiency.
-The methodology for examining mesh dependency for ellipses with $Ar$ other than 1 is consistent with the procedure established for $Ar=1$.
+(2019) reveals minimal differences, indicating a high level of accuracy.
+Additionally, the accuracy of the main mesh’s results falls within the upper and lower bounds provided by standard benchmarks Schäfer et al.
+(1996), which confirms that the computational precision of the main mesh is sufficient for further training purposes.
+The methodology for examining mesh dependency for ellipses with an $Ar$ value other than 1 follows the same procedure established for $Ar=1$.
 Table 1: Grid independence test for an elliptical cylinder with $Ar=1$ at $Re=100$.
@@
 Fine 25,624 3.228 3.207 0.992 0.301
-II.2 Deep Reinforcement learning
-Deep learning
-Deep learning, a subset of machine learning, employs multi-layer artificial neural networks to model and learn intricate nonlinear relationships.
-Fig.
-3 demonstrates how deep learning captures the representation of input data across multiple layers of a neural network, including the input layer, several hidden layers, and the output layer.
-Each layer is composed of numerous neural units, with information transmitted between nodes via connection weights.
-The fundamental concept behind deep learning is its ability to autonomously learn data representations through successive layers.
-Utilizing the backpropagation algorithm, the network autonomously discerns the transformation process from raw data to sophisticated, high-level abstract features, eliminating the need for manual feature extraction.
-Figure 3: DRL combines deep learning’s ability to interpret complex data with reinforcement learning’s strategy to make informed decisions.
-Reinforcement learning
-In the domain of Reinforcement Learning, the interaction between an agent and its environment is the core mechanism by which learning is facilitated.
-As illustrated in Figure 3, this interaction proceeds in a cyclical pattern.
-At each timestep $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the relevant information necessary for making decisions.
-Based on this observation, the agent selects and executes an action $a_{t}$, according to its policy $\pi$.
-This policy, which is essentially a strategy that maps states to actions, may function deterministically or probabilistically.
-After the action is executed, the environment transitions to a new state $s_{t+1}$ and awards the agent with an immediate reward $r_{t+1}$.
-This reward, quantified as a numerical value, evaluates the action’s efficacy in progressing towards the agent’s objectives.
-Subsequently, the agent assimilates this feedback—comprised of both the reward and the new state—to refine its policy, employing RL algorithms designed to optimize the accumulation of long-term rewards.
-Deep Reinforcement learning
-Deep learning, a specialized subset of machine learning, excels at identifying patterns in extensive datasets through its multi-layer neural network architecture.
-This approach enables computers to derive complex representations directly from raw data, leading to revolutionary advances in fields like image recognition, natural language processing, and predictive analytics.
-DRL marries the potent perceptual capabilities of deep learning with the decision-making and optimization capabilities of reinforcement learning, effectively overcoming their individual limitations.
-Deep learning addresses the challenge of processing original, high-dimensional input data, which is problematic for traditional reinforcement learning, while reinforcement learning provides a structured decision-making framework and a goal-oriented learning mechanism that enhances the capabilities of deep learning.
-As depicted in Figure 3, a deep learning network transforms the intricate state of the environment into a compressed and efficient feature representation.
-This transformation yields a more manageable and interpretable state space for reinforcement learning algorithms.
-In DRL, reinforcement learning leverages these high-level features to learn and optimize decision-making policies.
-These policies are typically parameterized by deep neural networks, implying that the selection of actions is influenced not only by the current observation of the environmental state but also by the learned weights within the network.
-II.3 Proximal Policy Optimization algorithm
-PPO is an advanced DRL algorithm suitable for addressing reinforcement learning challenges within both discrete and continuous action spaces.
-The fundamental concept of PPO revolves around enhancing performance by maximizing a surrogate objective function while ensuring the stability of policy updates.
-In PPO, an agent interacts with an environment to gather experience data, which is then utilized to construct the objective function.
-This function incorporates dominant estimates from the experience data to gauge the direction of policy improvement.
-During these interactions, the agent generates a sequence of trajectories ($\tau$).
-A trajectory consists of a series of state-action pairs produced as the agent operates under the current policy, continuing until a terminal state is reached or a predetermined maximum number of steps is exceeded.
-Symbolically, a trajectory $\tau$ can be represented as follows:
-$\tau=(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},s_{2},\ldots,s_{T-1},a_{T-1},r_{T},s_{T}),$ (7)
-where, $s_{t}$ represents the state at time step $t$; $a_{t}$ represents the action at time step $t$; $r_{t+1}$ represents the immediate reward obtained after taking action $a_{t}$; $T$ represents the length of the trajectory; $s_{0}$ is the initial state, and $s_{T}$ is the terminal state.
-The cumulative reward $R(\tau)$ of a trajectory $\tau$ is obtained by summing the rewards $r_{t}$ obtained at each time step $t$:
-$R(\tau)=\sum_{t=0}^{T}r_{t},$ (8)
-The objective function $J(\pi_{\theta})$ in the PPO algorithm aims to maximize the expected cumulative reward obtained by following the policy $\pi_{\theta}$.
-Mathematically, this objective function is expressed as:
-$J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[R(\tau)],$ (9)
-where $J(\pi_{\theta})$ represents the expected return under policy $\pi_{\theta}$, with the expectation taken over trajectories sampled from the policy distribution $\pi_{\theta}$.
-The term $R(\tau)$ denotes the total reward accumulated along a trajectory $\tau$.
-The policy gradient $\nabla_{\theta}J(\pi_{\theta})$, indicating the direction of steepest ascent in the objective function $J(\pi_{\theta})$, is computed using the policy gradient theorem.
-The gradient of the expected return is expressed as the expectation of the gradient of the log-probability of trajectories weighted by the return:
-$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[\nabla_{\theta}\log P(\tau\mid\theta)R(\tau)],$ (10)
-where $\nabla_{\theta}\log P(\tau\mid\theta)$ represents the gradient of the log-probability of trajectories with respect to the policy parameters $\theta$.
-This computation involves sampling trajectories from the policy distribution $\pi_{\theta}$ and weighting the log-probability gradients by their respective returns.
-The policy parameters $\theta$ are updated iteratively using gradient ascent, with the update rule:
-$\theta_{k+1}=\theta_{k}+\alpha\nabla_{\theta}J(\pi_{\theta}),$ (11)
-where $\theta_{k}$ represents the policy parameters at the $k$-th iteration, $\alpha$ is the learning rate controlling the step size, and $\nabla_{\theta}J(\pi_{\theta})$ denotes the policy gradient.
-In the PPO algorithm, the trajectories generated by the agent are instrumental in shaping the policy update process.
-The objective function, $J(\pi_{\theta})$, is designed to enhance the agent’s performance by maximizing the expected cumulative reward under the policy $\pi_{\theta}$.
-This optimization involves iteratively updating the policy parameters $\theta$ using the policy gradient $\nabla_{\theta}J(\pi_{\theta})$, while maintaining stability and preventing substantial policy shifts.
-The policy gradient is estimated using a collection of trajectories, $\mathcal{D}=\{\tau_{i}\}_{i=1,\ldots,N}$, each generated by the agent interacting with the environment according to the policy.
-The policy gradient can be calculated as follows:
-$\hat{g}=\frac{1}{|\mathcal{D}|}\sum_{\tau\in\mathcal{D}}\sum_{t=0}^{T}\nabla_{\theta}\log\left(\pi_{\theta}(a_{t}\mid s_{t})\right)R(\tau),$ (12)
-where $|\mathcal{D}|$ represents the number of trajectories in $\mathcal{D}$, and $T$ is the length of each trajectory.
-This formulation offers a simplified approach to computing the policy gradient.
-Moreover, PPO commonly employs a clipped surrogate objective function to facilitate more stable policy updates.
-This objective is expressed as:
-$L^{CLIP}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{\tau\in\mathcal{D}}\sum_{t=0}^{T}\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right),$ (13)
-where $r_{t}(\theta)$ is the likelihood ratio between the old and new policies, $A_{t}$ denotes the advantage function, and $\epsilon$ is a hyperparameter that defines the clipping range.
-The inclusion of this clipped surrogate objective aids in moderating the extent of policy updates, thereby stabilizing the training process.
-In summary, the PPO algorithm presents a robust framework for the training of DRL agents.
-It operates through an iterative process of updating policy parameters aimed at maximizing the expected cumulative reward.
-Crucially, PPO maintains stability and mitigates the risk of substantial policy deviations.
-Employing policy gradients along with clipped surrogate objectives, PPO ensures effective and stable training across a diverse array of reinforcement learning tasks.
-This methodology not only enhances the efficiency of learning but also contributes significantly to the consistency and reliability of the training process.
+II.3 Deep reinforcement learning
+Model structure
+The structure of our employed DRL model is illustrated in Fig.
+3.
+the DRL marries the perceptual power of deep learning with the decision-making and optimization capabilities of reinforcement learning, effectively overcoming their individual limitations and enhancing overall performance.
+Reinforcement Learning serves as the core mechanism for facilitating learning through the interaction between an agent and its environment.
+At each time step $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the necessary information for decision-making.
+Based on this observation, the agent selects and executes an action $a_{t}$ according to its policy $\pi$.
+The policy, serving as a strategy mapping states to actions, can be deterministic or probabilistic in nature.
+After the action is executed, the environment transitions to a new state $s_{t+1}$ and provides the agent with an immediate reward $r_{t+1}$.
+This reward, expressed as a numerical value, assesses the effectiveness of the action in progressing towards the agent’s objectives.
+Consequently, the agent incorporates this feedback-comprising both the reward and the new state—to refine its policy, utilizing Reinforcement Learning algorithms designed to optimize the accumulation of long-term rewards.
+Figure 3: DRL combines deep learning’s capability to interpret complex data with reinforcement learning’s strategy to make informed decisions.
+Proximal policy optimization
+PPO algorithm algorithm is an improvement over the policy gradient methods.
+The core idea of the PPO algorithm is to update the policy through two crucial steps: data sampling and policy optimization.
+In the data sampling phase, we interact with the environment using the current policy and collect a batch of trajectory data denoted as $\tau$, which can be mathematically expressed as:
+$\tau=(s_{0},a_{0},r_{1},s_{1},a_{1},r_{2},s_{2},\ldots,s_{T-1},a_{T-1},r_{T},s_{T}).$ (11)
+Here, $s_{t}$ represents the state at time step $t$, $a_{t}$ represents the action taken at time step $t$, and $r_{t}$ represents the reward received at time step $t$.
+The trajectory $\tau$ consists of a sequence of state-action-reward tuples observed during the interaction between the agent and the environment over a certain time horizon $T$.
+For a given trajectory $\tau$, at each interaction between the agent and the environment, the agent receives a reward value $r_{t}$ at time step $t$.
+Hence, the cumulative reward $R(\tau)$ of the trajectory $\tau$ is obtained by summing up the rewards obtained at each time step:
+$R(\tau)=\sum_{t=0}^{T}r_{t}.$ (12)
+Here, $T$ represents the time horizon or the length of the trajectory, and $r_{t}$ represents the reward received at time step $t$.
+The cumulative reward $R(\tau)$ provides a measure of the total reward obtained by the agent throughout the trajectory $\tau$.
+In the policy optimization phase, we use the sampled data obtained in the previous phase to update the policy.
+The method used to update the policy is gradient ascent.
+To perform gradient ascent, we compute the gradient of the PPO objective function with respect to the policy parameters.
+The objective function of the PPO algorithm is a surrogate objective designed to strike a balance between exploration and exploitation.
+It can be expressed as follows:
+$\left.L^{CLIP}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(r_{t}(\theta)\right)\hat{A}_{t},\operatorname{clip}\left(r_{t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{t}\right)\right],$ (13)
+where $\theta$ represents the policy parameters that are being optimized during the policy update process.
+$\hat{\mathbb{E}}_{t}$ denotes the empirical expectation, indicating that the expectation is estimated using the samples collected from the environment.
+$r_{t}(\theta)$ represents the ratio of the probabilities between the new policy and the old policy, evaluated at time step $t$.
+$\hat{A}_{t}$ corresponds to the advantage estimate at time step $t$.
+The advantage represents the advantage of taking a particular action compared to the expected value of the action under the current policy.
+The $\operatorname{clip}(.)$ function performs clipping on the policy ratio to ensure that the update does not deviate too much from the old policy.
+It restricts the ratio within the range $[1-\varepsilon,1+\varepsilon]$, where $\varepsilon$ is a small value.
+$[\min\left(r_{t}(\theta)\right)\hat{A}_{t},\operatorname{clip}\left(r_{t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{t}]$ expression represents the clipped surrogate objective.
+It consists of two terms: the first term is the minimum between $r_{t}(\theta)\hat{A}_{t}$ and $\operatorname{clip}\left(r_{t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{t}$, and the second term is the clipped version of $r_{t}(\theta)\hat{A}_{t}$.
+The purpose of this formulation is to balance policy improvement while ensuring a controlled update magnitude.
 II.4 DRL-enhanced active flow control
-This section outlines the design of a framework utilizing DRL algorithms to address AFC problems.
-Firstly, it is necessary to define two fundamental elements in DRL: the agent and the environment.
-The agent can be designed as algorithms suitable for solving continuous action control problems, such as PPO agent or SAC agent.
-The CFD simulation of the flow around an elliptical cylinder is chosen as the environment for the agent to interact with.
-Rabault et al.Rabault et al.
-(2020) presented the main structure of the custom environment class, encouraging readers to delve into the detailed content.
-This framework design applies DRL methods to solve the AFC problem, enabling the agent to continuously learn and optimize its behavior through interaction with the environment to achieve the control goal.
-Fig.
-4 illustrates the fundamental framework of DRL, the computational process of a single episode during training, and the formulas utilized.
-Figure 4: DRL leverages the concurrent execution of multiple environments within a parallel computational framework.
-Each iteration consists of two distinct phases: exploration and optimization.
-During the exploration phase, the agent engages with the environment via its current policy $\pi$, executing actions $a_{t}$, receiving states $s_{t}$ and rewards $r_{t}$, and noting terminal states.
-Upon the result of the iteration, the parallel computation across $n$ environments yields a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, with each trajectory $\tau_{i}$ comprising the sequence of states, actions, and rewards as $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1},r_{1}),\ldots,(s_{T},a_{T},r_{T})\}$, where $(s_{t},a_{t},r_{t})$ represents the state, action , and reward at time-step $t$, respectively, for the $i$-th environment.
-In the optimization phase, the agent samples a minibatch of trajectories from the experience replay buffer to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$.
-The parameters $\theta$ of the policy network are updated using gradient ascent, i.e., $\theta\leftarrow\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $\alpha$ denotes the learning rate.
-This process iterates, with the agent deploying the updated policy $\pi_{\theta^{\prime}}$ to interact with the environment, computing the advantage function $A(s_{t},a_{t})$, and recalculating the policy gradient for subsequent updates of $\theta$.
-Through iterative refinement, the policy network incrementally enhances its action selection strategy within the environment.
-  • State $s_{t}$: Traditional state spaces are often too expansive for effective policy learning, requiring operations in partially observable environments.
-As a result, an agent’s performance heavily relies on the quality and relevance of the data it can observe.
-The impact of partial observability of states $s_{t}$ on control results was mentioned in Viquerat et al.Viquerat et al.
-(2022)’s research.Integrating instantaneous velocity or pressure information from CFD simulations to characterize the environmental state $s_{t}$, the DRL agent can effectively perceive and analyze the current flow conditions.
-To achieve the control objectives of reducing drag and lift around the elliptical cylinder and controlling the shedding of vortices in the wake field, multiple probes are strategically placed around the elliptical cylinder and within its wake field.
-This method is advantageous for gathering comprehensive data pertaining to the flow dynamics around the cylinder and subsequent wake characteristics.
-The design of probes placement in this study was based on the design principles articulated by Rabault et al.Rabault et al.
-(2019), which advocate strategic positioning of probes around the cylinder and within the wake region of the cylinder.In their studies, Wang and XuWang and Xu (2024) and Wang et al.Wang et al.
-(2022a) all referred to the probes arrangement scheme proposed by Rabault et al.Rabault et al.
+Framework of DRL
+Our AFC framework utilizes DRL algorithms, consisting of the agent and the environment.
+The agent, designed for continuous action control, employs algorithms like PPO or SAC.
+The environment chosen for interaction with the agent is the CFD simulation of the flow around an elliptical cylinder.
+The custom environment class, as presented by Rabault et al.Rabault et al.
+(2020), provides the main structure, and further details can be explored in their work.
+Figure 4 illustrates the fundamental framework of DRL, including the computational process of a single training episode and the utilized formulas.
+Figure 4: The Proximal Policy Optimization algorithm is trained using a parallel computing framework to simultaneously run multiple environments.
+In the interaction stage between agent and environment, an agent interacts with its environments by following the current policy $\pi$, where it executes actions $a_{t}$, receives states $s_{t}$, and gathers rewards $r_{t}$, identifying terminal states in the process.
+This interaction generates a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, each $\tau_{i}$ being a sequence of state-action-reward tuples ending at a terminal state $T$.
+The agent uses these trajectories to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$ and applies gradient ascent to update the policy parameters $\theta$, aiming to maximize cumulative rewards.
+Through repeated iterations, leveraging updated policies and recalculated gradients, the agent continuously enhances its policy effectiveness, optimizing its performance across various environments.
+Fundamental components configuration
+In DRL-Enhanced active flow control, it involves state ($s_{t}$), action ($a_{t}$), reward ($r_{t}$), and parameter, which are described as below.
+The traditional state space $s_{t}$ poses challenges in learning policies due to its large size, emphasizing the reliance of the agent’s performance in partially observable environments on the quality and relevance of observed data, as highlighted by Viquerat et al.Viquerat et al.
 (2022).
-  • Action $a_{t}$ : The control action involves the blowing and suction of two synthetic jets on the elliptical cylinder.
-The flow rates of the two synthetic jets are constrained such that $Q_{\text{lower}}+Q_{\text{upper}}=0$.
-Therefore, the RL agent manipulates the velocity of the upper synthetic jets, while the lower synthetic jets has the same velocity but in the opposite direction.
-The velocity of the synthetic jets is set to be less than 2% of the inlet mean velocity.
-This restriction is imposed to prevent the agent from taking shortcuts and using excessively high jet velocities to achieve drag reduction, which contradicts the fundamental principle of AFC: energy efficiency.
-Furthermore, each control action lasts for 50 time steps, termed as a viscous action.
-Allowing faster actions could harm overall performance as they need time to influence the environment.
-Ensuring smooth transitions between current and updated actions is crucial.
-Let $a_{t}$ and $a_{t+1}$ represent the action magnitudes at consecutive time steps.
-To ensure a gradual transition between these actions and mitigate abrupt changes in jet velocity, a smoothing function $S$ can be defined mathematically.
-The function $S$ interpolates between the previous action magnitude $V_{\Gamma_{1},T_{i-1}}$ and the target action magnitude $a$ at time step $i$, resulting in an updated value $V_{\Gamma_{1},T_{i}}$.
-This smoothing function can be expressed as:
+To achieve the control objectives of reducing drag and lift and controlling vortex shedding in the wake field of the elliptical cylinder, strategically placed probes are utilized.
+The design of probe placement follows the principles proposed by Rabault et al.Rabault et al.
+(2019).
+This scheme was also adopted by Wang and XuWang and Xu (2024a) and Wang et al.Wang et al.
+(2022a) in their respective studies.
+The control action $a_{t}$ involves blowing and suction of two synthetic jets on the elliptical cylinder.
+The flow rates of the upper and lower synthetic jets are constrained such that their sum is zero, i.e., $Q_{\text{lower}}+Q_{\text{upper}}=0$.
+The RL agent controls the velocity of the upper synthetic jet, while the lower synthetic jet has the same velocity but in the opposite direction.
+To ensure energy efficiency in AFC, the velocity of the synthetic jets is limited to less than 2% of the inlet mean velocity, preventing the agent from utilizing excessively high jet velocities for drag reduction.
+Each control action, known as a ‘viscous action’, has a duration of 50 time steps.
+To ensure smooth transitions between consecutive actions, a smoothing function $S$ is introduced:
 $S(V_{\Gamma_{1},T_{i}},a,V_{\Gamma_{1},T_{i-1}})=V_{\Gamma_{1},T_{i}}+\beta\cdot(a-V_{\Gamma_{1},T_{i-1}}),$ (14)
@@
 This function ensures a smooth transition between consecutive actions, with the parameter $\beta$ controlling the smoothness of the transition.
-  • Reward $r_{t}$ : This study incorporates the average drag coefficient (without control), the time-averaged value of the drag coefficient, and the time-averaged value of the lift coefficient (averaged over the duration of the action).
-The primary objective is to minimize the drag while maintaining a low lift coefficient; hence, the weight of the drag coefficient in the reward function is set to 1, while the lift coefficient is assigned a weight $w$, typically chosen to be 0.1 or 0.2.
-This hyperparameter $w$ is determined through tuning.
-Including a constant (the average drag coefficient without control) in the reward function helps guide it towards positive rewards, which positively affects the convergence speed of the learning curve.
-The reward function $f(\cdot)$ translates the drag and lift coefficients into a numerical reward assessing the agent’s behavior.
+The reward $r_{t}$ evaluates the agent based on the average drag coefficient, time-averaged drag coefficient, and time-averaged lift coefficient during the action, with the objective of minimizing drag while keeping the lift coefficient low.
+The weight of the drag coefficient in the reward function is set to 1, indicating its importance, while the lift coefficient is assigned a weight $w$.
 The specific reward function takes the following form:
 $r_{T_{i}}=C_{D,0}-\left(C_{D}\right)_{T_{i}}-\omega\left|\left(C_{L}\right)_{T_{i}}\right|,$ (15)
-    $C_{D,0}$ represents the baseline drag coefficient, serving as a reference point for drag reduction.
+where $C_{D,0}$ represents the baseline drag coefficient, serving as a reference point for drag reduction.
 $(C_{D})_{T_{i}}$ denotes the drag coefficient at time step $T_{i}$, with the objective of minimizing this value relative to the baseline.
@@
 $\omega$ is a weighting factor that quantifies the trade-off between minimizing drag and controlling lift fluctuations.
-  • Parameter.
-Time granularity in the DRL framework is crucially set based on the fluid dynamics problems studied and significantly affects agent performance and learning task complexity.
-Precise adjustments of this hyperparameter, informed by prior research and our ongoing strategy refinement, are vital for achieving optimized control outcomes.
-Each action is applied for a duration of 0.025 seconds ( Corresponds to 25 numerical simulation time steps ), and each episode consists of 100 control actions.
-The maximum total time for an episode is defined as $T_{\max}=T_{100}=2.5$ nondimensional time units.
-The shedding frequency of vortices around the elliptical cylinder in this study ranges from 0.3 to 0.4, corresponding to 6.25 to 8.33 vortex shedding cycles per episode.
-The length of episodes is intentionally chosen to encompass multiple shedding cycles, facilitating effective learning of the control algorithm by the agent through observation and adaptation of flow dynamics over several cycles.
-II.5 High-Performance Computing in DRL and CFD
-The DRL training and numerical simulation in this study were conducted on a high-performance computing (HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60 GHz and comprising a total of 64 cores distributed evenly across two sockets, each containing 32 cores.
-Numerical simulations were performed using the eighth version of the open-source software platform, OpenFOAM® version 8.Jasak et al.
-(2007) The deployment of DRL algorithms was based on the Tensorforce library.Schaarschmidt et al.
-(2018) The ANN definitions and gradient descent algorithms were facilitated using the TensorFlow open-source library.Abadi et al.
-(2016) For DRL implementation, agent was loaded using the Tensorforce platform with the PPO algorithm, and the environment was constructed using the Gym interface.Brockman et al.
-(2016) Fig.
+The time granularity parameter significantly affects the agent’s performance and the complexity of the learning task.
+Precise adjustment to this hyperparameter, based on prior research and ongoing strategy refinement, is essential for achieving optimized control outcomes.
+In particular, each action in our analysis is applied for a duration of 0.025 seconds, which corresponds to 25 numerical simulation time steps.
+Each episode consists of 100 control actions, resulting in a maximum total time for an episode of $T_{\max}=T_{100}=2.5$ nondimensional time units.
+The shedding frequency of vortices around the elliptical cylinder in this study ranges from 0.3 to 0.4, which corresponds to 6.25 to 8.33 vortex shedding cycles per episode.
+The length of episodes is intentionally chosen to encompass multiple shedding cycles, which facilitates effective learning of the control algorithm by allowing the agent to observe and adapt to the flow dynamics over several cycles.
+Computing specifications
+The DRL training and numerical simulation are conducted on a high-performance computing (HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60 GHz, and comprising a total of 64 cores evenly distributed across two sockets, each containing 32 cores.
+Specifically, the numerical simulation is performed using version 8 of the open-source software platform OpenFOAM® version 8 Jasak et al.
+(2007).
+The deployment of DRL algorithms is based on theTensorforce library Schaarschmidt et al.
+(2018), while the definitions of artificial neural networks (ANNs) and gradient descent algorithms are facilitated using the TensorFlow open-source library Abadi et al.
+(2016).
+For the DRL implementation, the agent is loaded using the Tensorforce platform with the PPO algorithm, and the environment is constructed using the Gym interface Brockman et al.
+(2016) .
+Fig.
 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process.
-This parallel strategy was inspired by the research of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al.
-(2022a), and also incorporated insights from the work of Wang et al.Wang et al.
+This parallel strategy is inspired by the studies of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al.
+(2022a), and also incorporateds insights from the work of Wang et al.Wang et al.
 (2022a) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.
@@
 5 depicts the vortex shedding patterns within the wake field of elliptical cylinders with $Ar$ from 0 to 2.
-Starting from Fig.
-5(a) depicting an elliptical cylinder with $Ar=2$, the wake field does not exhibit vortex shedding.
+For an elliptical cylinder with $Ar=2$, the wake field does not exhibit vortex shedding.
 This is expected as the elongated elliptical cylinder results in weak adverse pressure gradients that allow the boundary layers to stay attached for a prolonged distance.
@@
 For a stable flow field with no vortex shedding, we expect the DRL-based flow control algorithm to conform to the underlying flow physics and to not introduce any actuation since the current baseline is essentially the optimal flow configuration.
-To the best of our knowledge, this sanity test that verifies minimal DRL-based control is often overlooked in literature.
-We include this to showcase the robustness of DRL-based control across a spectrum of flow physics.
-We delay the discussion of DRL control to Sec.
-III.2.
+Based on our current knowledge, this sanity test that verifies minimal DRL-based control is often overlooked in the literature.
+We include it to emphasize the robustness of DRL-based control across a spectrum of flow physics.
+The detailed discussion of DRL control will be presented in Sec.
+III.2, where we will provide a comprehensive analysis of its principles and application.
 Figure 5: Contours of instantaneous vorticity for flow around an elliptic cylinder at $Re=100$, varying from $Ar=2$ to $Ar=0$.
 Purple and orange indicate positive and negative vorticity, respectively.
-In all the depicted vorticity contours, purple represents the clockwise vortices shed from the upper side of the elliptical cylinder, while orange represents the counterclockwise vortices shed from the lower side.
+Purple represents the clockwise vortices shed from the upper side of the cylinder, while orange represents the counterclockwise vortices shed from the lower side.
+(a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1.0$; (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (g) $Ar=0$.
 As the $Ar$ decreases to 1.75, the recirculation region behind the cylinder starts to become unstable and begins to exhibit vortex shedding phenomena.
@@
 This serves as a foundation for subsequent flow control studies.
-Velocity fluctuation and design of probe layout
-In addition to the baseline flow field, we analyze the fluctuating component of velocity, which serves as a good indicator to inform the location of probes during control.
-The judicious incorporation of fluid physics knowledge within the control agents can streamline their architecture without compromising performance, as exemplified by Pastoor et al.Pastoor et al.
-(2008), where phase control in drag reduction of a bluff body is primarily founded upon the understanding of decoupled shear layer development and wake processes.
-We also utilize the physical information from the wake field to guide the flow control strategies of the agents.
-This aspect is particularly reflected in the practical design of the probe positions.
-The study by Protas and Wesfreid Protas and Wesfreid (2002) elaborates on the adjustable component of the drag coefficient, $C_{D}^{0}$, which reflects the influence of oscillatory flow phenomena on the drag coefficient.
-This provides a pathway for strategic manipulation of the von Kármán vortex street to adjust the drag through targeted interventions.
-Building on this foundation, we tailor the physical information observed by the agents to specifically face the factors contributing to instability in the wake field.
-Figure 6: The distribution of probes within the fluctuating velocity magnitude fields for different $Ar$ : (a) $Ar=0.25$; (b) $Ar=0.1$; (c) $Ar=0$.
+Velocity fluctuation and probe layout optimization
+In addition to investigating the baseline flow field, we conduct an analysis of the velocity fluctuation component, which serves as a valuable guiding indicator for probe placement during control.
+By integrating our knowledge of fluid physics into the control agents, we can optimize their architecture while maintaining superior performance.
+An illustrative example of this approach is demonstrated in the work of Pastoor et al.Pastoor et al.
+(2008), where phase control for drag reduction of a bluff body is achieved by using the decoupled development of shear layers and wake processes.
+Similarly, we utilize the physical information from the wake field to guide the flow control strategies of the agents, with a focus on the practical design of probe positions.
+Moreover, Protas and Wesfreid Protas and Wesfreid (2002) highlighted the influence of oscillatory flow phenomena on the drag coefficient, represented by the adjustable component $C_{D}^{0}$.
+This understanding provides a strategic pathway for manipulating the von Kármán vortex street and achieving drag adjustment through targeted interventions.
+Expanding on this foundation, we tailor the physical information observed by the agents to specifically face the factors contributing to instability in the wake field.
+Figure 6: The distribution of probes within the fluctuating velocity magnitude fields for different $Ar$ : (a) $Ar=1$; (b) $Ar=0$.
+To provide a clearer visualization of the probe distribution, we display to a specific region within the computational domain.
+The x-axis ranges from -2D to 10D, and the y-axis ranges from -2D to 2.1D, where D represents a reference length scale.
 As depicted in Fig.
-6, the fluctuating magnitude velocity fields around elliptical cylinders with $Ar$ of 0, 0.1, and 0.25 are illustrated.
+6, the fluctuating magnitude velocity fields around elliptical cylinders with $Ar$ of 0 and 1 are illustrated.
 The positions of the black dots on the diagram represent the locations of the probes.
@@
 The extraction of physical information from these locations enables the agents to acutely observe changes in the flow field.
+In addition, Fig.
+6 illustrates two tests with different $Ar$, the configurations for other $Ar$ are similar and not further elaborated upon.
+III.2 Effectiveness of AFC
+Using the DRL algorithm, we investigate the performance of actively controlling the magnitude of synthetic jets for targeted flow control throughout the process of elliptical cylinders transitioning from elongated bodies to circular cylinders, and finally to a normal flat plate, with aspect ratios ranging from 2 to 0 at $Re=100$.
+The evaluation of the control’s effectiveness and performance is based on several criteria: intensity of vortex shedding after control, reductions in $C_{D}$, weakening of $C_{L}$, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
+For a detailed demonstration of vortex suppression, we present in Fig.
+7 an example using an elliptical cylinder with $Ar=0.75$, observing the evolution of the instantaneous vorticity field throughout the DRL training process.
+We select four representative snapshots of the instantaneous vorticity field to depict changes in vortex shapes and positions.
+Labels A,B,C,D and E in Fig.
+7(a) denote the five vortices initially shed.
+By the time depicted in Fig.
+7(b), these initial five vortices have progressed downstream without new vortices being shed under the influence of flow control.
 In Fig.
-6, three tests with different $Ar$ are shown as examples.
-The configurations for other $Ar$ are fundamentally similar and are not elaborated further.
-III.2 Effectiveness of AFC
-This section describes the performance of AFC using the DRL algorithm, which manipulates the magnitude of synthetic jets to achieve targeted flow control.
-At $Re=100$, AFC is applied to elliptical cylinders with $Ar$ ranging from 2 to 0, from elongated body to circular cylinder and then to a normal flat plate.
-The evaluation of the control’s effectiveness and performance is based on several criteria: intensity of vortex shedding after control, reductions in drag coefficient, variations in lift coefficient, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
-Fig.
-7 and Fig.
-8 display the velocity magnitude snapshots of elliptical cylinders with $Ar$ ranging from 2 to 0 under baseline and controlled conditions.
-More specifically, Fig.
-7 shows those for $Ar$ from 2 to 0.75, the vortices in the wake field of the controlled elliptical cylinders can be completely suppressed.
-In contrast, Fig.
-8 illustrates those for $Ar$ from 0.5 to 0, the vortices in the wake field of the controlled elliptical cylinders are mitigated by not fully suppressed.
+7(c), labeled as A,B,C,D and E, the vortices have developed near the exit of the computational domain, exhibiting elongated shapes as they are about to be expelled from the domain.
+At this stage, the shape of the recirculation bubble has stabilized, and no new vortices are shed.
+By the time represented in Fig.
+7(d), the originally labeled five vortices have all been expelled from the computational domain, the area of the recirculation bubble no longer expands, and the entire computational domain has reached a stable state.
+Under the conditions of $Ar=1.75$, 1.5, 1.25, and 1 for the elliptical cylinder, the DRL-based AFC demonstrates similar control effectiveness.
+Figure 7: For an elliptical cylinder with $Ar=0.75$, observed the evolution of the instantaneous vorticity snapshots when DRL training.
+Specifically, focused on four representative training episodes: (a) Episode = 0, (b) Episode = 450, (c) Episode = 800, and (d) Episode = 2100.
+The initial shedding vortex are labeled as A, B, C, D and E, and them deformation and displacement are subsequently tracked.
+The variations in the instantaneous vorticity snapshots at these episodes describe the influence of flow control techniques on vortex shedding.
 Figure 8: Comparison of the velocity magnitude snapshots for an elliptical cylinder before and after control.
@@
 (a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$.
-Starting from Fig.
-7(a) for the elliptical cylinder with $Ar=2$, the baseline flow field (left) exhibits high stability with no observable vortex shedding.
-We applied a DRL-based flow control to this stable, non-shedding wake field, and the controlled flow field (right) behind the elliptical cylinder remained stable and nearly identical to the baseline.
-This sanity test indicates that the DRL-based model is consistent with flow physics and does not generate any unreasonable perturbation during the control process.
-As the $Ar$ decreases to $Ar=1.75$, the application of DRL-based AFC technology successfully suppresses the vortices that originally separated alternately from both sides of the elliptical cylinder.
-Further reducing the $Ar$ results in a more pronounced and frequent shedding of vortex structures in the uncontrolled wake, while the recirculation bubble downstream of the elliptical cylinder gradually decreases and becomes oscillatory.
-Although the flow instability gradually increases, under the conditions of $Ar=1.5$, 1.25, 1, and 0.75, the alternating vortex streets behind the elliptical cylinders are transformed into a stable wake field after the implementation of AFC in all cases.
-The DRL-based AFC technology demonstrates significant vortex shedding suppression capabilities under these conditions, highlighting the effectiveness of the AFC control strategy in enhancing the flow stability of the wake field.
+To demonstrate the suppression effect of AFC technology on vortex shedding phenomena, velocity magnitude snapshots of elliptical cylinders with $Ar$ ranging from 2 to 0.75 and from 0.5 to 0 are displayed in Fig.
+8 and Fig.
+9, respectively.
+As shown in Fig.
+8, the DRL-based AFC approach demonstrates consistent and effective suppression of vortex shedding and enhancement of flow stability in the wake field.
+At $Ar$ = 2, the baseline flow field is stable without vortex shedding, and the controlled flow field remains stable and similar to the baseline.
+As $Ar$ decreases, the AFC successfully suppresses vortices that were previously formed on both sides of the elliptical cylinder.
+Despite increased vortex shedding and oscillatory behavior in the uncontrolled wake, AFC implementation at $Ar$ values of 1.5, 1.25, 1, and 0.75 transforms the alternating vortex streets into a stable wake field.
+Subsequently, as the $Ar$ continuously decreases, the baseline flow becomes more unstable, characterized by a diminishing recirculation region and increased vortex shedding intensity and frequency.
+This phenomenon is visually depicted in Fig.
+9.
+In particular, at $Ar$ = 0.5, the application of flow control successfully suppresses the alternating vortex shedding, resulting in an elongated recirculation region.
+However, the wake flow still exhibits oscillations.
+For $Ar$ = 0.25, flow control elongates the recirculation bubble similar to the $Ar$ = 0.5 case, but with prominent oscillations.
+Vortex shedding persists in an alternating manner, resembling elongated and stretched vortex structures.
+Although complete suppression of vortex shedding is not achieved at $Ar$ = 0.1 or 0 due to the inherent chaotic nature of the wake behind vertical thin bodies, the recirculation bubbles are elongated, and the initiation point of shedding is shifted further downstream.
+The controlled flow exhibits reduced shedding frequency and intensity compared to the baseline conditions.
+Fig.
+8 illustrates the effectiveness of AFC technology in eliminating vortex shedding and transitioning flow from unstable to stable modes within the range of $Ar$ from 2 to 0.75, while Fig.
+9 indicates that at lower $Ar$ values (0.5 to 0), this technology can effectively alleviate vortex shedding and enhance flow stability to some extent.
 Figure 9: Comparison of the velocity magnitude snapshots for an elliptical cylinder before and after control.
@@
 (a) $Ar=0.5$; (b) $Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
-In Fig.
-8, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the baseline flow becomes increasingly unstable, and the area of the recirculation region gradually decreases while the strength and frequency of vortex shedding increase.
-When $Ar=0.5$, the frequent alternating vortex shedding is effectively suppressed after flow control is applied.
-The recirculation region behind the elliptical cylinder is significantly elongated, and the shedding of vortices is controlled, although the wake flow still exhibits alternating oscillations.
-For the controlled flow at $Ar=0.25$, the recirculation bubble behind the elliptical cylinder is elongated, resembling the case at $Ar=0.5$.
-However, there is a notable difference: the recirculation bubble now demonstrates prominent oscillations.
-The vortex shedding behind the cylinder continues alternately, akin to the elongated vortex structures being stretched longitudinally.
-Despite these instabilities, vortex shedding in the wake fields at $Ar=0.5$ and $Ar=0.25$ is reasonably mitigated compared to the baseline state.
-At $Ar=0.1$ or 0, compared to baseline conditions, the recirculation bubbles behind the cylinder are elongated, and the initiation point of vortex shedding is extended further downstream.
-We do not observe complete suppression of vortex shedding due to the highly chaotic nature of these wake behind vertical thin bodies.
-Nevertheless, the vortex shedding is reduced to a lower frequency and intensity.
-These observations reflect the effectiveness of the control strategies in suppressing vortex shedding and stabilizing the wake field.
 Figure 10: Comparison of $C_{D}$ and $C_{L}$ before and after control.
@@
 (a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
-Fig.
-9 presents a comparative analysis of the $C_{D}$ and $C_{L}$ before and after control intervention over $Ar$ ranging from 2 to 0.75.
-At an $Ar$ of 2, both $C_{D}$ and $C_{L}$ are stable, indicating that AFC does not have any effect.
-However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, upon activating synthetic jets control, there is an immediate, sharp response in $C_{D}$, which quickly descends to a value near its minimum.
-After half the total control duration, $C_{D}$ stabilizes and remains constant at this minimal value.
-Similarly, $C_{L}$ experiences significant fluctuations at the start of control but stabilizes near zero after half the total control duration.
-The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ are detailed in Tab.
+Further, Fig.
+10 and Fig.
+11 provide a comparative analysis of the $C_{D}$ and $C_{L}$ before and after implementing the DRL-based AFC strategy, showcasing its effectiveness across a range of $Ar$ from 2 to 0.75 and from 0.5 to 0, respectively.
+As shown in Fig.
+10, the AFC strategy has no effect on the stability of both $C_{D}$ and $C_{L}$ at an $Ar$ of 2.
+However, when synthetic jets control is activated at $Ar$ values of 1.75, 1.5, 1.25, 1, and 0.75, there is an immediate and substantial reduction in $C_{D}$, which quickly reaches a minimum value.
+After half of the total control duration, $C_{D}$ stabilizes and remains constant at this minimum value.
+Similarly, $C_{L}$ experiences significant fluctuations initially but stabilizes near zero after half of the total control duration.
+The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ values can be found in Tab.
 2.
-These observations substantiate the effectiveness of the AFC strategy underpinned by DRL algorithms.
-The activation of control swiftly alters the periodic behavior of both $C_{D}$ and $C_{L}$, facilitating the stabilization of $C_{D}$ and $C_{L}$ in the subsequent stages.
-This demonstrates the capability of the DRL-based AFC approach to significantly reduce the drag around an elliptical cylinder to its lowest feasible value and diminish the lift to zero.
-Fig.
-10 illustrates the $C_{D}$ and $C_{L}$ for both baseline and controlled flow conditions around an elliptical cylinder at $Ar$ of 0.5, 0.25, 0.1, and 0.
-Under AFC, at $Ar=0.5$ and $Ar=0.25$, both $C_{D}$ and $C_{L}$ are reduced compared to the baseline flow, yet they do not achieve the stable state observed at $Ar=0.75$ and higher.
-Initially, both the $C_{D}$ and the $C_{L}$ experienced abrupt changes.
-However, they did not fully stabilize in the subsequent period.
-In the latter half of the control duration, $C_{D}$ and the $C_{L}$ exhibited fluctuations, but both the amplitude and frequency of these fluctuations were lower than those observed in the baseline flow.
-At $Ar=0.1$ and $Ar=0$, the controlled flow experiences a sharp decrease in the drag coefficient at the onset of control, followed by oscillations that persist at a lower frequency and amplitude than the baseline flow.
-Although complete stability is not achieved, the average value of the $C_{D}$ is significantly reduced.
-Similarly, the $C_{L}$ undergoes an initial sharp decrease and approaches near zero, then later starts to oscillates, but at a lower frequency than the baseline $C_{L}$ oscillation.
+Clearly, the periodic oscillations of $C_{D}$ and $C_{L}$ originally present in the baseline flow are significantly reduced, ultimately transitioning to a stable state.
 Figure 11: Comparison of $C_{D}$ and $C_{L}$ before and after control.
 The background color indicates that the drag or lift coefficient is still in an unstable stage after flow control.
-(a) $Ar=0.5$; (b)$Ar=0.25$; (c) $Ar=0.1$; (d) $Ar=0$.
-Figure 11: Action, which is the mass flow rate of the synthetic jets.
-The background color represents that the value of the action is in an unstable stage, and the absence of a background color means that the action has been in a stable stage during this period.
-(a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$.
-(f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
-Fig.
-11 illustrates the time history curve of the action (the mass flow rate of synthetic jets), When performing AFC using DRL for flow around an elliptical cylinder with $Ar$ between 0 and 2.
-When $Ar=2$, the AFC is negligible, and the value of the action, pertaining to the control input, remains close to zero.
-This is a strong indication that the DRL-based model is not generating nonphysical perturbations for a stable flow field.
-From $Ar=1.75$ to $Ar=0.75$, during the initial phases of control, the agent selects larger action values, corresponding to higher mass flow rates of synthetic jets.
-This indicates a deliberate strategy to assert flow control by introducing significant external energy inputs.
-As control progresses, the agent gradually diminishes the action values, thereby reducing the mass flow rate of the synthetic jets.
-Subsequently, in the mid to later stages, the agent’s action values converge towards zero, corresponding to a near-zero mass flow rate of synthetic jets.
-This behavior signifies the agent’s optimization towards an energy-saving state, maintaining flow characteristics with minimal control input.
-When $Ar=0.5$ and $Ar=0.25$, the value of the action initially surges but does not stabilize near zero in the later stages.
-For $Ar=0.1$ and $Ar=0$, the action value remains in a state of periodic fluctuation.
-Overall, the results presented in Fig.
-7 and Fig.
-9 illustrate that the DRL-based control strategy exhibits energy efficiency and effectiveness within the range of elliptical $Ar=2$ to $Ar=0.75$.
-By progressively decreasing the mass flow rate of synthetic jets, the strategy achieves flow control, reduces drag and lift forces, and does so with reduced external energy input.
-However, at $Ar=0.5$ and $Ar=0.25$, the energy efficiency of the AFC strategy diminishes.
-At $Ar=0.1$ and $Ar=0$, the external energy utilized by the control strategy is relatively high compared to the inlet velocity, which will be detailed in Tab.
-2.
-Figure 12: When $Ar$ of the elliptical cylinder is in the range of 2-0, the learning curve of PPO algorithm training.
-(a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$.
-(f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
-In Fig.
-12, we observe the learning curve during training of flow around elliptical cylinders with different $Ar$.
-At $Ar=2$, the learning curve has entered the plateau phase after the initial improvement phase, indicating that the DRL training has converged.
-For $Ar=1.75$ and $Ar=1.5$, the reward function initially increases from negative to positive values and then after 4,000 episodes converges near the maximum reward value.
-At $Ar=1.25$, despite some initial fluctuations, the training reaches a relatively stable state after 3,000 episodes and eventually settles into a stable plateau.
-For $Ar=1$ and $Ar=0.75$, there are significant initial fluctuations with the learning curve rising swiftly, indicating rapid progress in policy learning.
-After 2,000 episodes, the training demonstrates a stable convergence trend, with minor fluctuations in the reward value between 2,000 and 5,000 episodes.
-This shows that the PPO algorithm has successfully learned an effective strategy, exhibiting commendable convergence and stability in performance.
-At $Ar=0.5$, the speed of training convergence begins to slow, trending towards convergence after 4,000 episodes, with the reward values stabilizing between 4,000 and 10,000 episodes.
-At $Ar=0.25$, training approaches convergence after 8,000 episodes, reaching a total of 20,000 episodes, with a mid-to-late plateau marking stable convergence.
-At $Ar=0.1$ and $Ar=0$, the training convergence is extremely slow, with signs of convergence only appearing after nearly 12,000 episodes.
-The DRL algorithm exhibits rapid initial convergence and reaches a stable plateau during the mid-to-late stages for $Ar=1.25$, 1 and 0.75, showcasing excellent convergence and stability throughout the training process.
-Training also achieves stable convergence within 5,000 episodes for $Ar=1.5$ and $Ar=1.75$.
-However, for $Ar=0.5$, 0.25, 0.1 and 0, as the flow phenomena become increasingly complex, the control challenge intensifies, leading to a noticeable slowdown in convergence speed.
-This indicates that for flow scenarios with higher degrees of instability, the convergence performance of DRL training is compromised, requiring a larger number of training iterations to reach stable convergence.
-Table 2: The control effect of elliptical cylinder under DRL-based AFC.
+(a) and (b) $Ar=0.5$; (c) and (d) $Ar=0.25$; (e) and (f) $Ar=0.1$; (g) and (h) $Ar=0$.
+The impact of the AFC strategy on $C_{D}$ and $C_{L}$ at $Ar=0.5$ and $Ar=0.25$ is described in Fig.
+11.
+While both coefficients experience reductions compared to the baseline flow, they do not achieve the same level of stability observed at $Ar=0.75$ and higher.
+Initially, there are abrupt changes in both $C_{D}$ and $C_{L}$, but they do not fully stabilize in the subsequent period.
+In the latter half of the control duration, there are fluctuations in both coefficients, but with lower amplitude and frequency compared to the baseline flow.
+At $Ar=0.1$ and $Ar=0$ are described in also in Fig.
+11, the controlled flow exhibits a sharp decrease in drag coefficient at the beginning of the control, followed by oscillations at a lower frequency and amplitude compared to the baseline flow.
+Although complete stability is not achieved, the average value of $C_{D}$ is significantly reduced.
+Similarly, $C_{L}$ initially decreases sharply and approaches zero, then exhibits oscillations at a lower frequency than the baseline $C_{L}$ oscillations.
+Table 2: Quantified control effects of elliptical cylinders under DRL-based AFC.
 $Ar$ Re $\overline{C}_{D,\text{Baseline}}$ $\overline{C}_{D,\text{Controlled}}$ Drag Reduction (%) $\overline{C}_{L,\text{Baseline}}$ $\overline{C}_{L,\text{Controlled}}$ Lift Reduction (%) $\overline{a}$ $\overline{a}$ ratio (%)
@@
 1.25 100 2.918 2.804 3.9 0.541 0.006 98.9 0.006 0.6
-1 100 3.207 2.951 8.0 1.022 0.011 98.9 0.001 0.1
+1.00 100 3.207 2.951 8.0 1.022 0.011 98.9 0.001 0.1
 0.75 100 3.792 3.197 15.7 1.603 0.076 95.2 0.010 1.0
@@
 0 100 7.679 4.331 43.6 1.513 0.605 60.0 0.521 52.1
-  • a. $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged drag coefficient of the elliptical cylinder once the baseline flow has fully developed.
-  • b.
+The DRL-based AFC implemented in our study is demonstrated across Figs.
+8, 9, 10 and 11.
+This control strategy not only achieves high control performance but also maintains robust effectiveness, as evidenced by the quantified variations in $C_{D}$, $C_{L}$, and the ratio of synthetic jet mass flow rate to incoming flow rate, presented in Tab.
+2.
+$\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged $C_{D}$ and $C_{L}$ of the elliptical cylinder once the baseline flow has fully developed.
 $\overline{C}_{D,\text{Controlled}}$, $\overline{C}_{L,\text{Controlled}}$, and $\overline{a}$ denote the time-averaged values under stable conditions after AFC is applied, excluding the transient response phase.
-  • c. $\overline{a}$ ratio (%) represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the inlet.
-Tab.
-2 presents the effectiveness of AFC techniques based on DRL algorithms, targeting elliptical cylinders with $Ar$ ranging from 0 to 2.
-It meticulously details the values of $C_{D}$ and $C_{L}$ before and after the application of control, along with their respective rates of change.
-Additionally, it describes the magnitude of the mass flow rate of the synthetic jets utilized for control and its ratio relative to the inflow rate.
-When the $Ar$ is set to 2, the baseline flow has reached a steady state.
-During the application of flow control, the agent does not use external energy to disrupt this stable flow field, which demonstrates the intelligence of the DRL-based AFC.
-This reflects the adaptability and intelligent decision-making of AFC control in maintaining flow stability.
-For elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the $C_{D}$ reduction rates are 0.9%, 2.1%, 3.9%, 8.0%, and 15.7% respectively, while the $C_{L}$ reduction rates exceed 95% in each case.
-These results highlight the efficacy of the DRL-based AFC technology in effectively suppressing both $C_{D}$ and $C_{L}$ within the $Ar$ range of 1.75 to 0.75, demonstrating the efficiency of the DRL-based intelligent control strategy.
-However, as $Ar$ decreases to 0.5, 0.25, 0.1, and 0, even though the $C_{D}$ reduction rates significantly increase to 26.9%, 35.9%, 37.6%, and 43.6%, the $C_{L}$ reduction rates are 68.0%, 66.1%, 50.2%, and 60.0% respectively.
-This indicates that while the DRL-based control strategy achieves significant reductions in $C_{D}$, its effect on suppressing $C_{L}$ is limited and does not completely eliminate $C_{L}$ around the elliptical cylinder.
-As the $Ar$ further decreases, the AFC technique based on the DRL algorithm does not fully suppress vortex shedding behind the elliptical cylinder, but merely delays the shedding process.
-The reduction in $C_{L}$ coefficients aligns with the flow snapshots presented in Fig.
-8, showing that vortex shedding still occurs, albeit with controlled frequency and intensity.
-These phenomena suggest that at very low $Ar$ values, the control strategy faces challenges in managing more complex and unstable flow conditions.
-[…]
-In their research, Protas and WesfreidProtas and Wesfreid (2002) elucidated that the drag coefficient consists of two components: (i) the invariant drag ($C_{D}^{b}$), which remains constant at a given Reynolds number; and (ii) the amendable drag correction ($C_{D}^{0}$), associated with oscillatory flow phenomena within the fluid, representing the additional resistance arising from fluid oscillations or fluctuations.
-This portion of drag can be adjusted through external interventions, indicating the influence of unstable flow regions and the potential for control.
-In our research, the flow field is controlled by utilizing external excitation, such as synthetic jets, to mitigate the additional resistance caused by instability, thereby optimizing the overall fluid dynamic performance.
-As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, thereby enhancing the controllable component of the drag.
-This is the reason why the drag reduction rate gradually increases as $Ar$ decreases in Tab.
+At $Ar=2$, the results of AFC control still maintain a stable baseline flow, without applying external excitation to disrupt flow stability.
+In the range of $Ar$ values from 1.75 to 0.75, the AFC control achieves substantial reductions in $C_{D}$ ranging from 0.9% to 15.7%, while the reduction rates in $C_{L}$ exceed 95% in all cases.
+Notably, the reduction in the coefficient of $C_{L}$ for elliptical cylinders, ranging from 99.7% to 95.2%, aligns with the achievement of a completely stable state (free from oscillations and fluctuations) for both $C_{D}$ and $C_{L}$ as illustrated in Fig.
+10.
+This consistency is also mirrored in the suppression of alternating vortex shedding in the Baseline, as depicted in Fig.
+8.
+Figure 12: Action, which is the mass flow rate of the synthetic jets.
+The background color represents that the value of the action is in an unstable stage, and the absence of a background color means that the action has been in a stable stage during this period.
+(a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$; (f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
+Interestingly, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the $C_{D}$ significantly increase to 26.9%, 35.9%, 37.6%, and 43.6% respectively, while the reduction rates of the $C_{L}$ are observed at 68.0%, 66.1%, 50.2%, and 60.0%.
+Protas and WesfreidProtas and Wesfreid (2002) delineated the instantaneous drag coefficient into two components: the invariant drag ($C_{D}^{b}$) and the amendable drag ($C_{D}^{0}$).
+The $C_{D}^{b}$ remains constant at a given Reynolds number, while the $C_{D}^{0}$ is related to oscillatory flow phenomena, representing the additional resistance caused by fluid oscillations.
+This amendable part can be adjusted through external interventions, demonstrating the potential for flow control.
+To mitigate the $C_{D}^{0}$ caused by flow instabilities, we utilize external excitations such as synthetic jets to manipulate the flow field, thereby optimizing the overall fluid dynamics performance.
+As the $Ar$ of the elliptical cylinder decreases, the pulsations, oscillations, and instabilities within the wake intensify, enhancing the controllable component of the drag.
+This explains why the drag reduction ratio gradually increases with the decrease in $Ar$ as shown in Tab.
 2.
+However, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the $C_{L}$ are significantly lower than when $Ar$ is greater than or equal to 0.75.
+Within this range of $Ar$, only 50.2% to 68% of the $C_{L}$ is reduced, consistent with the fact that neither the $C_{L}$ nor the $C_{D}$ achieve the complete stability observed in Fig.
+11.
+Additionally, the baseline instability is not fully mitigated as shown in Fig.
+9.
+This highlights the fact that there is still room for improvement in the current DRL-based AFC framework for controlling the flow instability of elliptical cylinders when $Ar$ is less than or equal to 0.5.
+Addressing this challenge remains a significant hurdle in the field of active flow control technology for slender structures.
+Fig.
+12 illustrates the time history curve of the action (the mass flow rate of synthetic jets) during the application of DRL for AFC in the flow around an elliptical cylinder with $Ar$ ranging from 0 to 2.
+At $Ar=2$, the AFC has minimal impact, with the action value close to zero, indicating stable flow control without nonphysical perturbations.
+From $Ar=1.75$ to $Ar=0.75$, the agent employs a deliberate strategy of selecting larger action values for higher mass flow rates, aiming to assert flow control through significant energy inputs.
+As the control progresses, the action values gradually decrease, reducing the mass flow rate.
+In the later stages, the action values converge towards zero, optimizing for energy efficiency while maintaining flow characteristics.
+However, at $Ar=0.5$ and $Ar=0.25$, stable flow control is challenging, as the mass flow rate does not stabilize near zero.
+The non-zero action values indicate the difficulty in suppressing the chaotic wake for vertical slender bodies under these conditions.
+For $Ar=0.1$ and $Ar=0$, the action values exhibit periodic fluctuations, showcasing the DRL agent’s ability to recognize and exploit dominant flow frequencies.
+Furthermore, in Tab.
+2, we normalize the mass flow rates of the synthetic jets against the incoming far-field flow rate, reflecting their relative strength compared to the incoming flow.
+$\overline{a}$ ratio (%) represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the far-field inflow boundary.
+For $Ar$ greater than 0.75, the ratio of synthetic jets to the inflow is less than 1%, indicating that the mass flow rate utilized in the flow control strategy proposed by the PPO agent is minimal, in line with the energy efficiency requirements of AFC technologies.
+For $Ar$ ranging from 0.75 to 1.75, the elliptical cylinder achieves a reduction in $C_{D}$ and attenuates up to 95% of $C_{L}$ with very low energy consumption from external excitation.
+However, as $Ar$ decreases from 0.5 to 0.1, the ratio of synthetic jets to the inflow increases significantly.
+Despite this increase in energy input, the capability to attenuate $C_{L}$ decreases.
+This indicates that at lower $Ar$ values, the instability in the wake field cannot be fully controlled, even with higher energy from external excitation.
 The control strategies derived from DRL-based AFC technology exhibit greater complexity compared to conventional harmonic forcing implemented in previous works (Bergmann, Cordier, and BrancherBergmann, Cordier, and Brancher (2005)), highlighting the value of employing ANN as controllers.
@@
 Furthermore, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.
+Figure 13: When $Ar$ of the elliptical cylinder is in the range of 2-0, the learning curve of PPO algorithm training.
+(a) $Ar=2$; (b) $Ar=1.75$; (c) $Ar=1.5$; (d) $Ar=1.25$; (e) $Ar=1$.
+(f) $Ar=0.75$; (g) $Ar=0.5$; (h) $Ar=0.25$; (i) $Ar=0.1$; (j) $Ar=0$.
+The above content details the effectiveness of the flow control strategy, with the corresponding DRL training process curves for elliptical cylinders with different $Ar$ values presented in Fig.
+13.
+First, we observed the overall trend of the reward function.
+For $Ar$ values ranging from 1.75 to 0, we noted that the reward function increases with the number of training iterations, eventually converging to a stable value.
+However, for $Ar=2$, after 5,000 training iterations, the reward function reaches a plateau but remains below zero, indicating that although the reward function has converged, the agent is unable to achieve the desired objective.
+Next, we examined the convergence performance of the reward function.
+For $Ar$ values ranging from 1.75 to 0, the reward function exhibits clear convergence points, stabilizing after a plateau phase.
+This indicates that the agent has learned effective strategies, achieving stable returns under these conditions.
+However, the convergence speed of DRL training varies significantly across different $Ar$ values.
+For $Ar$ values of 1.75, 1.5, and 1.25, convergence signs appear around 2,000 iterations.
+For $Ar$ values of 1 and 0.75, convergence signs emerge around 1,000 iterations.
+Despite setting the total training iterations to 5,000 for $Ar$ values between 1.75 and 0.75, we observed that the reward function converges much earlier for $Ar$ values of 1 and 0.75, indicating a faster convergence speed and the agent’s ability to quickly learn effective strategies.
+For $Ar=0.5$, convergence signs appear after 4,000 iterations.
+For $Ar=0.25$, convergence signs emerge after 8,000 iterations.
+For $Ar=0.1$, convergence signs are observed after 10,000 iterations.
+For $Ar=0$, convergence signs appear after 12,000 iterations.
+In the range of $Ar=0.5$ to 0, the DRL training convergence speed is significantly slower compared to $Ar$ values greater than 0.5.
+This reflects a lower training efficiency and greater difficulty in exploring effective strategies for the agent when $Ar$ is less than or equal to 0.5.
+Furthermore, in the case of different $Ar$, DRL training was conducted in parallel using 60 CFD environments, allowing the DRL algorithm to collect data simultaneously from multiple independent simulations.
+The parameters of the ANN were updated every 60 iterations, which means that the agent and the environment only interacted without updating the policy network for the first 60 episodes.
+After 60 episodes, the ANN network parameters started to update, resulting in an initial learning curve that exhibits a noticeable step-like increase.
+This parallel acceleration effect has also been observed in the research conducted by Rabault and KuhnleRabault and Kuhnle (2019a) and Wang and XuWang and Xu (2024a).
 III.3 Analysis of the Controlled Flow
@@
 Figure 14: When $Ar=2$, 1.75, 1.5, 1.25, 1, 0.75, the vortex shedding of the elliptical cylinder can be completely suppressed.
-The streamwise snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$.
+The instantaneous vorticity snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$.
 These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the result of control, respectively.
@@
 By the third time point ($T_{3}$), the vortices behind the elliptical cylinder are completely suppressed, and the regions of positive and negative vorticity in the flow field nearly achieve symmetry, demonstrating high stability.
-Figure 14: When $Ar=0.5$, 0.25, 0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed.
-The streamwise snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$.
-These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.
 Fig.
@@
 At $T_{1}$, the vortex shedding behind the cylinder is fully developed, $T_{2}$ represents the control phase, and $T_{3}$ provides a snapshot of the flow field after control cessation.
+Figure 15: When $Ar=0.5$, 0.25, 0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed.
+The instantaneous vorticity snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$.
+These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.
 For $Ar=0.5$ and $Ar=0.25$, During the $T_{1}$ phase, vortices are observed detaching from the cylinder with certain frequency and intensity, forming a Kármán vortex street in the downstream wake.
@@
 Multi-level vortex shedding occurs, with primary and secondary vortices differing in shape, frequency, and intensity.
-At T2, vortices behind the cylinder are elongated and delayed, and the recirculation area significantly enlarges.
+At $T_{2}$, vortices behind the cylinder are elongated and delayed, and the recirculation area significantly enlarges.
 The instability of the recirculation area manifests as asymmetry, swinging, and oscillation phenomena.
@@
 IV CONCLUSIONS
-In this study, we investigate the use of DRL-based AFC to suppress vortex shedding, reduce drag, and mitigate lift fluctuations of elliptical cylinders at a Reynolds number of 100.
-The $Ar$ of the elliptical cylinder ranges from an ellipse ($Ar=2.0$) to a circle ($Ar=1.0$), and finally to a flat plate ($Ar=0$).
-We employ the PPO algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of the elliptical cylinder.
-The primary objective of this work is to gain further insights into the control capabilities of DRL algorithms for complex flow systems.
+In this study, we investigated the use of DRL-based AFC to suppress vortex shedding, reduce drag, and mitigate lift fluctuations of elliptical cylinders at a $Re=100$.
+The $Ar$ of the elliptical cylinder ranged from an elliptical cylinder ($Ar=2.0$) to a circle cylinder ($Ar=1.0$), and finally to a flat plate ($Ar=0$).
+We employed the PPO algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of the elliptical cylinder.
+The primary objective of this work was to gain further insights into the control capabilities of DRL algorithms for complex flow systems.
 The main findings of this study can be summarized as follows:
@@
 However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the reduction rates in the drag coefficient around the cylinder are respectively 0.9%, 2.1%, 3.9%, 8.0%, and 15.7%, while the reduction rates in lift coefficient are 99.7%, 99.4%, 98.9%, 98.9%, and 95.2%.
-These data indicate that the AFC strategy based on DRL is capable of significantly minimizing the periodically fluctuating drag coefficient to minimal values and stabilizing it, while also maintaining the oscillating lift coefficient near zero.
-This demonstrates the efficacy of DRL in AFC applications, effectively reducing the lift and drag around elliptical cylinder with $Ar$ ranging from 0.75 to 1.75.
-Visualizations of velocity fields reveal that the original periodically shedding vortices transition to stable flow fields with no vortex shedding after flow control.
-Instantaneous vorticity maps during the control process show that with the activation of synthetic jets, the vortices downstream of the elliptical cylinder are progressively elongated.
-The originally unstable recirculation area behind the cylinder transforms into a stable, symmetrical recirculation bubble, ceasing the production of new vortices.
+The originally periodic oscillations of the drag coefficient are reduced to their minimum value and stabilized.
+Simultaneously, over 95% of the lift coefficient is suppressed, ultimately stabilizing close to zero.
+With the activation of synthetic jets, the oscillating recirculation bubble in the baseline flow is gradually elongated and transformed into a stable and symmetrical recirculation bubble.
+Simultaneously, the existing vortices in the wake region become progressively elongated, ceasing the generation of new vortices.
+The originally periodic shedding of vortices transitions into a stable flow field, eliminating the phenomenon of vortex shedding.
 These results underscore the robust adaptability of DRL-based AFC across different elliptical cylinder shapes ($Ar$ between 0.75 and 1.75), and its capability of learning optimal control strategies directly through interaction with the environment to achieve a reduction in drag and lift coefficients around the cylinder.
 Moreover, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the mass flow rate ratios of synthetic jets relative to the inlet flow rate are 0.7%, 0.9%, 0.6%, 0.1%, and 1.0% respectively.
-This underscores that the flow control strategy efficiently utilizes less than 1% of external excitation energy, exemplifying its energy-saving capability for effective flow control.
+This indicates the effectiveness of the flow control strategy, which utilizes less than 1% of external excitation energy, showcasing its significant energy-saving attribute.
   • When the $Ar$ of the elliptical cylinder decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the drag coefficient and lift coefficient are 26.9%, 35.9%, 37.6%, and 43.6%, respectively, while the corresponding reduction rates for the lift coefficient are 68.0%, 66.1%, 50.2%, and 60.1%.
 Despite the effective reduction in both drag and lift coefficients achieved by the DRL-based AFC strategy, vortex shedding phenomena have not yet stabilized compared to cases with larger $Ar$ values.
-The initial phase of control exhibits distinct fluctuations, followed by low-frequency and low-amplitude oscillations in the drag and lift coefficients, which are lower than the baseline.
-Particularly, for $Ar$ values of 0.1 and 0, the drag and lift coefficients experience a sharp decrease initially, followed by moderate-amplitude, low-frequency periodic oscillations.
-This indicates that for elliptical cylinders with $Ar$ ranging from 0.5 to 0, the AFC control strategy can partially suppress lift and reduce drag.
-From the vorticity plots of the control process, for $Ar$ values of 0.5 and 0.25, the recirculation region behind the elliptical cylinder is elongated during the initial phase of control, with symmetric recirculation bubbles.
-As the control progresses, the recirculation region becomes unstable, exhibiting lateral oscillations and the shedding of long, narrow vortices.
-Similarly, for $Ar$ values of 0.1 and 0, the recirculation region undergoes elongation during the control phase, with high-frequency and large-amplitude oscillations, accompanied by the shedding of new vortices.
-After a significant period of control, the shedding of vortices is somewhat alleviated, with a decrease in both the frequency and intensity of vortex shedding.
+The drag and lift coefficients exhibit significant fluctuations during the initial stages of control, eventually settling into oscillatory states with smaller amplitudes and frequencies compared to the baseline flow.
+In particular, for elliptical cylinders with $Ar$ values of 0.1 and 0, the lift and drag coefficients continue to exhibit periodic oscillations.
+This indicates that for elliptical cylinders within the $Ar$ range of 0.5 to 0, the AFC control strategy can partially suppress lift and reduce drag.
+For $Ar$ of 0.5 and 0.25, the recirculation region behind the elliptical cylinder initially elongates, exhibiting symmetric recirculation bubbles.
+As control progresses, the recirculation region becomes unstable, displaying lateral oscillations and the shedding of long, narrow vortices.
+For $Ar$ values of 0.1 and 0, the recirculation region remains unstable throughout the control phase, accompanied by the shedding of new vortices.
+However, compared to the baseline flow, the frequency and intensity of vortex shedding decrease.
 Correspondingly, for elliptical cylinders with $Ar$ values of 0.5, 0.25, 0.1, and 0, the ratios of the mass flow rates of the synthetic jets relative to the incoming flow rate are 4.9%, 12.6%, 49.9%, and 52.1%, respectively.
-This suggests that despite the use of higher external energy, the lift coefficient cannot be completely controlled to prevent oscillations and fluctuations.
-In conclusion, for elliptical cylinders with $Ar$ ranging from 1.75 to 0.75, intelligent agents trained using the PPO algorithm have adeptly learned to adopt an energy-efficient flow control strategy.
-The control policies based on the PPO algorithm significantly reduce lift around the elliptical cylinder, greatly diminish drag, and completely suppress vortex shedding in the wake.
-Regarding elliptical cylinder with $Ar$ values ranging from 0.5 to 0, the control strategies obtained from agents trained with the PPO algorithm can moderately alleviate vortex shedding and reduce the amplitude and frequency of fluctuations in lift and drag coefficients.
+This indicates that despite using higher external energy, it is still not possible to suppress the oscillations and fluctuations in the lift coefficient.
+To summarize, for elliptical cylinders with $Ar$ ranging from 1.75 to 0.75, intelligent agents trained using the PPO algorithm have adeptly learned to adopt an energy-efficient flow control strategy.
+The control policies based on the PPO algorithm significantly reduce $C_{L}$ around the elliptical cylinder, greatly diminish $C_{D}$, and completely suppress vortex shedding in the wake.
+Regarding elliptical cylinder with $Ar$ values ranging from 0.5 to 0, the control strategies obtained from agents trained with the PPO algorithm can moderately alleviate vortex shedding and reduce the amplitude and frequency of fluctuations in $C_{D}$ and $C_{L}$.
 These control results demonstrate the adaptability of DRL-based AFC strategies to various geometric shapes, their robustness against uncertainties, and their ability to handle complex problems.
-This makes DRL-based AFC particularly suitable for controlling complex fluid dynamic systems.
 ACKNOWLEDGMENTS

### v4

@@
 Abstract
-This study investigates the effectiveness of active flow control (AFC) technology supported by deep reinforcement learning (DRL) applied to flows around an elliptical cylinder at $Re=100$.
-We vary the aspect ratio ($Ar$) of the elliptical cylinder from an elliptical cylinder ($Ar=2.0$) to a circular cylinder ($Ar=1.0$), and ultimately to a flat plate ($Ar=0$).
+We study the adaptability of deep reinforcement learning (DRL)-based active flow control (AFC) technology for bluff body flows with complex geometries.
+It is extended from a cylinder with $Ar=1$ to a flat elliptical cylinder with $Ar=2$, slender elliptical cylinders with $Ar$ less than 1, and a flat plate with $Ar=0$.
+The robustness and adaptability of DRL-based control technology will be assessed under varying levels of flow instability around bluff bodies.
 We utilize the Proximal Policy Optimization (PPO) algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of a cylinder to achieve reduction in drag, minimization of lift, and suppression of vortex shedding.
-The robustness and adaptability of DRL-based control techniques are examined across different geometric configurations.
 Our research findings indicate that, for elliptical cylinders with $Ar$ between 1.75 and 0.75, the reduction in drag coefficient ranges from 0.9% to 15.7%, and the reduction in lift coefficient ranges from 95.2% to 99.7%.
@@
 This reflects the control strategy’s significant reduction in both drag and lift coefficients, while also alleviating vortex shedding.
-It is expected underscore the adaptability and potential of DRL-based AFC in controlling complex fluid dynamics across diverse geometric configurations.
+The interaction and nonlinear development of vortices in the wake of elliptical cylinders lead to complex flow instability, and DRL-based AFC technology shows adaptability and potential in addressing flow control problems for this type of bluff body flow.
 preprint: AIP/123-QED
 I INTRODUCTION
-The development of active flow control (AFC) technology has progressed from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineering.Collis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021) This technology actively intervenes the flow field to improve flow characteristics, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiency.Collis and Joslin (2004); Wang et al.
+The development of active flow control (AFC) technology progresses from theoretical exploration to practical application, gradually becoming an important branch within the domains of fluid mechanics and engineering.Collis and Joslin (2004); Jahanmiri (2010); Awada, Younes, and Ilinca (2021) This technology actively intervenes the flow field to improve flow characteristics, such as reducing drag, controlling vortex shedding, enhancing mixing efficiency, or preventing flow separation, thereby enhancing system performance and efficiency.Collis and Joslin (2004); Wang et al.
 (2018) In its early stages, AFC research primarily focused on theoretical analysis and simple experiments to explore the influence of external inputs, such as jets, vibrations, or magnetic fields, on fluid flow dynamics.Jahanmiri (2010); Wang et al.
-(2018); Awada, Younes, and Ilinca (2021) With the advancement of computational fluid dynamics (CFD) and increased computational power, researchers have been able to delve deeper into complex flow responses through numerical simulations and develop more effective control strategies.Aram et al.
+(2018); Awada, Younes, and Ilinca (2021) With the advancement of computational fluid dynamics (CFD) and increased computational power, researchers are able to delve deeper into complex flow responses through numerical simulations and develop more effective control strategies.Aram et al.
 (2018); Jain, Yeo, and Chopra (2010); Li et al.
-(2022) However, as fluid problems are typically characterized by a high degree of nonlinearity and multiscale nature, involving high-dimensional state spaces and complex system dynamicsPozorski and Wacławczyk (2020); Reichstein et al.
-(2019), many engineering problems are often intractable using existing models.
-Sophisticated control strategies that require consideration of a wide range of spatiotemporal scales are indispensable when addressing phenomena such as unstable flows, turbulence, and multiphase flows.Subramaniam (2020)
-Recent advances in machine learning and deep learning technologies have revolutionized the field of artificial fluid control Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020).
-Deep Reinforcement Learning (DRL) combines deep learning and reinforcement learning to optimize decision-making through interactions with the environment Arulkumaran et al.
+(2022) However, due to the high degree of nonlinearity and multiscale nature of fluid problems, which involve high-dimensional state spaces and complex system dynamics, many engineering problems are often intractable using existing models.Pozorski and Wacławczyk (2020); Reichstein et al.
+(2019); Vinuesa (2024) Sophisticated control strategies that require consideration of a wide range of spatiotemporal scales are indispensable when addressing phenomena such as unstable flowsVinuesa (2024), turbulenceRen, Rabault, and Tang (2021), and multiphase flows.Subramaniam (2020)
+State-of-the-art machine learning technologies profoundly impact the field of active fluid control.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Ren, Hu, and Tang (2020) Deep Reinforcement Learning (DRL) combines deep learning and reinforcement learning to optimize decision-making through interactions with the environment.Arulkumaran et al.
 (2017); François-Lavet et al.
 (2018); Wang et al.
-(2020).
-DRL excels at extracting complex features from large datasets and making optimal decisions based on these features Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015).
-By using deep neural networks, DRL is capable of handling high-dimensional state spaces and enables machines to tackle previously intractable decision-making challenges Arulkumaran et al.
+(2020) DRL excels at extracting complex features from large datasets and making optimal decisions based on these features.Janiesch, Zschech, and Heinrich (2021); LeCun, Bengio, and Hinton (2015) By using deep neural networks, DRL handles high-dimensional state spaces and enables machines to tackle previously intractable decision-making challenges.Arulkumaran et al.
 (2017); François-Lavet et al.
 (2018); Wang et al.
-(2020).
-DRL’s nonlinear modeling capabilities and adaptive learning mechanisms allow it to learn complex strategies directly from raw sensory data without manual feature engineering Dargazany (2021).
-It continuously improves its strategies through interactions with the environment Wang et al.
-(2020), making it advantageous for dynamic decision-making problems François-Lavet et al.
+(2020) DRL’s nonlinear modeling capabilities and adaptive learning mechanisms allow it to learn complex strategies directly from raw sensory data without manual feature engineering.Dargazany (2021) It continuously improves its strategies through interactions with the environment, making it advantageous for dynamic decision-making problems.François-Lavet et al.
 (2018); Wang et al.
-(2020).
-DRL has found applications in various fields including gaming, autonomous vehicles, robotics, natural language processing, and computer vision.
-In particular, it has achieved distinguished success, such as AlphaGo’s victory over world champions in gaming Granter, Beck, and Papke (2017), route planning and decision-making optimization in autonomous driving Sallab et al.
-(2017), and complex physical tasks learning in robotics Zhao, Queralta, and Westerlund (2020).
-With progressive advancement in the DRL technology, there has been a significant increase in the application of DRL in the field of fluid dynamics as well.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al.
+(2020) DRL finds applications in various fields including gaming, autonomous vehicles, robotics, natural language processing, and computer vision.
+In particular, it achieves distinguished success, such as AlphaGo’s victory over world champions in gaming, route planning and decision-making optimization in autonomous driving, and learning complex physical tasks in robotics.Granter, Beck, and Papke (2017); Zhao, Queralta, and Westerlund (2020); Sallab et al.
+(2017)
+With progressive advancements in DRL technology, there is a significant increase in the application of DRL in the field of fluid dynamics as well.Brunton, Noack, and Koumoutsakos (2020); Brunton and Noack (2015); Garnier et al.
 (2021); Razdan and Shah (2022) DRL methods are highly suited to handle the high-dimensional and complex state spaces intrinsic to fluid dynamics.Xie et al.
 (2023) Through the function approximation capabilities of deep neural networks, DRL is adept at processing fluid systems and discerning flow patterns and structures.Zhao et al.
-(2024) DRL models, through their interaction with the environment and trial-and-error processes, can independently learn and refine strategies to meet diverse fluid conditions.Husen, Chaudary, and Ahmad (2022); Collis and Joslin (2004); Jahanmiri (2010) For instance, DRL has been applied to shape optimization of fluid systems and enabled automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal strategies.Bhola et al.
+(2024) DRL models, through their interaction with the environment and trial-and-error processes, independently learn and refine strategies to meet diverse fluid conditions.Husen, Chaudary, and Ahmad (2022); Collis and Joslin (2004); Jahanmiri (2010) For instance, DRL is applied to shape optimization of fluid systems, enabling the automatic discovery and application of effective shape adjustments to achieve optimization goals by intelligently exploring the design space and learning optimal strategies.Bhola et al.
 (2023) Moreover, the long-term reward mechanism intrinsic to DRL allows for the consideration of future impacts when formulating strategies.
-This unique feature has allowed DRL to shine in many AFC problems where DRL’s focus on long-term rewards enables control models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanics.Vinuesa et al.
+This unique feature allows DRL to excel in many AFC problems, where DRL’s focus on long-term rewards enables control models to balance short-term gains with future outcomes, thereby better addressing control and optimization challenges within fluid mechanics.Vinuesa et al.
 (2022); Vignon, Rabault, and Vinuesa (2023); Viquerat et al.
 (2022); Vinuesa (2024)
-In literature, DRL has attracted considerable interest for flow control around blunt bodies, as flow around blunt body tends to involve strong non-linearities such as flow separation and recirculation, traditional control methods often struggle in achieving control objectives whereas DRL models have shown promising results.
+In the literature, DRL attracts considerable interest for flow control around bluff bodies, as flow around these bodies tends to involve strong nonlinearities such as flow separation and recirculation.
+Traditional control methods often struggle to achieve control objectives, whereas DRL models show promising results.
 Pioneered by Rabault et al.
@@
 Heess et al.
-(2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024a) applied the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder.
-Consistently, they achieved an approximate 8% reduction in drag in their validation experiments.
+(2017), Ren, Rabault, and Tang Ren, Rabault, and Tang (2021), and Wang and Xu Wang and Xu (2024a) apply the Proximal Policy Optimization (PPO) algorithm to AFC in the flow around a circular cylinder.
+Consistently, they achieve an approximate 8% reduction in drag in their validation experiments.
 Furthermore, Li and ZhangLi and Zhang (2022), Wang et al.Wang et al.
-(2022a), Wang and XuWang and Xu (2024b) and Wang et al.Wang et al.
-(2022b)He et al.
-(2023) respectively conducted flow control studies on confined circular cylinder, square cylinder and airfoil using DRL.
+(2022a), Wang and XuWang and Xu (2024b), Wang et al.Wang et al.
+(2022b), and He et al.He et al.
+(2023) respectively conduct flow control studies on confined circular cylinders, square cylinders, and airfoils using DRL.
 In addition, Fan et al.Fan et al.
-(2020) initiated the application of DRL to AFC experiments, optimizing power efficiency by adjusting the rotational speed of auxiliary circular cylinders positioned behind another circular cylinder.
-Ren, Wang, and TangRen, Wang, and Tang (2021) proposed an AFC strategy to conceal the hydrodynamic characteristics of circular cylinders, such as strong shear forces and periodic shedding vortices.
-Studies on flow around constrained circular cylinders under different blockage ratios have also received extensive attention in the field Rabault et al.
+(2020) initiate the application of DRL to AFC experiments, optimizing power efficiency by adjusting the rotational speed of auxiliary circular cylinders positioned behind another circular cylinder.
+Ren, Wang, and TangRen, Wang, and Tang (2021) propose an AFC strategy to conceal the hydrodynamic characteristics of circular cylinders, such as strong shear forces and periodic shedding vortices.
+Studies on flow around constrained circular cylinders under different blockage ratios also receive extensive attention in the field.Rabault et al.
 (2019); Tang et al.
 (2020); Li and Zhang (2022); Wang et al.
-(2022a); Wang and Xu (2024a).
-Although active flow control strategies for both circular and square cylinders are pivotal in various engineering applications for managing aerodynamic and hydrodynamic forcesVignon, Rabault, and Vinuesa (2023); Rabault et al.
-(2020); Garnier et al.
-(2021); Razdan and Shah (2022), their extension to more intricate, non-ideal geometries poses significant challenges.
-The effectiveness of DRL strategies in such contexts remains largely unexplored, marked by uncertainties regarding their capacity to generate physically reasonable and adaptable control measures.
-Current literature indicates a distinct lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies.
-Based on the studies conducted by Jia and Xu Wang and Xu (2024a, b), we focus on the development of the DRL technqiue for active flow control around elliptical cylinders, with the objective of developing effective control strategies for drag reduction, flow stability enhancement, and overall performance improvement, addressing the research gap in applying DRL to different geometric shapes of elliptical cylinders.
+(2022a); Wang and Xu (2024a) The application of DRL-based AFC technology in bluff body flows is summarized in Tab.
+1.
+DRL-based AFC control technology is successfully applied to square, circular, and airfoil flows.
+Additionally, multiple studies demonstrate that DRL-based flow control strategies exhibit robustness across different Reynolds numbers and achieve favorable control outcomes in fluid mechanics experiments.
+This indicates that DRL-based AFC technology provides effective control strategies in challenging high-dimensional and nonlinear scenarios.
+Table 1: Application of AFC technology based on DRL in the flow around bluff body.
+Re Bluff Body Control measures DRL algorithm Drag Reduction Reference
+100 Cylinder Synthetic Jets PPO 8% Rabault et al.Rabault et al. (2019)
+100 Cylinder Synthetic Jets PPO 8% Wang and XuWang and Xu (2024a)
+100 Cylinder Synthetic Jets PPO 8% Wang et al.Wang et al. (2022a)
+120 Cylinder Synthetic Jets S-PPO-CMA 18.4% Paris, Beneddine, and DandoisParis, Beneddine, and Dandois (2021)
+100-300 Cylinder Synthetic Jets PPO 6%-24% He et al.He et al. (2023)
+100-400 Cylinder Synthetic Jets PPO 6%-39% Tang et al.Tang et al. (2020)
+1,000 Cylinder Synthetic Jets PPO 30% Ren, Rabault, and TangRen, Rabault, and Tang (2021)
+10,160 Cylinder Two fast-rotating smaller cylinders TD3 30% Fan et al.Fan et al. (2020)
+100 Square Synthetic Jets SAC 14% Wang et al.Wang et al. (2022a)
+100 Square Synthetic Jets SAC 14% Chen et al.Chen et al. (2023)
+100 Square Synthetic Jets SAC and TQC 17% Xia et al.Xia et al. (2024)
+100-400 Square Synthetic Jets SAC 14%-47% Wang and XuWang and Xu (2024b)
+100-500 Square Synthetic Jets SAC 14%-51% Jia and XuJia and Xu (2024)
+500-2,000 Square Synthetic Jets SAC 44%-61% Yan et al.Yan et al. (2023)
+The application of AFC strategies based on DRL is widely validated across various bluff body flow scenarios.
+Among these, research on confined circular cylinder flow under different blockage ratios is the most thorough.
+In contrast, studies related to elliptical cylinder flow are exceedingly rare in this domain.
+The aspect ratio of an elliptical cylinder significantly affects flow stability and vortex structure development, with the slender shape’s structural effects directly increasing the difficulty of flow control.
+While active flow control strategies for circular and square cylinders are crucial for managing aerodynamic and hydrodynamic performance in various engineering applications, extending these strategies to more complex, non-ideal geometries remains a major challenge.Vignon, Rabault, and Vinuesa (2023); Rabault et al.
+(2020) Under these circumstances, the effectiveness of DRL strategies is largely unexplored, particularly regarding whether they can produce physically reasonable and adaptable control measures.
+Current literature indicates a lack of comprehensive studies on the variability of DRL-based flow control strategies across different geometries, such as from streamlined bodies to various bluff bodies.Garnier et al.
+(2021); Razdan and Shah (2022) Therefore, this study preliminarily investigates whether it is possible to effectively reduce drag and potentially suppress vortex shedding behind an elliptical cylinder at $Re=100$.
+This lower Reynolds number allows for a detailed examination of the fundamental principles and mechanisms involved in the control strategy, enabling a rigorous evaluation without the additional complexities introduced by turbulence.
+The experiences and insights gained from this work are crucial for applying DRL to more practical engineering problems in fluid dynamics.
 The structure of this paper is organized as follows: sec.
-II introduces the problem description and provides a comprehensive procedure for implementing DRL.
+II details the problem addressed in this study and describes the numerical simulation methods and deep reinforcement learning techniques used.
+Additionally, it presents the framework for solving the AFC problem using DRL, providing a thorough description of the fundamental elements involved.
 Sec.
 III presents the main results and analysis.
+In sec.
+III.1, the baseline flow for elliptical cylinders with different $Ar$ is described, and probe positions and distributions are designed using physical information from the flow field.
+Sec.
+III.2 provides a detailed description of the controlled process for an elliptical cylinder with $Ar=0.75$ during DRL training.
+We also illustrate the impact of synthetic jet actions, whether blowing or suction, on the streamlines around the elliptical cylinder using detailed streamline.
+Sec.
+III.3 systematically tests the flow control performance of elliptical cylinders with $Ar$ ranging from 0 to 2 using DRL-based AFC, verifying the robustness of DRL-based AFC technology for elliptical cylinders with different geometric shapes.
+Sec.
+III.4 analyzes the controlled flow to explore the physical mechanisms of control.
 Finally, sec.
@@
 The solver implemented in OpenFOAM uses the finite volume method to discretize the computational domain into a mesh consisting of control volumes.
-To ensure numerical stability, a time step size of $\Delta t=0.0005$ is selected in the simulation.
+We choose the pisoFoam solver based on OpenFOAM, which is a transient solver for incompressible flow using the PISO algorithm.
+PISO algorithm is to iteratively correct the pressure and velocity fields to satisfy the incompressibility condition at each time step.
+The time and spatial derivatives are discretized using the Crank-Nicolson scheme and the second-order central difference scheme, respectively.
+A non-dimensional time step of $\Delta t=0.0005$ is selected.
+The CFL number, defined as $\left(\frac{\Delta t|U|}{\Delta x}\right)$—where $|U|$ represents the magnitude of the velocity vector per grid unit and $\Delta x$ is the grid length in the direction of the velocity—is less than 1, meeting the accuracy requirements.
 Grid scheme
@@
 Figure 4 illustrates the fundamental framework of DRL, including the computational process of a single training episode and the utilized formulas.
+We refer to the work of Rabault et al.Rabault et al.
+(2020) and use a multilayer perceptron with two fully connected layers of 512 neurons each as the policy network.
+This network structure is sufficient to handle the complexity of the flow control problem in this study.
+Simpler network structures lack the modeling capacity to capture the mapping relationships, while more complex network structures are harder to train.
+The architecture of the ANN needs to be designed based on the dimensions of the input and output data, which often relies on empirical knowledge.
 Figure 4: The Proximal Policy Optimization algorithm is trained using a parallel computing framework to simultaneously run multiple environments.
@@
 The flow rates of the upper and lower synthetic jets are constrained such that their sum is zero, i.e., $Q_{\text{lower}}+Q_{\text{upper}}=0$.
+They also conducted an evaluation of the momentum injected into the flow field by the four jets.
 The RL agent controls the velocity of the upper synthetic jet, while the lower synthetic jet has the same velocity but in the opposite direction.
@@
 $\omega$ is a weighting factor that quantifies the trade-off between minimizing drag and controlling lift fluctuations.
+Rabault and KuhnleRabault and Kuhnle (2019a) found that in the absence of a lift penalty, DRL produced a "cheating" strategy that resulted in reduced drag but was accompanied by a significant increase in lift.
+By incorporating a lift penalty in the reward function, this issue can be effectively mitigated.
+In previous work, the lift coefficient penalty factor is often set to 0.1Wang and Xu (2024a), 0.2Rabault et al.
+(2019); Tang et al.
+(2020) or 1Ren, Rabault, and Tang (2021).
+In this study, we use a penalty factor of 0.2, consistent with Rabault et al.Rabault et al.
+(2019)
 The time granularity parameter significantly affects the agent’s performance and the complexity of the learning task.
@@
 In particular, each action in our analysis is applied for a duration of 0.025 seconds, which corresponds to 25 numerical simulation time steps.
-Each episode consists of 100 control actions, resulting in a maximum total time for an episode of $T_{\max}=T_{100}=2.5$ nondimensional time units.
-The shedding frequency of vortices around the elliptical cylinder in this study ranges from 0.3 to 0.4, which corresponds to 6.25 to 8.33 vortex shedding cycles per episode.
+Considering that the Strouhal number for the flow around the elliptical cylinder in this study is between 0.3 and 0.4, we refer to the work of Rabault and KuhnleRabault and Kuhnle (2019a), where the duration of an episode corresponds to approximately 8 vortex shedding periods.
+Additionally, the work of Rabault et al.Rabault et al.
+(2019) considers an episode duration equivalent to about 6.5 vortex shedding periods.
+Therefore, in this study, the duration of an episode for the DRL training for AFC on the elliptical cylinder corresponds to approximately 6.25 to 8.33 vortex shedding periods.
+Each episode consists of 100 control actions, resulting in a maximum total time for an episode of $T_{\max}=T_{100}=2.5$ non-dimensional time units.
 The length of episodes is intentionally chosen to encompass multiple shedding cycles, which facilitates effective learning of the control algorithm by allowing the agent to observe and adapt to the flow dynamics over several cycles.
+Rabault and KuhnleRabault and Kuhnle (2019a) studied the impact of action update frequency on control strategies and found that approximately $T_{a}/T_{K}=10\%$, where $T_{K}$ represents the vortex shedding frequency and $T_{a}$ represents the duration of the action.
+In this study, the ratio of action duration to vortex shedding frequency is around 6% to 8%.
+This ensures that the action update frequency is sufficiently high for the ANN to control the system effectively while allowing the system enough time to respond to the control actions.
 Computing specifications
@@
 The deployment of DRL algorithms is based on theTensorforce library Schaarschmidt et al.
-(2018), while the definitions of artificial neural networks (ANNs) and gradient descent algorithms are facilitated using the TensorFlow open-source library Abadi et al.
+(2018), while the definitions of artificial neural networks and gradient descent algorithms are facilitated using the TensorFlow open-source library Abadi et al.
 (2016).
@@
 III RESULTS
-III.1 Baseline Flow
+III.1 Flow characteristics of baseline flow
 Vortex shedding phenomenon
@@
 6 illustrates two tests with different $Ar$, the configurations for other $Ar$ are similar and not further elaborated upon.
-III.2 Effectiveness of AFC
+III.2 Training of DRL-based flow control
+In this section, we use an elliptical cylinder with $Ar=0.75$ as an example to illustrate the interaction process between the agent and the environment during DRL training and the trial-and-error learning process to maximize the objective function.
+Fig.
+7 shows the evolution of the vorticity field of the elliptical cylinder throughout the DRL training process, clearly demonstrating how the vortices shed from the elliptical cylinder in the baseline flow are gradually expelled from the computational domain while suppressing the emergence of new vortices during the execution of active flow control.
+During the training process, six characteristic episodes are marked, and the values of $C_{D}$, $C_{L}$, and the synthetic jet velocity for the elliptical cylinder are extracted for these episodes.
+Fig.
+8 provides a more intuitive demonstration of the active flow control process and control strategies.
+Finally, Fig.
+9 offers a detailed analysis of the streamline patterns before and after the execution of synthetic jet actions, with a local magnification of the streamlines near the synthetic jet location.
+This clearly shows the changes in streamlines around the elliptical cylinder when the synthetic jet is in the blowing or suction state.
+This section provides a comprehensive description of the DRL training process and the operational mechanism of the synthetic jet.
+To facilitate understanding of how the agent learns the optimal flow control strategy through trial and error, we use the learning curve from the training process and the changes in the vorticity field around the elliptical cylinder to describe it.
+During DRL training, the agent first observes the state of the environment.
+We use several probes to provide physical information (velocity or pressure) of the flow field to characterize the state of the environment.
+The agent then calculates the initial reward function value based on the observed state and feeds the corresponding action (i.e., the jet velocity value) back to the CFD environment.
+Once the synthetic jets begin executing the action, they immediately influence the physical field around the elliptical cylinder, causing a change in its state.
+Next, the agent observes the state of the CFD environment again and calculates the reward value corresponding to that action.
+If the reward value increases, it indicates that the control strategy is beneficial for the objective function.
+If the reward value decreases, the agent will adjust the control strategy in the opposite direction.
+Through this iterative interaction and feedback process, the objective function is gradually optimized to the best control strategy.
+Figure 7: For an elliptical cylinder with $Ar=0.75$, observed the evolution of the instantaneous vorticity snapshots when DRL training.
+Specifically, focused on four representative training episodes: (a) Episode = 1, (b) Episode = 650, (c) Episode = 1000, and (d) Episode = 2000.
+The variations in the instantaneous vorticity snapshots at these episodes describe the influence of flow control techniques on vortex shedding.
+Figure 8: For an elliptical cylinder with $Ar=0.75$, the time series changes of the action, $C_{L}$, and $C_{D}$ during the characteristic episodes of DRL training.
+(a) The learning curve of DRL training for the elliptical cylinder with $Ar=0.75$.
+The values of the synthetic jet actions and the time series of $C_{L}$ and $C_{D}$ for episodes 1, 450, 650, 800, 1000, and 2000 are summarized in (b), (c), and (d), respectively.
+(b) The variation of the synthetic jet velocity with time.
+(c) The time series of $C_{L}$ for the characteristic episodes.
+(d) The time series of $C_{D}$ for the characteristic episodes.
+Note that the characteristic episodes in figures (b), (c), and (d) are consistent, with labels shown in (d).
+Figure 9: Instantaneous streamlines of an elliptical cylinder with $Ar=0.75$.
+(a) In the baseline flow, streamline around the elliptical cylinder.
+To highlight the changes in streamlines during the synthetic jet blowing or suction process, localized streamline near 90° and 270° of the elliptical cylinder are presented on the right side.
+(b) In the baseline flow, with the synthetic jets turned off, the streamlines flow along the elliptical cylinder’s surface and separate near the aft side.
+(c) In the baseline flow, the fluid flows closely along the surface and separates near the aft side.
+(d) Instantaneous streamlines of the elliptical cylinder when flow control is activated.
+The synthetic jets perform suction on the upper side and blowing on the lower side of the elliptical cylinder.
+(e) and (f) display localized streamline near the synthetic jets.
+(e) Streamlines when the upper jet is performing suction.
+(f) Streamlines when the lower jet is performing blowing.
+To demonstrate the suppression of wake vortices in an elliptical cylinder during DRL training, we provide an example using an elliptical cylinder with $Ar=0.75$ in Fig.
+7, observing the evolution of the instantaneous vorticity field throughout the DRL training process.
+The DRL training consists of 6000 episodes, including a rapid growth phase and a convergence phase.
+After an initial rapid rise during the first 2000 episodes, the reward function value approaches its maximum.
+In the 2000-6000 episode phase, the reward function value fluctuates near its maximum, indicating that the training has stably converged.
+We select four typical episodes during the surge phase to describe the changes in vortex shapes and positions.
+Fig.
+7(a) shows the wake field of the elliptical cylinder at episode 1, where regular, periodic vortex shedding is observed.
+DRL training begins with control applied to an already fully developed flow around the elliptical cylinder.
+In Fig.
+7(b), no new vortices shed under the influence of synthetic jet blowing or suction, and the recirculation region behind the elliptical cylinder elongates.
+The initially unstable and oscillating recirculation region behind the elliptical cylinder becomes symmetric and stable during flow control execution.
+Fig.
+7(c) shows that the vortices shed in the baseline flow gradually move downstream, stretching and interacting with each other.
+At this stage, the shape of the recirculation bubble stabilizes, and no new vortices shed.
+By episode 2000, Fig.
+7(d) illustrates that no vortex shedding exists in the computational domain, the area of the recirculation bubble no longer expands, and the entire computational domain reaches a stable state.
+This indicates that when the DRL training reward value converges near its maximum, the agent has learned the optimal control strategy, effectively suppressing vortex shedding in the wake field through the actions of the synthetic jet.
+To more clearly illustrate the variations in synthetic jet velocity, $C_{L}$, and $C_{D}$ during DRL training, we selected six characteristic episodes to extract the corresponding actions, $C_{L}$, and $C_{D}$.
+Fig.
+8(a) divides the DRL training into two phases: the reward function value rapidly increases between episodes = 0 and episodes = 2000, and after episode = 2000, the reward function enters a plateau, indicating that the training has converged.
+We selected episodes = 1, episodes = 450, episodes = 650, episodes = 800, episodes = 1000, and episodes = 2000 as characteristic episodes to observe the changes in $C_{D}$, $C_{L}$, and the control strategies for the elliptical cylinder.
+Fig.
+8(b) describes the variation of the synthetic jet velocity over time, representing the development of the flow control strategy.
+It can be observed that at episode 1, the agent attempts to use a relatively low jet velocity.
+As the number of episodes increases, the amplitude of the jet velocity gradually increases.
+By episode = 2000, the control strategy involves using a high jet velocity in the initial phase to control the flow field, followed by maintaining stability with much lower jet velocities.
+Fig.
+8(c) shows the time series of $C_{L}$ for the elliptical cylinder.
+At episode = 1, $C_{L}$ exhibits periodic fluctuations.
+As the number of episodes increases, the amplitude of $C_{L}$’s periodic fluctuations decreases.
+By episode = 2000, $C_{L}$ shows significant fluctuations in the initial phase of jet activation, then quickly reduces to near zero, maintaining zero for more than half the time, indicating that the lift around the elliptical cylinder is fully controlled.
+Fig.
+8(d) shows the time series of $C_{D}$ for the elliptical cylinder.
+At episode = 1, $C_{D}$ exhibits periodic oscillations.
+As the number of episodes increases, the amplitude and frequency of $C_{D}$ fluctuations decrease.
+By episode = 2000, $C_{D}$ only shows severe fluctuations at the beginning of control and then stabilizes near its minimum value.
+In Fig.
+9, we illustrate how the synthetic jets disrupt the streamlines along the sidewalls of an elliptical cylinder with $Ar=0.75$ during active flow control.
+Fig.
+9(a) shows the streamlines around the elliptical cylinder without active flow control.
+Fig.
+9(b) and Fig.
+9(c) provide zoomed-in views of the streamlines near the elliptical cylinder at the locations of the synthetic jets, which are not active at this point.
+Without control, the streamlines follow the contours of the elliptical cylinder and separate near the aft side.
+Fig.
+9(d) shows the streamlines around the elliptical cylinder with flow control activated.
+Fig.
+9(e) and Fig.
+9(f) again provide zoomed-in views of the streamlines at the synthetic jet locations, but this time with the jets active.
+When control is activated, the synthetic jets either blow or suction relative to the sidewalls, disrupting the previously continuous streamlines that followed the solid boundary of the elliptical cylinder.
+The noticable difference between uncontrolled baseline and controlled flow is that the flow stays attached after applying jet control.
+In Fig.
+9(b) and Fig.
+9(c), distinct flow separation can be observed in the rear-end.
+With synthetic jet, the flow stays attached after the suction/blowing point in Fig.
+9(e) and Fig.
+9(f).
+This illustrates the mechanism by which the synthetic jets influence the flow, breaking the streamlines and altering the flow pattern around the elliptical cylinder.
+III.3 DRL-based AFC for various ellipses
 Using the DRL algorithm, we investigate the performance of actively controlling the magnitude of synthetic jets for targeted flow control throughout the process of elliptical cylinders transitioning from elongated bodies to circular cylinders, and finally to a normal flat plate, with aspect ratios ranging from 2 to 0 at $Re=100$.
 The evaluation of the control’s effectiveness and performance is based on several criteria: intensity of vortex shedding after control, reductions in $C_{D}$, weakening of $C_{L}$, the actions (magnitude of the synthetic jets velocities), and the learning curves observed during the training process.
-For a detailed demonstration of vortex suppression, we present in Fig.
-7 an example using an elliptical cylinder with $Ar=0.75$, observing the evolution of the instantaneous vorticity field throughout the DRL training process.
-We select four representative snapshots of the instantaneous vorticity field to depict changes in vortex shapes and positions.
-Labels A,B,C,D and E in Fig.
-7(a) denote the five vortices initially shed.
-By the time depicted in Fig.
-7(b), these initial five vortices have progressed downstream without new vortices being shed under the influence of flow control.
-In Fig.
-7(c), labeled as A,B,C,D and E, the vortices have developed near the exit of the computational domain, exhibiting elongated shapes as they are about to be expelled from the domain.
-At this stage, the shape of the recirculation bubble has stabilized, and no new vortices are shed.
-By the time represented in Fig.
-7(d), the originally labeled five vortices have all been expelled from the computational domain, the area of the recirculation bubble no longer expands, and the entire computational domain has reached a stable state.
-Under the conditions of $Ar=1.75$, 1.5, 1.25, and 1 for the elliptical cylinder, the DRL-based AFC demonstrates similar control effectiveness.
-Figure 7: For an elliptical cylinder with $Ar=0.75$, observed the evolution of the instantaneous vorticity snapshots when DRL training.
-Specifically, focused on four representative training episodes: (a) Episode = 0, (b) Episode = 450, (c) Episode = 800, and (d) Episode = 2100.
-The initial shedding vortex are labeled as A, B, C, D and E, and them deformation and displacement are subsequently tracked.
-The variations in the instantaneous vorticity snapshots at these episodes describe the influence of flow control techniques on vortex shedding.
 Figure 10: Comparison of the velocity magnitude snapshots for an elliptical cylinder before and after control.
@@
 10 and Fig.
-9, respectively.
+11, respectively.
 As shown in Fig.
-8, the DRL-based AFC approach demonstrates consistent and effective suppression of vortex shedding and enhancement of flow stability in the wake field.
+10, the DRL-based AFC approach demonstrates consistent and effective suppression of vortex shedding and enhancement of flow stability in the wake field.
 At $Ar$ = 2, the baseline flow field is stable without vortex shedding, and the controlled flow field remains stable and similar to the baseline.
@@
 This phenomenon is visually depicted in Fig.
-9.
+11.
 In particular, at $Ar$ = 0.5, the application of flow control successfully suppresses the alternating vortex shedding, resulting in an elongated recirculation region.
@@
 Figure 12: Comparison of $C_{D}$ and $C_{L}$ before and after control.
-The background color represents the unstable stage of control, and the absence of background color means that the corresponding drag or lift has been in a stable stage during this period.
+The background color represents the unstable stage of control, while the absence of background color indicates that the corresponding $C_{D}$ or $C_{L}$ remains in a stable stage during this period.
 (a) and (b) $Ar=2$; (c) and (d) $Ar=1.75$; (e) and (f) $Ar=1.5$; (g) and (h) $Ar=1.25$; (i) and (j) $Ar=1$; (k) and (l) $Ar=0.75$.
@@
 13 provide a comparative analysis of the $C_{D}$ and $C_{L}$ before and after implementing the DRL-based AFC strategy, showcasing its effectiveness across a range of $Ar$ from 2 to 0.75 and from 0.5 to 0, respectively.
+The training duration $T_{\text{max}}$ is set to 2.5.
+Time is normalized using $T_{\text{max}}$, with the control period starting at Time = 0 and ending at Time = 1.
 As shown in Fig.
-10, the AFC strategy has no effect on the stability of both $C_{D}$ and $C_{L}$ at an $Ar$ of 2.
+12, the AFC strategy has no effect on the stability of both $C_{D}$ and $C_{L}$ at an $Ar$ of 2.
 However, when synthetic jets control is activated at $Ar$ values of 1.75, 1.5, 1.25, 1, and 0.75, there is an immediate and substantial reduction in $C_{D}$, which quickly reaches a minimum value.
@@
 The reductions in $C_{D}$ and $C_{L}$ at different $Ar$ values can be found in Tab.
-2.
+3.
 Clearly, the periodic oscillations of $C_{D}$ and $C_{L}$ originally present in the baseline flow are significantly reduced, ultimately transitioning to a stable state.
@@
 The impact of the AFC strategy on $C_{D}$ and $C_{L}$ at $Ar=0.5$ and $Ar=0.25$ is described in Fig.
-11.
+13.
 While both coefficients experience reductions compared to the baseline flow, they do not achieve the same level of stability observed at $Ar=0.75$ and higher.
@@
 At $Ar=0.1$ and $Ar=0$ are described in also in Fig.
-11, the controlled flow exhibits a sharp decrease in drag coefficient at the beginning of the control, followed by oscillations at a lower frequency and amplitude compared to the baseline flow.
+13, the controlled flow exhibits a sharp decrease in drag coefficient at the beginning of the control, followed by oscillations at a lower frequency and amplitude compared to the baseline flow.
 Although complete stability is not achieved, the average value of $C_{D}$ is significantly reduced.
@@
 The DRL-based AFC implemented in our study is demonstrated across Figs.
-8, 9, 10 and 11.
+10, 11, 12 and 13.
 This control strategy not only achieves high control performance but also maintains robust effectiveness, as evidenced by the quantified variations in $C_{D}$, $C_{L}$, and the ratio of synthetic jet mass flow rate to incoming flow rate, presented in Tab.
-2.
+3.
 $\overline{C}_{D,\text{Baseline}}$ and $\overline{C}_{L,\text{Baseline}}$ represents the time-averaged $C_{D}$ and $C_{L}$ of the elliptical cylinder once the baseline flow has fully developed.
@@
 Notably, the reduction in the coefficient of $C_{L}$ for elliptical cylinders, ranging from 99.7% to 95.2%, aligns with the achievement of a completely stable state (free from oscillations and fluctuations) for both $C_{D}$ and $C_{L}$ as illustrated in Fig.
+12.
+This consistency is also mirrored in the suppression of alternating vortex shedding in the Baseline, as depicted in Fig.
 10.
-This consistency is also mirrored in the suppression of alternating vortex shedding in the Baseline, as depicted in Fig.
-8.
 Figure 14: Action, which is the mass flow rate of the synthetic jets.
@@
 This explains why the drag reduction ratio gradually increases with the decrease in $Ar$ as shown in Tab.
-2.
+3.
 However, as the $Ar$ decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the $C_{L}$ are significantly lower than when $Ar$ is greater than or equal to 0.75.
 Within this range of $Ar$, only 50.2% to 68% of the $C_{L}$ is reduced, consistent with the fact that neither the $C_{L}$ nor the $C_{D}$ achieve the complete stability observed in Fig.
+13.
+Additionally, the baseline instability is not fully mitigated as shown in Fig.
 11.
-Additionally, the baseline instability is not fully mitigated as shown in Fig.
-9.
 This highlights the fact that there is still room for improvement in the current DRL-based AFC framework for controlling the flow instability of elliptical cylinders when $Ar$ is less than or equal to 0.5.
@@
 Furthermore, in Tab.
-2, we normalize the mass flow rates of the synthetic jets against the incoming far-field flow rate, reflecting their relative strength compared to the incoming flow.
+3, we normalize the mass flow rates of the synthetic jets against the incoming far-field flow rate, reflecting their relative strength compared to the incoming flow.
 $\overline{a}$ ratio (%) represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the far-field inflow boundary.
@@
 The above content details the effectiveness of the flow control strategy, with the corresponding DRL training process curves for elliptical cylinders with different $Ar$ values presented in Fig.
-13.
+15.
 First, we observed the overall trend of the reward function.
@@
 However, the convergence speed of DRL training varies significantly across different $Ar$ values.
-For $Ar$ values of 1.75, 1.5, and 1.25, convergence signs appear around 2,000 iterations.
-For $Ar$ values of 1 and 0.75, convergence signs emerge around 1,000 iterations.
+For $Ar$ values of 1.75, 1.5, and 1.25, signs of convergence appear around 2,000 iterations.
+For $Ar$ values of 1 and 0.75, signs of convergence emerge around 1,000 iterations.
 Despite setting the total training iterations to 5,000 for $Ar$ values between 1.75 and 0.75, we observed that the reward function converges much earlier for $Ar$ values of 1 and 0.75, indicating a faster convergence speed and the agent’s ability to quickly learn effective strategies.
-For $Ar=0.5$, convergence signs appear after 4,000 iterations.
-For $Ar=0.25$, convergence signs emerge after 8,000 iterations.
-For $Ar=0.1$, convergence signs are observed after 10,000 iterations.
-For $Ar=0$, convergence signs appear after 12,000 iterations.
+For $Ar=0.5$, convergence appears after 4,000 iterations.
+For $Ar=0.25$, convergence is reached after 8,000 iterations.
+For $Ar=0.1$, convergence can be observed after 10,000 iterations.
+For $Ar=0$, convergence appears after 12,000 iterations.
 In the range of $Ar=0.5$ to 0, the DRL training convergence speed is significantly slower compared to $Ar$ values greater than 0.5.
@@
 This parallel acceleration effect has also been observed in the research conducted by Rabault and KuhnleRabault and Kuhnle (2019a) and Wang and XuWang and Xu (2024a).
-III.3 Analysis of the Controlled Flow
+III.4 Analysis of the Controlled Flow
 Fig.
@@
 These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the result of control, respectively.
+(a) $Ar=2$, (b) $Ar=1.75$, (c) $Ar=1.5$, (d) $Ar=1.25$, (e) $Ar=1.0$, (f) $Ar=0.75$.
 In the study of elliptical cylinders, the phenomenon of vortex shedding and its control are crucial for understanding the structure of flow fields.
@@
 These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.
+(a) $Ar=0.5$, (b) $Ar=0.25$, (c) $Ar=0.1$, (d) $Ar=0$.
 For $Ar=0.5$ and $Ar=0.25$, During the $T_{1}$ phase, vortices are observed detaching from the cylinder with certain frequency and intensity, forming a Kármán vortex street in the downstream wake.